### v1

# Semiparametric Preference Optimization: Your Language Model is Secretly a Single-Index Model

Abstract
Aligning large language models to preference data is commonly implemented by assuming a known link function between the distribution of observed preferences and the unobserved rewards (e.g., a logistic link as in Bradley-Terry). If the link is wrong, however, inferred rewards can be biased and policies be misaligned. We study policy alignment to preferences under an unknown and unrestricted link. We consider an $f$-divergence-constrained reward maximization problem and show that realizability of the solution in a policy class implies a semiparametric single-index binary choice model, where a scalar-valued index determined by a policy captures the dependence on demonstrations and the rest of the preference distribution is an unrestricted function thereof. Rather than focus on estimation of identifiable finite-dimensional structural parameters in the index as in econometrics, we focus on policy learning, focusing on error to the optimal policy and allowing unidentifiable and nonparametric indices. We develop a variety of policy learners based on profiling the link function, orthogonalizing the link function, and using link-agnostic bipartite ranking objectives. We analyze these and provide finite-sample policy error bounds that depend on generic functional complexity measures of the index class. We further consider practical implementations using first-order optimization suited to neural networks and batched data. The resulting methods are robust to unknown preference noise distribution and scale, while preserving the direct optimization of policies without explicitly fitting rewards.
1 Introduction
Modern large language models (LLMs) are tuned using human or AI feedback (RLHF/RLAIF) to better align outputs with user preferences and safety desiderata (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022b, a; Nakano et al., 2021; Wu et al., 2021). A common setup interprets pairwise preferences as discrete choice under a latent reward, which is inferred and optimized, while constraining or penalizing the deviation from a reference model. This balances quality improvements with preservation of language abilities.
Linking preferences to rewards is usually done by assuming a particular choice model, such as Bradley-Terry (logistic link) or Thurstone (probit link), so that, given demonstrations and a reward function, the distribution of preferences becomes fully specified (Rafailov et al., 2023; Zhan et al., 2023; Glaese et al., 2022; Ziegler et al., 2019; Ibarz et al., 2018). A prominent example is Direct Preference Optimization (DPO, Rafailov et al., 2023), which uses the Bradley-Terry choice model. This, however, imposes a lot of structure on choice behavior, and misspecifying this link can bias inferred rewards and misalign policy optimization (Hong et al., 2025; Xu and Kankanhalli, 2024). Alternative approaches to alignment from preferences depart from structural/generative modeling of choice, focusing on loss aversion and label corruption (Ethayarajh et al., 2024; Liang et al., 2024; Kong et al., 2024).
However, specifying a known reward-preference link is not actually necessary for a structural/generative interpretation (meaning one that assumes preferences optimize a random utility or equivalently that preferences are generated by some conditional probability distribution given demonstrations and reward). Letting the link be arbitrary and unknown gives rise to semiparametric discrete choice models, which have been studied extensively in the econometrics literature (Cosslett, 1983; Klein and Spady, 1993; Manski, 1975, 1985; Horowitz, 1992; Kim and Pollard, 1990; Han, 1987; Sherman, 1993). This literature, however, focuses on reward (utility) models that are prescribed by finite-dimensional structural parameters, imposes stringent assumptions so that these structural parameters become identifiable from choice data, and develops methods and theory for estimation of and inference on these structural parameters, as they are themselves their target of interest.
In this paper, we study instead LLM alignment from preference data generated by a completely unknown link function and unknown reward function, where the target is the optimal reward-maximizing divergence-constrained policy. A key insight is that mere realizability of this optimal policy in our LLM policy class implies a so-called semiparametric single-index model for preferences, which has an arbitrary unknown link function. The policy class induces a class of indices (corresponding to reward or utility functions), which can be an arbitrary nonparametric class, and moreover the true index (corresponding to the true reward function) is generally not identifiable. We study policy learning methods and analyze bounds on the policy difference to the optimal policy (rather than convergence of structural estimates). Finally, we give practical gradient-based optimization algorithms that can be used to align neural-net-based models such as transformer LLMs.
We call the resulting general approach Semiparametric Preference Optimization (SPO). The primary algorithm we study is Orthogonalized SPO (OSPO), which performs empirical risk minimization with a loss function that depends on a family of univariate regression functions related to the unknown link. The analysis shows we can simply plug in estimates for these regression (for example, kernel regression estimates, since the regression is just univariate) and the method is locally invariant to the estimation: the error bounds scale quadratically in the error of the plug-in. This shows how we can perform almost as well as if we had known the link exactly (as, for example, DPO assumes we do) while being robust to not knowing the link at all (e.g., even when it is not logistic). In summary, SPO keeps the divergence-constrained policy search without an explicit reward model of DPO, while removing unverifiable assumptions on the reward-to-preference link and still provably optimally aligning policies to preferences.
Summary of contributions
  • From RLHF to single-index model. We formulate realizability of the $f$-divergence-constrained optimal policy and preferences observed with unknown link function as a single-index model of binary choice, with an index given by our policy class.
  • Semiparametric policy learners for RLHF. We develop the policy learners Profiled SPO (PSPO), Orthogonalized SPO (OSPO), Ranking SPO (RSPO) as three ways to search over policies that would maximize the likelihood or rank of their implied index in this model, without explicitly learning a reward model.
  • Policy-learning theory. We formalize an affine-invariant metric on indices and prove it controls policy difference to the optimal policy. We analyze guarantees of our policy learners in terms of this metric, showing consistency of PSPO (without a rate) and establishing rates for OSPO and RSPO that naturally depend on the functional complexity of the given classes.
  • Practical algorithms. We provide practical first-order algorithms suited to neural networks and batched data and study a simple empirical recalibration step at the end that enforces the divergence budget.
Roadmap
Section 2 derives the single-index formulation and the implications for policy learning. Section 3 presents and analyzes PSPO. Section 4 presents and analyzes OSPO. Section 5 presents and analyzes RSPO. Section 6 analyzes empirical calibration to the divergence budget. Section 7 discusses possible extensions. Section 8 surveys related work. Full proofs appear in the appendices.
2 RLHF as a Semiparametric Single-Index Model
This section formalizes the policy optimizations objective, describes the preference data we observe, and shows how they jointly yield a single-index model. We define an affine-invariant metric for error in the index error and show how it controls error in the divergence-constraint-calibrated policy. We then articulate the precise identification property needed to have hope to control this metric and prove the assumption holds under a simple condition when we have at least three actions.
2.1 The policy target
The target we focus on is the policy maximizing average rewards subject to a constraint on its deviation from a reference policy. This is a standard approach to align models to a new task while staying close to a pre-trained policy, preserving its language ability and preventing catastrophic forgetting. While the common divergence is Kullback-Leibler (KL), we here consider any $f$-divergence, of which KL is an example.
We let $x\in\mathcal{X}$ denote context (e.g., user query) and $y\in\mathcal{Y}$ action (e.g., LLM response). The mean reward of taking action $y$ in context $x$ is given by the unknown reward function $r^{\star}(x,y)$. For simplicity we focus on a finite action space throughout the paper, $|\mathcal{Y}|<\infty$, while $\mathcal{X}$ can be general Borel. Given convex $f:\mathbb{R}^{+}\to\mathbb{R}$ with $f(1)=0$, for any two probability mass functions $p,q\in\Delta^{\mathcal{Y}}$ on $\mathcal{Y}$ we define $D_{f}(p\|q)=\infty$ whenever $p(y)>0,q(y)=0$ for some $y$ and $D_{f}(p\|q)=\sum_{y:q(y)>0}q(y)f(p(y)/q(y))$ otherwise. Henceforth we will assume throughout that $f$ is twice continuously differentiable and strictly convex with derivative approaching -$\infty$ near 0. KL is given by setting $f(u)=u\log u$. Other examples satisfying these conditions include $\alpha$-divergences (which include KL), Jensen-Shannon divergence, and the combination of KL and $\chi^{2}$-divergence as used in Huang et al. (2024).
Given a reference policy $\pi_{\rm ref}:\mathcal{X}\to\Delta^{\mathcal{Y}}$, we are interested in the divergence-constrained reward-maximizing policy:
$\pi^{\star}\in$ $\operatornamewithlimits{argmax}_{\pi:\mathcal{X}\to\Delta^{\mathcal{Y}}}\mathbb{E}_{x}\sum_{y\in\mathcal{Y}}\pi(y\mid x)r^{\star}(x,y)$ (1)
$\text{s.t.}\quad\mathbb{E}_{x}D_{f}(\pi(\cdot\mid x)\|\pi_{\rm ref}(\cdot\mid x))\leq\kappa,$
where the expectations are taken with respect to a context distribution $x\sim P_{x}$.
Combining a simple convexity argument with the results of (Wang et al., 2023) yields a closed form for $\pi^{\star}$.
Assumption 1.
Let $\omega(x)=\sum_{y\in\operatornamewithlimits{argmax}_{y}r^{\star}(x,y)}\pi_{\rm ref}(y\mid x)$. Suppose $\mathbb{E}_{x}\omega(x)f(1/\omega(x))>\kappa$.
Theorem 1 (Closed form for $\pi^{\star}$).
Suppose Assumption 1 holds. There exist $\beta^{\star}>0$ and a function $\lambda^{\star}:\mathcal{X}\to\mathbb{R}$ such that
$\pi^{\star}(y\mid x)\;=\;\pi_{\rm ref}(y\mid x)\,(f^{\prime})^{-1}\!\left(\beta^{\star-1}\big(r^{\star}(x,y)-\lambda^{\star}(x)\big)\right),$ (2)
where $\mathbb{E}_{x}D_{f}(\pi^{\star}(\cdot\mid x)\|\pi_{\rm ref}(\cdot\mid x))=\kappa$.
Assumption 1 simply requires that a pure-reward-maximizing policy is not already feasible in the divergence constraint thus rendering it irrelevant.
Remark 1 (The case of KL).
The solution simplifies a lot for $f(u)=u\log u$, in which case $\pi^{\star}(y\mid x)\propto\pi_{\rm ref}(y\mid x)\exp(\beta^{\star-1}r^{\star}(x,y))$ for some $\beta^{\star}$.
2.2 Preference data
Toward optimizing Eq. 1, we assume we have access to preference data consisting of $n$ tuples $(x,y_{0},y_{1},z)\sim P,$ with $w=(x,y_{0},y_{1})\in\mathcal{W}$ drawn from some joint distribution with $x$-marginal matching $P_{x}$ (otherwise we can do importance sampling) and $z\in\{0,1\}$ being the indicator that $y_{1}$ is preferred to $y_{0}$. We index our data $(w_{i},z_{i})=(x_{i},y_{i0},y_{i1},z_{i})$ for $i=1,\dots,n$ and assume they are drawn iid.
We assume that preferences are drawn according to
$z\mid x,y_{0},y_{1}\sim\mathrm{Bernoulli}\left(\Phi^{\star}\left(r^{\star}(x,y_{1})-r^{\star}(x,y_{0})\right)\right),$ (3)
where $\Phi^{\star}$ is an unknown cumulative distribution function (CDF).
The key in this paper is that we let $\Phi^{\star}$ be completely unknown, rather than setting it to a known function like sigmoid or the normal CDF. This is important since we only care to understand rewards so that we can optimize policies, rather than fully model the distribution of $z$.
Remark 2 (Random utility interpretation).
Suppose $\epsilon_{1},\epsilon_{2}$ are independent of $x,y_{0},y_{1}$ with $\mathbb{P}(\epsilon_{1}-\epsilon_{2}\leq t)=\Psi^{\star}(t)$. Then $z=\operatornamewithlimits{argmax}_{i=0,1}r^{\star}(x,y_{i})+\epsilon_{i}$ has the distribution above. This corresponds to the feedback provider choosing the response with higher perceived reward, which is random with a conditional mean given by $r^{\star}(x,y_{i})$ and idiosyncratic shocks given by $\epsilon_{i}$. The shocks can represent variation in the population of feedback providers or any variation in a single individual’s assessment.
For simplicity, we assume that $P$ is symmetric in actions (i.e., measure of a set is preserved when we replace $(x,y_{0},y_{1})$ by $(x,y_{1},y_{0})$) so that $\mathbb{E}[h(x,y_{1})-h(x,y_{0})]=0$ for any function $h(x,y)$.
2.3 Reformulation as a single-index model
In policy optimization we search over a policy class $\{\pi_{\theta}:\theta\in\Theta\}$. We emphasize that throughout, we let $\theta$ be a completely abstract parameter, not necessarily finite dimensional. In this sense, the policy class can be parametric or nonparametric, and is generally understood to be large and rich.
Let us assume realizability of this class for our learning task in Eq. 1 in that $\pi^{\star}=\pi_{\theta^{\star}}$ for some $\theta^{\star}\in\Theta$. We will now show this essentially induces what is known as a single-index model (Ichimura, 1987) on the preference data.
Let us posit the following model for preference data: for $\theta\in\Theta$ and $\Psi$ a CDF,
$z\mid x,y_{0},y_{1}$ $\sim\mathrm{Bernoulli}\left(\Psi\left(t_{\theta}(x,y_{0},y_{1})\right)\right),$ (4)
$h_{\theta}(x,y)$ $=f^{\prime}\left(\frac{\pi_{\theta}(y\mid x)}{\pi_{\rm ref}(y\mid x)}\right),\qquad t_{\theta}(x,y_{0},y_{1})=h_{\theta}(x,y_{1})-h_{\theta}(x,y_{0}).$ (5)
Combining Theorem 1 with Eq. 3 and rearranging establishes that this is a valid model under policy realizability.
Assumption 2 (Policy realizability).
$\pi^{\star}=\pi_{\theta^{\star}}$ for some $\theta^{\star}\in\Theta$.
Proposition 1 (Policy realizability implies a single-index model).
Under Assumptions 1 and 2 we have that Eq. 4 holds for $\theta=\theta^{\star}$ and $\Psi(u)=\Psi^{\star}(u)=\Phi^{\star}(\beta^{\star}u)$.
Specifically, Eq. 2 can be rearranged as $r^{\star}(x,y)=\beta^{\star}h_{\theta^{\star}}(x,y)+\lambda^{\star}(x)$. Then $\lambda^{\star}$ is canceled by differencing the rewards of $y_{1}$ and $y_{0}$, and $\beta^{\star}$ is swallowed into the CDF as it is unrestricted.
Equation 4 is referred to as a single-index model (of binary choice) since it posits the existence of a parametric univariate composite statistic that summarizes all the relevant aspects of the inputs (here, $x,y_{0},y_{1}$), which is often known as an index. The rest is a completely unrestricted nonparametric dependence on this univariate quantity $t^{\star}=t_{\theta^{\star}}$. We will refer to $h_{\theta}$ as potential functions and to $t_{\theta}$ as indices.
We henceforth impose Assumptions 1 and 2 so that, by Proposition 1, $\theta^{\star}$, $t^{\star}$, and $\Psi^{\star}$ are well defined and characterize the data generating distribution.
Recall that, since $\theta$ is just an abstract parameter, the class of indices
$\mathcal{T}=\left\{t_{\theta}(x,y_{0},y_{1})=f^{\prime}\left(\frac{\pi_{\theta}(y_{1}\mid x)}{\pi_{\rm ref}(y_{1}\mid x)}\right)-f^{\prime}\left(\frac{\pi_{\theta}(y_{0}\mid x)}{\pi_{\rm ref}(y_{0}\mid x)}\right):\theta\in\Theta\right\},$
may be parametric or nonparametric and is in general as rich as the policy class $\{\pi_{\theta}:\theta\in\Theta\}$. As such, we will characterize $\mathcal{T}$ in terms of its own functional complexity, permitting rich and growing function classes.
2.4 From a single-index model to a policy
While policies are the ones parameterizing the index, via the parameterization by $\theta$, not every policy is feasible in the divergence constraint. Policies are invariant to additive transforms of rewards but not scaling, as that determines the temperature or divergence. Nonetheless, once we fix the divergence constraint we obtain scaling invariance as well.
Given a parameter $\theta$, which gives a potential function $h_{\theta}$, let us define the corresponding $\beta$-regularized policy as
$\pi_{\beta,\theta}(y\mid x)\;=\;\pi_{\rm ref}(y\mid x)\,(f^{\prime})^{-1}\!\left(\beta^{-1}\big(h_{\theta}(x,y)-\lambda_{\beta,\theta}(x)\big)\right),$
where $\lambda_{\beta,\theta}(x)$ is chosen so that $\pi_{\beta,\theta}(y\mid x)$ is a valid conditional probability mass function. With a slight abuse of notation let us further define the population-divergence-calibrated policy as
$\pi_{\kappa,\theta}(y\mid x)\;=\;\pi_{\beta_{\kappa,\theta},\theta}(y\mid x),$
where $\beta_{\kappa,\theta}>0$ is chosen so that $\pi_{\kappa,\theta}(y\mid x)$ satisfies $\mathbb{E}_{x}D_{f}(\pi_{\kappa,\theta}(\cdot\mid x)\|\pi_{\rm ref}(\cdot\mid x))=\kappa$ (which we know that $\pi^{\star}$ satisfies under Assumption 1). Namely, we can think of $\lambda_{\beta}(x)$ as being chosen so that, for any choice of $\beta$ and for any context $x$, the policy sums to one over $y$, and of $\beta_{\kappa,\theta}$ being then chosen such that the resulting policy has exactly the desired divergence.
We emphasize that here the calibration defining $\pi_{\kappa,\theta}(y\mid x)$ is done exactly on the true population distribution. We consider an empirical version in Section 6, which is an actual algorithm that can be implemented on a data sample and which we show incurs an additive error. Note also that by construction, under Assumption 1, we have $\pi^{\star}=\pi_{\kappa,\theta^{\star}}$.
Note that because of the affine transform by $\beta_{\kappa,\theta}>0,\lambda_{\kappa,\theta}(x)$, we must have that $\pi_{\kappa,\theta}$ and $\pi_{\kappa,\theta^{\prime}}$ are the same whenever the indices $h_{\theta}$ and $h_{\theta^{\prime}}$ are affine transforms of one another (with $x$-dependent location and $x$-independent scaling). Therefore, we do not actually need to learn $\theta^{\star}$ – we only need to learn it up to this equivalence. Correspondingly, rather than phrase errors as $\ell_{2}$ distance in $\theta$ or even $L_{2}$ distance in $h_{\theta}$, we will seek to control distance up to this equivalence:
Definition 1 (Scale- and $x$-location-invariant metric).
Define
$\rho(h)\;=\;\inf_{a>0}\left(\mathbb{E}_{x}\inf_{b(x)}\sum_{y}\pi_{\rm ref}(y\mid x)\,\big(a\,h(x,y)-b(x)-h_{\theta^{\star}}(x,y)\big)^{2}\right)^{1/2}.$
Clearly, if $\rho(h_{\theta})=0$ then, under Assumption 1, we have $\pi_{\kappa,\theta}=\pi^{\star}$. We would further like small $\rho$ to translate to small policy error. Before establishing this, we collect some of the basic regularity conditions we will use throughout.
Definition 2 (Norms).
Throughout, for $w$ functions $f(w)=f(x,y_{0},y_{1})$ we define $\|f\|_{p}$ as the $L_{p}$ norm over $w\sim P$, and for $(x,y)$-functions $g(x,y)$ we define $\|g\|_{p}$ as $L_{p}$ norm over $P_{x}\times\pi_{\rm ref}$. For $p=2$ we omit the subscript and write $\|\cdot\|$.
Assumption 3 (Basic regularity and coverage).
$M=\sup_{\theta\in\Theta}\|t_{\theta}\|_{\infty}<\infty$, $\sup_{\theta\in\Theta}\|\log f^{\prime\prime}(\pi_{\theta}(y\mid x)/\pi_{\rm ref}(y\mid x))\|_{\infty}<\infty$, and $\operatorname{esssup}_{x\sim P_{x}}\sup_{y_{0},y_{1}\in\mathrm{Support}(\pi_{\rm ref}(\cdot\mid x))}\frac{\pi_{\rm ref}(y_{0}\mid x)\pi_{\rm ref}(y_{1}\mid x)}{P(y_{0},y_{1}\mid x)}<\infty$.
The first two conditions ensure we are dealing with bounded quantities and the smoothness/curvature of objectives. The very last condition ensures that the data covers $\pi_{\rm ref}$ so that we can translate an index fitting well to the data (generated by $P$) to it inducing a small $\rho$ (defined by $\pi_{\rm ref}$). It is trivially satisfied when demonstrations are generated by drawing twice independently from $\pi_{\rm ref}$.
Remark 3 (Importance weighting).
If demonstrations are not generated by drawing twice independently from $\pi_{\rm ref}$, we can alternatively also introduce weighting by $\pi_{\rm ref}(y_{0}\mid x)\pi_{\rm ref}(y_{1}\mid x)/P(y_{0},y_{1}\mid x)$ (possibly clipped for stability). We can additionally reweight by $\frac{dP_{x}}{dP_{x}^{\mathrm{data}}}(x)$ if contexts $x$ were not generated by the same context distribution that defines our policy target Eq. 1 (we here assumed the context distributions were the same). Our analyses will easily carry over with weighted loss functions assuming bounded weights.
Theorem 2 (Affine-invariant index error controls policy error).
Suppose Assumptions 1, LABEL:, 2 and 3 hold, that $\pi_{\rm ref}(y\mid x)$ takes values in $\{0\}\cup[\epsilon,1]$ for $\epsilon>0$, and that, for each $\theta\in\Theta$ and $x\in\mathcal{X}$, define $\omega_{\theta}(x)=\sum_{y\in\operatornamewithlimits{argmax}_{y\in\mathcal{Y}}h_{\theta}(x,y)}\pi_{\mathrm{ref}}(y\mid x),$ and assume $\mathbb{E}_{x}\Bigl[\omega_{\theta}(x)\,f\!\bigl(1/\omega_{\theta}(x)\bigr)\Bigr]>\kappa$ for every $\theta\in\Theta$. Then there exists a constant $C<\infty$ such that
$\|\pi_{\kappa,\theta}-\pi^{\star}\|_{1}\ \leq\ C\,\rho(h_{\theta})\qquad\text{for all $\theta\in\Theta$}.$
Equivalently, we can bound the expected total variation distance between $\pi_{\kappa,\theta}(\cdot\mid x)$ and $\pi^{\star}(\cdot\mid x)$. We therefore henceforth focus on convergence and rates $\rho(h_{\hat{\theta}}) for$ a learner $\hat{\theta}$ as that would imply convergence and rates for policy error.
2.5 Identification of the target policy
In the previous section we argued that we need only learn $\theta^{\star}$ up to affine transformations of the potential function $h_{\theta^{\star}}$ it parameterizes, since all of those will lead to the same policy after divergence calibration. At the same time, however, under the single-index model Eq. 4, two indices are observationally equivalent if differences therein are monotonic transformations of one another, since this transformation can be captured by the freely varying CDF $\Psi$.
Therefore, to have any hope of controlling $\rho(h_{\theta})$ by making the potential $h_{\theta}$ explain the preference data, we need to assume that observational equivalence implies policy equivalence. We next posit this assumption and proceed to show it is in fact quite weak.
Assumption 4 (Observational equivalence implies policy equivalence).
For any $\theta\in\Theta$, if there exists a monotonic non-decreasing $m:\mathbb{R}\to\mathbb{R}$ with
$t_{\theta^{\star}}(x;y_{1},y_{0})=m\!\left(t_{\theta}(x;y_{1},y_{0})\right)\quad\text{$P$-a.s.},$
then there exist $a\geq 0$ and a measurable $b:\mathcal{X}\to\mathbb{R}$ such that
$h_{\theta^{\star}}(x,y)=a\,h_{\theta}(x,y)+b(x)\qquad\text{$P_{x}\times\pi_{\rm ref}$-a.s.}$
Note that this assumption is weaker than assuming that observational equivalence implies $\theta=\theta^{\star}$, as assumed in the econometrics literature that is interested in estimating $\theta^{\star}$ itself as a structural parameter (e.g., Cosslett, 1983; Klein and Spady, 1993).
In fact, unlike those strong identification conditions, Assumption 4 is quite natural and is in fact almost necessary when we have at least three actions, $|\mathcal{Y}|\geq 3$. The key observation is that whenever we have $x,y_{1},y_{2},y_{3}$ with $t_{\theta^{\star}}(x;y_{i},y_{j})=m(t_{\theta}(x;y_{i},y_{j}))$ for $(i,j)\in\{(1,2),(2,3),(1,3)\}$ then
$m(t_{\theta}(x;y_{1},y_{2}))+m(t_{\theta}(x;y_{2},y_{3}))$ $=t_{\theta^{\star}}(x;y_{1},y_{2})+t_{\theta^{\star}}(x;y_{2},y_{3})$
$=t_{\theta^{\star}}(x;y_{1},y_{3})$
$=m(t_{\theta}(x;y_{1},y_{3}))$
$=m(t_{\theta}(x;y_{1},y_{2})+t_{\theta}(x;y_{2},y_{3})).$
This imposes Cauchy’s functional equation (additivity) on a monotonic function. When the domain spanned by $t_{\theta}(x;y_{i},y_{j})$ is sufficiently rich, the only feasible solution for $m$ is a linear function (Aczél, 1966). Note that we need $\left\lvert\mathcal{Y}\right\rvert\geq 3$ to be able to make this rich enough, else all we can possibly get is the oddness of $m$. In the application of interest, of course, $\mathcal{Y}$ is quite big.
We now summarize one simple sufficient condition for using this to guarantee Assumption 4.
Theorem 3.
Suppose Assumption 3 holds. Suppose there exists $\delta>0$ such that for every $\theta\in\Theta$, $\left\lvert u\right\rvert<\|t_{\theta}\|_{\infty}$, $\left\lvert v\right\rvert<\delta$ with $\left\lvert u+v\right\rvert<\|t_{\theta}\|_{\infty}$ there exist some $x,y_{1},y_{2},y_{3}$ with $(x,y_{1},y_{2}),(x,y_{2},y_{3}),(x,y_{1},y_{3})\in\mathrm{Support}(P)$ and $u=t_{\theta}(x;y_{1},y_{2})$, $v=t_{\theta}(x;y_{2},y_{3})$. Then Assumption 4 holds.
3 Semiparametric RLHF Based on Profiling the Link (PSPO)
We now consider different policy learners motivated by a variety of existing semiparametric estimators of single-index models. We begin by adapting the method of Cosslett (1983), which profiles the link function $\Psi$ in a maximum likelihood estimation by letting it be freely chosen to maximize the likelihood for any given parameter $\theta$.
Intuitively, profiling removes the infinite-dimensional nuisance $\Psi$ by letting the data choose the monotone function that best explains outcomes for a given index. We then show uniform convergence of the profiled objective, population identification up to affine invariance, existence of empirical maximizers, and policy consistency.
3.1 The PSPO policy learner
Define the binary cross-entropy given both an index $t$ and a candidate link function $\Psi$:
$\ell(t,\Psi;w,z)=z\log\Psi\big(t(w)\big)+(1-z)\log\big(1-\Psi\big(t(w)\big).$
Recall we defined $w=(x,y_{0},y_{1})$.
If we knew $\Psi^{\star}$ we would plug it in and optimize the empirical average of the above likelihood over $t_{\theta}$e.g., this is the approach of DPO using the logistic link function. Instead, since we do not know $\Psi^{\star}$, we can consider both $t_{\theta}$ and $\Psi$ as parameters to maximize the likelihood over. Equivalently, for every $t_{\theta}$, we optimize the link, thus profiling it out of the objective. This profiled objective is defined as follows and its maximizer is the PSPO learner:
$L_{n}^{\rm PSPO}(\theta)=\sup_{\Psi\uparrow}\;\frac{1}{n}\sum_{i=1}^{n}\ell(t_{\theta},\Psi;w_{i},z_{i}),\quad\text{and}\quad\hat{\theta}_{\rm PSPO}\in\operatornamewithlimits{argmax}_{\theta\in\Theta}L_{n}^{\rm PSPO}(\theta),$ (6)
where $\Psi\uparrow$ denotes the restriction to monotonic non-decreasing functions $\mathbb{R}\to[0,1]$.
Similar to DPO, the approach here is to align the LLM policy by maximizing the likelihood of the reward function implied by the aligned policy. Recall, the policy is wrapped inside $t_{\theta}(w)=f^{\prime}\left(\frac{\pi_{\theta}(y_{1}\mid x)}{\pi_{\rm ref}(y_{1}\mid x)}\right)-f^{\prime}\left(\frac{\pi_{\theta}(y_{0}\mid x)}{\pi_{\rm ref}(y_{0}\mid x)}\right)$. Here the likelihood is Eq. 6, which profiles the link rather than assuming it is known and plugging it in.
3.2 Policy consistency of PSPO
We now establish rate-free consistency of PSPO, allowing for a nonparametric index class $\mathcal{T}$.
Theorem 4 (PSPO policy consistency).
[…]
$\rho(h_{\hat{\theta}_{\rm PSPO}})\to 0\quad\text{in probability.}$
[…]
3.3 First-order optimization
To optimize Eq. 6 in practice, especially when $\pi_{\theta}$ is a neural network, we can use a first-order (gradient-based) method. A simple approach would be to alternate taking a gradient step in $\theta$, taking $\Psi$ as fixed, with optimizing $\Psi$ for the current $\theta$. The former can be combined with any momentum and/or adaptive learning rate procedure. The latter is essentially a univariate monotone logistic regression and can be done using the Pool Adjacent Violators Algorithm (PAVA). This is akin to the Isotron algorithm (Kalai and Sastry, 2009; Kakade et al., 2011). Alternatively, differentiation through the isotonic regression can be done by implicit differentiation of PAVA KKT conditions or by training a separate network for $\Psi$ with monotonic constraints.
4 Semiparametric RLHF Based on Orthogonalizing the Link (OSPO)
We can only obtain rate-free consistency for PSPO as it can generally be ill-behaved because profiling the link function can lead to a rough loss surface. This was the observation of Klein and Spady (1993), who instead proposed a semiparametrically efficient approach. Intuitively, rather than aim to plug in the link $\Psi_{\theta}$ that profiles the likelihood at $\theta$ as in PSPO, instead we should consider how perturbing $\theta$ changes the likelihood, that is, consider the score. The score has two orthogonal components: the distribution of the index and the distribution of $z$ given the index. At $\theta^{\star}$, the latter is just the true link $\Psi^{\star}$, but away from it this is just a conditional probability. Since $z$ is binary, this is just a conditional expectation, which we can estimate this conditional expectation and plug it in. In success, the behavior is as though we used the true conditional expectation. This orthogonality phenomenon is akin to Double/Debiased Machine Learning (Chernozhukov et al., 2024). In the rest of the section we formally define the learner and analyze it.
4.1 The idealized quasi-likelihood
Consider the distribution of $z$ given an index $t_{\theta}$:
$g^{\star}_{\theta}(u)=P(z=1\mid t_{\theta}(w)=u).$
This orthogonalizes perturbations to $\theta$, which determine the index, and perturbations to the distribution of $z$ given the index. These live on orthogonal components of the tangent space: we can think of the former as “angle” (corresponding to changing the distribution of the index) and the latter as “magnitude” (corresponding to the distribution of $z$ given the index).
Solving the score equal to zero is conveniently equivalent to maximizing a quasi-likelihood given by plugging in $g^{\star}_{\theta}$ as the link:
$L_{n}^{\rm ideal}(\theta)$ $=\frac{1}{n}\sum_{i=1}^{n}\ell(t_{\theta},g^{\star}_{\theta};w_{i},z_{i}).$ (7)
Indeed $g^{\star}_{\theta^{\star}}=\Psi^{\star}$ is the true link function, but away from $\theta^{\star}$ note that $g_{\theta}$ may not even be monotonic. Fortunately, near $\theta^{\star}$, the difference between the true likelihood $\ell(t_{\theta},\Psi^{\star};w,z)$ and the quasi-likelihood $\ell(t_{\theta},g^{\star}_{\theta};w,z)$ vanishes very quickly as we get closer to $\theta^{\star}$, so that, under certain regularity, the estimator that maximizes $L_{n}^{\rm ideal}(\theta)$ behaves as though we knew the link function exactly. For example, if the link were logistic, the first order behavior would be the same for this estimator as for the one optimizing the well-specified DPO objective. Of course, the latter would no longer even be consistent when the link is no longer logistic. The former would adapt to whatever the link may be and behave as though we knew it exactly.
4.2 The OSPO policy learner with nuisance plug-in
Unfortunately, we do not know $g^{\star}_{\theta}$ in Eq. 7. Fortunately, we can easily estimate $g_{\theta}$ as a univariate regression of $z$ on $t_{\theta}$. Suppose we are given an estimator $\hat{g}_{\theta}$ for each $\theta\in\Theta$. Then we can define the plug-in quasi-likelihood as follows and its maximizer as the OSPO learner:
$L_{n}^{\rm OSPO}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(t_{\theta},\hat{g}_{\theta};w_{i},z_{i}),\quad\hat{\theta}_{\rm OSPO}\in\operatornamewithlimits{argmax}_{\theta\in\Theta}L_{n}^{\rm OSPO}(\theta).$ (8)
One possible choice for $\hat{g}_{\theta}$ is a kernel regression estimator: given a dataset $\mathcal{D}$, set
$\hat{g}_{\theta}(u)=\frac{\sum_{(w,z)\in\mathcal{D}}K\big((u-t_{\theta}(w))/b_{\theta}\big)\,z}{\sum_{(w,z)\in\mathcal{D}}K\big((u-t_{\theta}(w))/b_{\theta}\big)},$ (9)
where $K$ is a kernel such as Gaussian $K(u)=\exp(-\|u\|^{2}/2)$ or Epanechnikov $K(u)=\max\{1-u^{2},0\}$ and $b_{\theta}\to 0$ is a bandwidth. One can choose $b_{\theta}=h$ to be a common bandwidth. However, we will instead use $b_{\theta}=h\,\hat{\sigma}(t_{\theta}(w))$ to be scaled by the empirical standard deviation of $t_{\theta}$ on $\mathcal{D}$ so that $\hat{g}_{\theta}$ is scale invariant. Indeed, $g^{\star}_{\theta}$ is invariant to scale in that $g_{\theta}^{\star}(u)=g_{\theta^{\prime}}^{\star}(\beta u)$ whenever $t_{\theta}=\beta t_{\theta^{\prime}}$ for $\beta>0$. This is because $g_{\theta}^{\star}$ is a conditional expectation and scaling does not change the information content of the conditioning set.
The dataset $\mathcal{D}$ can be the same $n$ datapoints used in Eq. 8. Klein and Spady (1993) use a leave-one-out version where $\mathcal{D}=\{(w_{j},z_{j}):j\neq i\}$ is used to train the kernel regression plugged into the $i^{\text{th}}$ term in Eq. 8. That is the practical version we recommend. However, since the theoretical analysis only derives rates, we will focus the analysis on using a separate independent dataset, as splitting the data would only affect constants in the bound.
4.3 Identification up to scale and sign
There is a small wrinkle: if $t^{\star}$ is a function of $t_{\theta}$ then Eq. 7 cannot distinguish between them, as $g_{\theta}(t_{\theta})=g_{\theta^{\star}}(t_{\theta}^{\star})$ will be identical and so will the quasi-likelihood above. Klein and Spady (1993) get around this and achieve efficiency for $\theta^{\star}$ by requiring a very strong identification condition: that $g_{\theta}(t_{\theta})=g_{\theta^{\star}}(t_{\theta}^{\star})$ must imply $\theta=\theta^{\star}$. This can be unreasonably strong as it essentially means we know the sign in advance and precludes classes closed under negation ($t^{\star}=-t_{\theta}$ would imply $\theta=\theta^{\star}$ which would then imply $t^{\star}=0$). This is in contrast to PSPO’s objective Eq. 6, which restricts the profiling to monotonic nondecreasing link functions so that negating the index would not preserve the profile likelihood objective.
Focusing on identifying policies rather than underlying parameters, we will instead impose a slight strengthening of our up-to-scaling identification condition (Assumption 4) so that near-observational-equivalence translates to nearly pointing in the same (or opposite) directions. Quantifying the level of “near” equivalence is needed to ensure sufficient curvature in the quasi-likelihood.
Assumption 5 (MSE identification margin).
There exists a constant $c_{\rm MSE}>0$ with
$\inf_{f:\mathbb{R}\to\mathbb{R}}\|t^{\star}-f(t_{\theta})\|\geq c_{\rm MSE}\inf_{f\in\mathcal{F}_{\rm lin}}\|t^{\star}-f(t_{\theta})\|\qquad\forall\theta\in\Theta,$
where $\mathcal{F}_{\rm lin}=\{u\mapsto\alpha+\beta u:\alpha,\beta\in\mathbb{R}\}$.
Note that the left-hand side can also be written as $\mathbb{E}\bigl[\mathrm{Var}(t^{\star}(w)\mid t_{\theta}(w))\bigr]$. We can also rescale by $1/\mathrm{Var}(t^{\star}(w)$ to express the above as a relationship between a nonparametric $R^{2}$ and the linear $R^{2}$ (in the population). Assumption 5 posits that if $t_{\theta}$ contains nearly as much information as $t^{\star}$, then the linear regression error needs to be small. Since $\mathcal{F}_{\rm lin}$ permits negative coefficients, $t_{\theta}$ may point in the opposite direction from $t^{\star}$. Note that, under the Assumption 3’s coverage of $\pi_{\rm ref}$, Assumption 5 implies Assumption 4.
We now show that this means that nearly optimizing the quasi-likelihood at the population level means nearly aligning with $\pm t^{\star}$. This curvature in the quasi-likelihood is, in particular, needed for establishing a rate, as we do later in this section. We next introduce some additional regularity that various derivatives exist and are bounded, ensuring the quasi-likelihood is sufficiently smooth.
Assumption 6 (Additional regularity).
Suppose $\Psi^{\star}$ has a derivative bounded away from 0 on [M,M][-M,M]. Suppose $g_{\theta}^{\star}$ is twice continuously differentiable on [M,M][-M,M] with uniformly bounded first and second derivatives on [M,M][-M,M] over all $\theta\in\Theta$.
Lemma 1.
Under Assumptions 1, 2, 3, 5 and 6, there exists a constant $\lambda_{0}>0$ such that
$\mathbb{E}L_{n}^{\rm ideal}(\theta^{\star})-\mathbb{E}L_{n}^{\rm ideal}(\theta)\geq\lambda_{0}\inf_{f\in\mathcal{F}_{\rm lin}}\|t^{\star}-f(t_{\theta})\|^{2}\qquad\forall\theta\in\Theta.$
4.4 Identifying the sign
We focus on permitting flexibility in the parametrization and identify only policies, which are invariant to the index’s scale and $x$-based location. However, Assumption 5 still allows flipping the direction of the index while maintaining its informational content, since the acute angle will remain the same. There is a simple fix: ensure marginal sign agreement with $z$. As long as we are sufficiently well-aligned with $\pm t^{\star}$, even with the wrong sign, we can ensure that flipping the sign to align with $z$ will maintain error magnitude while fixing signs.
Lemma 2 (Fixing index sign).
Suppose Assumptions 1, 2, 3, 5 and 6 hold and that $t^{\star}\neq 0$. Then $\mathbb{E}[t^{\star}(w)z]>0$. Fix any $\theta\in\Theta$ with $t_{\theta}\neq 0$ and $\mathbb{E}L_{n}^{\rm ideal}(\theta^{\star})-\mathbb{E}L_{n}^{\rm ideal}(\theta)\leq\frac{c_{\rm MSE}}{2}\mathbb{E}[t^{\star}(w)z]$. Define $s=\mathbb{E}[t_{\theta}t^{\star}]$. Given a dataset $\mathcal{D}$ of iid draws of $(w,z)$, let $\hat{s}=\frac{1}{\left\lvert\mathcal{D}\right\rvert}\sum_{(w,z)\in\mathcal{D}}t_{\theta}(w_{i})z_{i}$. Then,
$\mathbb{P}\left(\hat{s}\,s\;\leq\;0\right)\leq\frac{4}{\left\lvert\mathcal{D}\right\rvert}\frac{\mathbb{E}[t^{\star}(w)^{2}]}{\mathbb{E}[t^{\star}(w)z]^{2}}.$
If we additionally assume a smoothness condition $\|t_{\theta}\|_{\infty}/\|t_{\theta}\|_{2}\leq c_{\rm smooth}$, the decay rate can be made exponential in $\left\lvert\mathcal{D}\right\rvert$:
$\mathbb{P}\left(\hat{s}\,s\;\leq\;0\right)\leq\exp(-\left\lvert\mathcal{D}\right\rvert\mathbb{E}[t^{\star}z]^{2}/(8c_{\rm smooth}\mathbb{E}[t^{\star}(w)^{2}])).$
The smoothness condition is analogous (although, incomparable) to the compactness condition in Theorem 4.
This means that it suffices to learn $t$ to align along the direction of $\pm t^{\star}$. We can always fix the sign afterward for any one given $t$. Notice that importantly both the minimum necessary alignment and the constants on the right-hand side above do not depend on the particular $t$. In practice, we use the same data to both train $\hat{\theta}^{\rm OSPO}$ and to align the sign. In theory, we simply split the data and fix the sign after learning some $\hat{t}$ since the rate is unchanged and this avoids having to impose more complexity assumptions on $\mathcal{T}$ so as to be able to prove Lemma 2 uniformly over $\mathcal{T}$.
4.5 Policy error for generic plug-in
We now analyze the policy error incurred by the OSPO learner, given a generic plug-in $\hat{g}_{\theta}$ in terms of its own error. Crucially, we will show that the error in $\hat{g}_{\theta}$ enters into our final bound quadratically, meaning it suffices to learn $g^{\star}_{\theta}$ quite slowly and it would still not affect the final rate. In this sense, the OSPO learner is quite robust to learning $\hat{g}_{\theta}$.
We need to assume some regularity for the plug-in $\hat{g}_{\theta}$: that is has almost the same range as $g^{\star}_{\theta}$ and certain Lipschitzness locally around $t^{\star}$.
Assumption 7 (Plug-in regularity).
$\inf_{u\in[-M,M]}\hat{g}_{\theta}(u)\geq c_{\rm range}\inf_{u\in[-M,M]}g^{\star}_{\theta}(u)$ and $\inf_{u\in[-M,M]}(1-\hat{g}_{\theta}(u))\geq c_{\rm range}\inf_{u\in[-M,M]}(1-g^{\star}_{\theta}(u))$ for some $c_{\rm range}>0$. There exists $\iota>0$ such that for any $\theta\in\Theta$ with $\inf_{a\in\mathbb{R}}\|t^{\star}-at_{\theta}\|\leq\iota$, we have $\|\hat{g}_{t^{\star}}(t_{t^{\star}}(w))-\hat{g}_{\theta}(t_{\theta}(w))\|_{4}\leq c_{\rm Lip}\inf_{a\in\mathbb{R}}\|t^{\star}-at_{\theta}\|$, $\|\hat{g}_{t^{\star}}(t_{t^{\star}}(w))-g^{\star}_{t^{\star}}(t_{t^{\star}}(w))-\hat{g}_{\theta}(t_{\theta}(w))+g^{\star}_{\theta}(t_{\theta}(w))\|_{4}\leq c_{\rm Lip}(\|\hat{g}_{t^{\star}}(t_{t^{\star}}(w))-g^{\star}_{t^{\star}}(t_{t^{\star}}(w))\|_{4}\vee\|\hat{g}_{\theta}(t_{\theta}(w))-g^{\star}_{\theta}(t_{\theta}(w))\|_{4})\inf_{a\in\mathbb{R}}\|t^{\star}-at_{\theta}\|$ for some $c_{\rm Lip}>0,c^{\prime}_{\rm Lip}>0$.
Note that using the scale-invariant distance $\inf_{a\in\mathbb{R}}\|t^{\star}-at_{\theta}\|$ on the right-hand side of the first Lipschitz equation makes sense because $g^{\star}$ is scale invariant: for any $\theta,\theta^{\prime}\in\Theta$ with $t_{\theta}=\beta t_{\theta^{\prime}}$ for $\beta>0$, we have $g^{\star}_{\theta}(u)=g^{\star}_{\theta^{\prime}}(\beta u)$. The same holds for the second Lipschitz equation if $\hat{g}$ is similarly scale invariant (such as is the case for the kernel regression in Eq. 9).
We also need to control the complexity of the index class $\mathcal{T}$. We do so in terms of rates on the covering number $N_{2}(\epsilon,\mathcal{T})$, the minimal number of square-integrable functions $t_{1},\dots,t_{N}$ needed to obtain $\sup_{t\in\mathcal{T}}\min_{i=1,\dots,N}\|t-t_{i}\|\leq\epsilon$.
Assumption 8 (Complexity of indices).
Suppose there exist constants $p\geq 0,A>0$ such that, if $p>0$ we have $\log N_{2}(\epsilon,\mathcal{T})\leq A\epsilon^{-p}$ and if $p=0$ we have $N_{2}(\epsilon,\mathcal{T})\leq(A/\epsilon)^{v}$ for some constant $v>0$.
This neatly packages two cases: the truly nonparametric case $p>0$ and the “simple” case $p=0$. In particular, if $\mathcal{T}$ were a VC-subgraph class of dimension $v$ then we would have Assumption 8 with $p=0$ and the given $v$ and some universal constant $A$ (van der Vaart and Wellner, 1996, theorem 2.6.7). In the appendix, the proof only uses Assumption 8 to bound a critical radius (Wainwright, 2019), so the assumption can be supplanted by bounds thereon when available instead.
We can now state our result.
Theorem 5 (Nonparametric rate for OSPO).
Suppose Assumptions 1, 2, 3, 5, 6 and 7 hold. Let $\hat{s}=\frac{1}{\left\lvert\mathcal{D}\right\rvert}\sum_{(w,z)\in\mathcal{D}}t_{\hat{\theta}_{\rm OSPO}}(w_{i})z_{i}$ be estimated on an independent data set with $\left\lvert\mathcal{D}\right\rvert\to\infty$. Suppose also $\sup_{\theta\in\Theta}\|\hat{g}_{\theta}-g^{\star}_{\theta}\|_{4}^{2}\to_{p}0$. Then for a constant $c>0$ depending on those in the assumptions, for all $\delta\in(0,0.5)$ and sufficiently large $n$, with probability at least $1-\delta$,
$\rho(\hat{s}\,h_{\hat{\theta}_{\rm OSPO}})\;\leq\;c\left(\varrho_{n}+\sqrt{\frac{\log(1/\delta)}{n}}+\sup_{\theta\in\Theta}\|\hat{g}_{\theta}-g^{\star}_{\theta}\|_{4}^{2}\right),\qquad\varrho_{n}=\left\{\begin{array}[]{lr}\sqrt{{v\log(n)}/{n}}&p=0\\ n^{-1/(2+p)}&\quad 0<p<2\\ n^{-1/4}\sqrt{\log n}&p=2\\ n^{-1/(2p)}&p>2\end{array}\right$.
The style of our analysis is similar to the approach of Foster and Syrgkanis (2023) for leveraging orthogonality of loss functions to nuisances in statistical learning, but we derive all bounds (except from the empirical process bounds) from first principles to avoid too stringent conditions since we have to deal with a parameter-dependent nuisance function.
Crucially, the dependence on the error in $\hat{g}_{\theta}$ is quadratic, meaning if we learn it a rate slightly faster than $n^{-1/4}$, or even just faster than $\sqrt{\varrho_{n}}$, then it will not affect the final rate at all. This is incredibly lax considering that $\hat{g}_{\theta}$ is just a univariate regression. Therefore, as we show next, under very mild regularity, even kernel regression can achieve a rate fast enough that it does not affect the first order behavior of the bound above.
4.6 Kernel regression plug-in error
Theorem 5 depended on having a plug-in estimator $\hat{g}$ with small uniform $L_{4}$ error. Any plug-in with such a guarantee can work. One proposal was the kernel regression estimator in Eq. 9 with scale-free bandwidth $b_{\theta}=h\,\hat{\sigma}(t_{\theta}(w))$, which we will analyze here. We show it attains a uniform $L_{\infty}$ (thus also uniform $L_{4}$) error rate of $m^{-1/3}$ under mild conditions, where we let $m=\left\lvert\mathcal{D}\right\rvert$ be the size of the data set used to fit the kernel regression. Since this error enters the bound in Theorem 5 quadratically, this means it has no influence on the rate of the bound and is eventually insignificant.
We will need some regularity conditions to establish this result. First, we impose conditions on the choice of kernel and bandwidth. These will be satisfied, e.g., if we use the Epanechnikov kernel and let $h\asymp(\log m/m)^{1/3}$.
Assumption 9 (Kernel conditions).
The kernel $K:[-1,1]\to\mathbb{R}_{+}$ is nonnegative, compactly supported, symmetric, bounded, and Lipschitz. It has $\int K(u)du=1$, $\int\left\lvert u\right\rvert K(u)du<\infty$, and $\int\left\lvert u\right\rvert K(u)^{2}du<\infty$. The bandwidth satisfies $\leq 1, h0h\to 0$ and $mh\to\infty$.
Next, we need some regularity, including a complexity condition on $\mathcal{T}$ that is more specific than the critical radius used in the last section.
Assumption 10 (Regularity for uniform kernel regression).
$g^{\star}_{\theta}(u)$ is equi-Lipschitz over $\theta\in\Theta$. $t_{\theta}(w)$ has a density bounded away from 0 and $\infty$ on [M,M][-M,M] with common bounds over $\theta\in\Theta$. For some $v\geq 1,A>0$, the minimum number of $\|\cdot\|_{\infty}$ balls needed to cover $\mathcal{T}$ satisfies $N_{\infty}(\epsilon,\mathcal{T})\leq(A/\epsilon)^{v}$ for all $\epsilon\in(0,1]$.
Theorem 6.
Suppose Assumptions 3, 9 and 10 hold. Then there exist constants $c_{0},c_{1},c_{2},c_{3}>0$ such that for every $\delta\in(0,0.5)$, whenever $mh\geq c_{0}(v\log(c_{3}m/h)+\log(1/\delta))$, we have that with probability at least $1-\delta$,
$\sup_{\theta\in\Theta}\|\hat{g}_{\theta}-g^{\star}_{\theta}\|_{\infty}\leq c_{1}h+c_{2}\sqrt{\frac{v\log(c_{3}m/h)+\log(1/\delta)}{mh}}+c_{2}\frac{v\log(c_{3}m/h)+\log(1/\delta)}{mh}.$
Consequently, if we choose $h\asymp(\log m/m)^{1/3}$, we obtain a rate of $\tilde{O}(m^{-1/3})$. If $m\asymp n$ (e.g., evenly split the data) this contributes an insignificant $\tilde{O}(m^{-2/3})=o(n^{-1/2})$ term to Theorem 5.
4.7 First-order optimization
Theoretically, to obtain the aligned policy, we must split the data four ways: one quarter to learn $\hat{g}_{\theta}$ for all $\theta\in\Theta$, one quarter to construct the quasi-likelihood and fit $\hat{\theta}^{\rm OSPO}$ to maximize it, one quarter to learn the sign of the index using $\hat{s}$, and finally one quarter to learn $\hat{\beta}_{\hat{\theta}^{\rm OSPO}}$ to calibrate to the divergence constraint (see Section 6). In practice, we suggest to just reuse the whole data for each of the latter steps, since the index sign and divergence calibration are single dimensional and should be quite stable (and the latter even only depends on $x$-data).
[…]
5 Semiparametric RLHF Based on Bipartite Ranking (RSPO)
Both PSPO and OSPO were based on maximizing something akin to a likelihood. We now consider an alternative based on optimizing a ranking objective. Han (1987) proposed a max rank correlation approach for learning an identifiable linear index in a family of observation models that includes binary choice, in which case the estimator maximizes the area under the curve (AUC). Sherman (1993) showed the estimated linear coefficients converge at a $n^{-1/2}$ rate and are asymptotically normally distributed. However, in the single-index binary choice setting, the asymptotic variance is generally larger than the efficiency bound derived by Cosslett (1983) and achieved by Klein and Spady (1993), meaning the max rank correlation estimator is not efficient and does not behave asymptotically like maximum likelihood with the true link function, as if it were actually known.
Nonetheless, a ranking objective for RLHF is appealing as it involves no additional nuisances (like PSPO) while still attaining a rate (like OSPO). In this section, we study how to leverage it for our alignment problem with nonlinear index classes implied by flexible policies, characterizing policy error in terms of general function complexity. To do so, we first establish an identification margin similar to Lemma 1, showing that small excess AUC translates to learning the index up to scale. Then we can leverage empirical risk minimization results of $U$-statistics (Clémençon et al., 2008) to establish a bound on the excess AUC for nonparametric index classes. Chaining the two together with Theorem 2, we obtain policy convergence.
5.1 From AUC maximization to an index and a policy
Define the AUC as
$\mathrm{AUC}(t)=$ $~\mathbb{P}_{(w_{0},z_{0}),(w_{1},z_{1})\sim P^{2}}(t(w_{1})>t(w_{0})\mid z_{0}=0,z_{1}=1)$ (10)
$+\frac{1}{2}\mathbb{P}_{(w_{0},z_{0}),(w_{1},z_{1})\sim P^{2}}(t(w_{1})=t(w_{0})\mid z_{0}=0,z_{1}=1),$
where $(w_{0},z_{0}),(w_{1},z_{1})$ are two independent draws from $P$. This is another way to express the area under the receiver operating characteristic curve given by thresholding $t(w)$ to classify $z$ (Menon and Williamson, 2016).
Notice this is distinct from averaging over single demonstration pairs $(x,y_{0},y_{1})$ how often the ranking between $h_{\theta}(x,y_{1})$ and $h_{\theta}(x,y_{0})$ matches the label $z$. This amounts to a 0-1 loss for the sign of $t_{\theta}(w)$ in classifying the label $z$, maximizing which corresponds to max-score estimators (discussed in Section 7.2). Importantly here we average over pairs of pairs.
In particular, taking the pairs-of-pair point of view further, notice that, since $P$ is symmetric in $y_{0},y_{1}$, if $\Phi^{\star}$ were also symmetric ($\Phi^{\star}(u)=1-\Phi^{\star}(-u)$), then we can rewrite Eq. 10 for our alignment-from-preferences setting without any conditioning:
$\mathrm{AUC}(t)=$ $~\mathbb{P}_{(x_{0},y_{00},y_{01},z_{0}),(x_{1},y_{10},y_{11},z_{1})\sim P^{2}}(t(x_{1},y_{1,(1-z_{1})},y_{1,z_{1}})>t(x_{0},y_{0,z_{0}},y_{0,1-z_{0}}))$ (11)
$+\frac{1}{2}\mathbb{P}_{(x_{0},y_{00},y_{01},z_{0}),(x_{1},y_{10},y_{11},z_{1})\sim P^{2}}(t(x_{1},y_{1,(1-z_{1})},y_{1,z_{1}})=t(x_{0},y_{0,z_{0}},y_{0,1-z_{0}})).$
Essentially, to get a draw from the conditional on $z=1$, take any observation $(x,y_{0},y_{1},z)$ and just make sure that the winning demonstration of the two per the observed label $z$ is placed in the $y_{1}$ position, and vice versa for $z=0$.
We next show that, in our setting, the set of maximizers of $\mathrm{AUC}(t)$ over $t\in\mathcal{T}$ are exactly $t^{\star}$, up to scale.
Proposition 2 (Maximizers of AUC are $t^{\star}$ up to scale).
$\mathrm{AUC}(t^{\star})\geq\mathrm{AUC}(t)$ for all indices $t:\mathcal{W}\to\mathbb{R}$, with equality if and only if $t^{\star}(w)=m(t(w))$ almost surely for a monotone nondecreasing function $m:\mathbb{R}\to\mathbb{R}$. Consequently, under Assumption 4, for any $t\in\mathcal{T}$, we have equality if and only if $t^{\star}(w)=at(w)$ almost surely for some $a\geq 0$.
In order to obtain rates, however, we need to show that small AUC suboptimality corresponds to being nearly equal to $t^{\star}$ up to scale, which correspondingly controls policy error, based on our results. We will again leverage Assumption 5 to quantify near-observational equivalence. We will also need some additional regularity similar but distinct from Assumption 6.
Assumption 11 (Additional regularity, alt.).
Suppose $\Psi^{\star}$ has a derivative bounded away from 0 and $\infty$ on [M,M][-M,M]. Suppose $t^{\star}$ is supported on an interval, on which it has a density bounded away from zero and $\infty$.
The latter part of Assumption 11 regarding the density of $t^{\star}$ will be used both to establish the next result this section as well as to establish a noise (margin) condition in the next section (for which only the upper bound is needed).
Theorem 7 (AUC suboptimality of policy-implied index controls policy error).
Suppose Assumptions 1, 2, 3, 5 and 11 hold. Then there exist constants $\iota>0,\lambda_{0}>0$ such that for every $\theta\in\Theta$ satisfying $\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t_{\theta})\leq\iota$, we have
$\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t_{\theta})\geq\lambda_{0}\inf_{a\geq 0}\|t^{\star}-at_{\theta}\|^{2}.$
5.2 The RSPO policy learner
Given Theorem 7, to align the policy to be quite close to the optimal policy, it suffices to find a policy whose implied index function attains near-optimal AUC. A reasonable approach to doing so is to maximize the empirical AUC, giving rise to the Ranking SPO (RSPO) learner:
$\widehat{AUC}(\theta)=\frac{1}{n_{1}n_{0}}\sum_{i,j:z_{i}=0,z_{j}=1}\phi(t_{\theta}(w_{j})-t_{\theta}(w_{i})),\qquad\hat{\theta}_{\rm RSPO}\in\operatornamewithlimits{argmax}_{\theta\in\Theta}\widehat{AUC}(\theta).$ (12)
where $\phi(u)=\mathbb{I}\left[u>0\right]+\tfrac{1}{2}\mathbb{I}\left[u=0\right]$ and $n_{1}=\sum_{i=1}^{n}z_{i}$, $n_{0}=n-n_{1}$.
Under the assumption that $\Phi^{\star}$ is symmetric, we can alternatively use an empirical version of Eq. 11:
$\widetilde{AUC}(\theta)=\frac{1}{n^{2}}\sum_{i,j}\phi(t_{\theta}(x_{j},y_{j,1-z_{j}},y_{j,z_{j}})-t_{\theta}(x_{i},y_{i,z_{i}},y_{i,1-z_{i}})),\qquad\tilde{\theta}_{\rm RSPO}\in\operatornamewithlimits{argmax}_{\theta\in\Theta}\widetilde{AUC}(\theta).$ (13)
Note that while Eq. 10 and Eq. 11 are exactly equal under the assumption that $\Phi^{\star}$ is symmetric, their sample versions (Eq. 12 and Eq. 13, respectively) are generally distinct. In practice, since the symmetric assumption is quite reasonable, we suggest to use Eq. 13 as it essentially gets double the data for free. In theory, in order to easily leverage existing results (Clémençon et al., 2008) and since the rates would be unaffected, we analyze Eq. 12.
Since the AUC objective involves a step function, to appropriately control the complexity we need to consider the sets of pairs that ranked in order or anti-order. To measure the complexity of a class $\mathcal{S}$ of subsets $S\subset\mathcal{W}\times\mathcal{W}$, define $N_{\triangle}(\epsilon,\mathcal{S})$ as the smallest number of subsets $S_{1},\dots,S_{N}\subset\mathcal{W}\times\mathcal{W}$ we need to take so that for every $S\in\mathcal{S}$ there exists a subset $S_{i}$ where $\mathbb{P}_{w,w^{\prime}\sim P^{2}}((w,w^{\prime})\in S,(w,w^{\prime})\notin S_{i}\vee(w,w^{\prime})\notin S,(w,w^{\prime})\in S_{i})\leq\epsilon$, that is, the symmetric difference has measure at most $\epsilon$.
Assumption 12 (Complexity of ranking sets).
Define $\mathcal{R}_{\mathcal{T}}=\{\{(w,w^{\prime}):t(w)-t(w^{\prime})\geq 0\}:t\in\mathcal{T}\}$. Suppose there exist constants $p\geq 0,A>0$ such that, if $p>0$ we have $\log N_{\triangle}(\epsilon,\mathcal{R}_{\mathcal{T}})\leq A\epsilon^{-p}$ and if $p=0$ we have $N_{\triangle}(\epsilon,\mathcal{R}_{\mathcal{T}})\leq(A/\epsilon)^{v}$ for some constant $v>0$.
This again neatly packages two cases: the truly nonparametric case $p>0$ and the “simple” case $p=0$. In particular, if $\mathcal{R}_{\mathcal{T}}$ were a VC class of dimension $v$ then we would have Assumption 12 with $p=0$ and the given $v$ and some universal constant $A$ (van der Vaart and Wellner, 1996, theorem 2.6.4).
We can now establish our main result for RSPO, where we permit flexible indices under Assumption 12.
Theorem 8.
Suppose Assumptions 1, 2, 3, 5, 11 and 12 hold. Then for a constant $c>0$ depending on all the assumptions, for all $\delta\in(0,0.5)$ and sufficiently large $n$, with probability at least $1-\delta$,
$\rho(h_{\hat{\theta}_{\rm RSPO}})\,\leq\,c\;\sqrt{\log(1/\delta)}\;\left\{\begin{array}[]{lr}\sqrt{{v\log(n)}/{n}}&p=0\\ n^{-1/(2+2p)}&\qquad\qquad 0<p<1\\ n^{-1/4}\sqrt{\log n}&p=1\\ n^{-1/4p}&p>1\end{array}\right$.
Importantly, we are able to attain policy error rates without having to estimate a nuisance parameter as in OSPO. The rates are generally incomparable, but we can sometimes obtain control on $N_{\triangle}(\epsilon,\mathcal{R}_{\mathcal{T}})$ via control on $N_{q}(\epsilon,\mathcal{T})$, the minimal number of $\|\cdot\|_{q}$ balls needed to cover $\mathcal{T}$.
Lemma 3.
Suppose for some $c_{\rm margin}>0$ and $\alpha\in(0,1]$ we have $\mathbb{P}_{w,w^{\prime}\sim P^{2}}(\left\lvert t(w)-t(w^{\prime})\right\rvert\leq u)\leq c_{\rm margin}u^{\alpha}$ for each $t\in\mathcal{T}$. Then for each $q\in[1,\infty]$, for a constant $c_{q}>0$, we have $N_{\triangle}(\epsilon,\mathcal{R}_{\mathcal{T}})\leq N_{q}((\epsilon/c_{q})^{1/\beta},\mathcal{T})$ with $\beta=\frac{\alpha q}{\alpha+q}$ for $q<\infty$ and otherwise $\beta=\alpha$.
If we take Lemma 3 with $\alpha=1$ (e.g., $t(w)$ has bounded densities) and assume Assumption 8 with exponent $p$, we obtain Assumption 12 with exponent $3/2p$. Thus, we maintain $\tilde{O}(1/\sqrt{n})$ rates in the “simple” case, but this can quickly deteriorate. For example, if $p=2$, OSPO would attain a rate of $\tilde{O}(n^{-1/4})$ compared to a much worse rate of $\tilde{O}(n^{-1/12})$ for RSPO. Nonetheless, the rates are indeed generally incomparable, and RSPO avoids nuisance estimation, circumventing the link function altogether by sidestepping the likelihood.
5.3 First-order optimization
The empirical AUC objective in Eq. 12 is somewhat difficult to optimize since $\phi$ is a step function. One approach, in the spirit of Horowitz (1992), is to replace $\phi(u)$ by $F(u/h)$ where $F$ is a smooth symmetric CDF and $h\to 0$ is a bandwidth. (Note that unlike the setting studied by Horowitz (1992), the smoothing is not needed for statistical reasons since the population AUC objective is already sufficiently smooth, unlike the objective they study; see Section 7.2.) This would make the objective differentiable so that we can use any off-the-shelf first-order optimizer, but the objective would still be non-convex in $t$.
A more appealing alternative is to use a surrogate loss function for bipartite ranking (Agarwal, 2014). In particular, we can obtain one from any surrogate loss for classification (Bartlett et al., 2006) by transforming the data into classification data. Let us focus on transforming Eq. 13, as that is the variant we recommend in practice; the transformation is analogous for Eq. 12. Create the $2n^{2}$ data points, where for each $i,j$, we label the features $(x_{j},y_{j,1-z_{j}},y_{j,z_{j}},x_{i},y_{i,z_{i}},y_{i,1-z_{i}})$ with $1$ and the features $(x_{j},y_{j,z_{j}},y_{j,1-z_{j}},x_{i},y_{i,1-z_{i}},y_{i,z_{i}})$ with 0. Then Eq. 13 is given by $2$ times the 0-1 loss of this classification problem over class scores $(w,w^{\prime})\mapsto t(w)-t(w^{\prime})$. Now use any convex classification surrogate loss, such as hinge or logistic. Now, rather than taking batches over the $n^{2}$ augmented data points, we can take batches of size $B$ from the original $n$ data points and simply consider the gradient for the corresponding $B^{2}$ terms.
6 Empirical Recalibration to the Divergence Budget
After training $\hat{\theta}$, we must choose a temperature $\beta$ that enforces the divergence budget. So far we wrapped this into the population-calibrated policy $\pi_{\kappa,\theta}$, but this cannot be done in practice. We now discuss how to calibrate the divergence using a validation sample of contexts $x$, and we show how this leads to an additive policy error that we can simply add on top of our previous bounds on the error in $\pi_{\kappa,\hat{\theta}}$ for a learner $\hat{\theta}$.
The procedure solves a one-dimensional root-finding problem on held-out contexts. We show a uniform law of large numbers in the temperature, a root-$m$ calibration rate, and the ensuing policy error guarantee. While the results require a separate validation set, so that we can treat $\hat{\theta}$ as a fixed input, in practice we may calibrate on the same data (using only the contexts).
Fix a compact interval $\mathcal{B}=[\underline{\beta},\overline{\beta}]\subset(0,\infty)$. Let $x_{1},\dots,x_{m}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}P_{x}$ and define
$\hat{\Phi}_{m}(\beta,\theta)=\frac{1}{m}\sum_{i=1}^{m}D_{f}\!\bigl(\pi_{\beta,h_{\theta}}(\cdot\mid x_{i})\,\big\|\,\pi_{\mathrm{ref}}(\cdot\mid x_{i})\bigr),\qquad\hat{\beta}_{\theta}\in\operatornamewithlimits{argmin}_{\beta\in\mathcal{B}}\bigl|\hat{\Phi}_{m}(\beta,\theta)-\kappa\bigr|.$
Theorem 9 (Empirical $\beta$-calibration).
Suppose the conditions of Theorem 2 hold. Then for some constant $c>0$, we have that for any $\theta\in\Theta$ and $\delta\in(0,0.5)$, with probability at least $1-\delta$,
$\|\pi_{\hat{\beta}_{\theta},\theta}-\pi_{\kappa,\theta}\|_{1}\leq\frac{c}{v}\sqrt{\frac{\log(m/\delta)}{m}},$
provided $v=\mathbb{E}_{x}\operatorname{Var}_{y\sim\pi_{\mathrm{ref}}(\cdot\mid x)}(h_{\theta}(x,y))>0$ (which implies $\beta_{\kappa,\theta}$ is positive and unique) and $\beta_{\kappa,\theta}\in\mathcal{B}$.
Therefore, we can simply add this additional $\tilde{O}(1/\sqrt{m})$ term to whatever policy error we derived on the population-calibrated policy by bounding $\rho(\theta)$ via Theorem 2. In theory, if we split the $n$ data points and use a constant fraction for calibration, we will never deteriorate the rate. In practice, we can use the same data since calibration should be rather stable and only ever uses the $x$-data.
7 Extensions
7.1 Localizing $\hat{g}_{\theta}$ in OSPO
OSPO as posited in Eq. 8 requires we estimate $\hat{g}_{\theta}$ for each $\theta\in\Theta$, and moreover the bound in Theorem 5 depends on its uniform convergence over $\theta\in\Theta$. One way to avoid this would be to localize the link estimate around an initial $\theta$ estimator in the style of Kallus et al. (2019). That is, first estimate an initial $\hat{\theta}$ (perhaps using PSPO), then estimate $\hat{\Psi}$ as the univariate regression of $z$ on $t_{\hat{\theta}}(w)$, then maximize the plug-in quasi-likelihood $\frac{1}{n}\sum_{i=1}^{n}\ell(t_{\theta},\hat{\Psi};w_{i},z_{i})$.
This has the benefit that in the end what we optimize looks like a standard classification loss with a fixed link function, as in DPO. While $\hat{\Psi}$ may not be monotonic, we only get a better estimate via rearrangement (Chernozhukov et al., 2009), ensuring monotonicity. It may, however, not be so smooth. Additionally, it may be challenging to ensure fast enough convergence of the initial $\hat{\theta}$ so that we get a fast enough rate on $\hat{\Psi}$, since our current analysis for PSPO gives no rate. Finally, the analysis will require adapting the approach of Kallus et al. (2019), which is developed for point estimation and inference, to orthogonal statistical learning in the vein of Foster and Syrgkanis (2023). This therefore remains a promising direction of future work.
7.2 Max-score estimators
[…]
This is the model studied by Manski (1975, 1985); Kim and Pollard (1990); Horowitz (1992) for linear indices $\mathcal{T}=\{w\mapsto\beta^{\top}w\}$. Indeed, in this model, any $t$ with $\mathbb{I}\left[t(w)\geq 0\right]=\mathbb{I}\left[t^{\star}(w)\geq 0\right]$ is observationally equivalent to $t^{\star}$. Therefore, it is generally impossible to identify up to scale anything much more complex than a linear index. That is, requiring that sign agreement of $t,t^{\prime}$ implies $t\propto t^{\prime}$ is highly restrictive. It, for example, prevents any class that admits “bump” functions, such as Hölder or Sobolev classes, since they would maintain sign while changing shape and scale. In contrast, the identification Assumption 4 is much more lax and can admit such nonparametric classes. For example, for any atomless CDF $F_{0}$, any subset of $\{af:a>0,f(w)\sim F_{0}\}$ (such as further restricting $f$ in some nonparametric function class) satisfies Assumption 4. Fixing the marginal distribution thus is exactly what is done with normalizing flows / invertible networks.
For learning linear indices up to scaling, Kim and Pollard (1990) establish $n^{-1/3}$ rates for the normalized coefficients that define the half-space that minimizes the binary 0-1 classification loss (the so-called max-score estimator). Horowitz (1992) replaces the 0-1 step function with a smooth CDF with a vanishing bandwidth (the so-called smoothed max-score estimator) and establishes a $n^{-s/(2s+1)}$ rate when $\Psi_{w}^{\star}$ is $s$-times Hölder-continuously differentiable.
These results can be directly applied to learning indices implied by realizability via Proposition 1 up to scale (and chained with Theorem 2 for policy error), provided they are linear. Although identification up to scale (and therefore of divergence-constrained policies) for more complex classes is doubtful, these methods may still be practical ways to align LLMs, that is, by aligning the sign of $t_{\theta}(w)$ with that of z0.5z-0.5. In particular, the smoothed objective of Horowitz (1992), while non-convex and potentially very rough, is differentiable so that first-order updates can be applied. This essentially ensures the LLM aligns (in the sense of divergence-constrained max-reward) with some reward whose differences align with the signs of true reward differences, even if the rewards are in fact different and lead to different policies – this is of course still valuable. Notice that if instead of smoothing, we used the logistic loss as a surrogate for the 0-1 loss we would have the DPO objective.
[…]
8 Related Literature
RLHF/RLAIF.
From the first demonstrations of reward learning from preferences (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020) to instruction-following models (Ouyang et al., 2022) and safety-oriented alignments (Bai et al., 2022b), RLHF/RLAIF is a primary ingredient in modern LLM practice. Zhan et al. (2023) prove regret bounds for offline RL from preference data. While RLHF for LLM alignment often employs KL constraints or penalties, Wang et al. (2023) study the use of other $f$-divergences, and Huang et al. (2024) analyze the use of $\chi^{2}$-divergence.
DPO (Rafailov et al., 2023) avoids explicit reward modeling by recognizing the optimal policy already encodes the reward, so the aligned policy can be trained by fitting the implied reward function to the data, assuming a known logistic link function. We follow the idea of implied reward function but drop the known link function (and generalize KL to any $f$-divergence). Assuming a known link between reward and preferences is generally common to RLHF (Rafailov et al., 2023; Glaese et al., 2022; Ziegler et al., 2019; Ibarz et al., 2018).
A variety of alignment methods depart from the structural/generative framework, where preferences are assumed to optimize a random utility or equivalently that preferences are generated by some conditional probability distribution given demonstrations and reward. Ethayarajh et al. (2024) consider loss-aversion in the spirit of the prospect theory of human decision making (as compared to utility-maximizing). Kong et al. (2024); Liang et al. (2024) consider non-stochastic corruptions to labels.
Econometric choice and semiparametrics.
Discrete choice modeling in econometrics was pioneered by McFadden (1974), who formalized random utility models and established the multinomial logit as a tractable framework for estimating preferences and welfare from choice data. Subsequent work expanded the basic logit to allow for richer substitution patterns and unobserved heterogeneity, including nested logit (McFadden, 1978), mixed logit and random coefficients models (Berry, 1995; Train, 2009), and more general GEV formulations (McFadden, 1981). Identification and estimation of structural parameters in the presence of price endogeneity and unobserved product characteristics led to instrumental-variables approaches such as the BLP framework (Berry et al., 1995, 2004), along with related control-function and inversion-based methods (Newey, 1999; Blundell and Powell, 2004). Discrete choice models have also been extended to dynamic settings to study intertemporal decision-making, beginning with Rust (1987) and followed by a large literature on estimation, identification, and computational methods for dynamic discrete choice (Hotz and Miller, 1993; Aguirregabiria and Mira, 2002; Arcidiacono and Miller, 2011).
Semiparametric binary choice with an unknown link function dates to Cosslett (1983), with efficient estimation first achieved by Klein and Spady (1993). Maximum score estimation (Manski, 1975, 1985) and its smoothed variants (Horowitz, 1992) provide robustness to link misspecification, while the cube-root asymptotics of such estimators were characterized by Kim and Pollard (1990). Closely related order-restricted and isotonic likelihood methods are developed in Barlow et al. (1972); Robertson et al. (1988). More broadly, these models fall within the class of single-index models, which extend beyond binary choice and admit alternative estimation strategies such as semiparametric least squares (Ichimura, 1993), average derivative estimators (Powell et al., 1989), and rank-based or rank-correlation estimators (Han, 1987; Sherman, 1993). This literature emphasizes identification and estimation of structural parameters under minimal distributional assumptions beyond an underlying parametric structural model.
9 Conclusion
Divergence-constrained alignment of LLMs naturally induces a semiparametric single-index model with the LLM policy class parametrizing the index. This requires only the probability one prefers one demonstration over another be some unknown increasing function of the difference in rewards. This significantly relaxes assuming the link function is known exactly, as in DPO. However, it means we no longer have a clean likelihood in the policy parameter alone. Inspired by Cosslett (1983), we considered an orthogonalized quasi-likelihood with a plug-in for a high-dimensional nuisance function characterizing the distribution of preference given any candidate index. Our analysis gave finite-sample bounds and admitted nonparametric function classes with indices identifiable only up to scale. The dependence on plug-in error was squared, meaning it generally does not affect the learning rate, and the learner behaves as though the link function were known. The semiparametric preference optimization framework developed thus allows to align LLMs to human and AI feedback in a way robust to how these are linked to rewards.
Appendix A Proof of Theorem 2
For each $\theta\in\Theta$, write $\beta_{\theta}=\beta_{\kappa,h_{\theta}}$ and $\pi_{\theta}=\pi_{\kappa,h_{\theta}}$. Define the canonical scores
$u_{\theta}(x,y)=f^{\prime}\bigl(r_{\theta}(x,y)\bigr),\qquad u_{\theta^{\star}}(x,y)=f^{\prime}\bigl(r_{\theta^{\star}}(x,y)\bigr),$
where $r_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}$. And for brevity, write $D(\theta)=\|\pi_{\kappa,\theta}-\pi^{\star}\|_{1}$.
Step 1: Existence and positivity of $\beta_{\theta}$
For a fixed $h$ and $x$, the set of policies maximizing $y\mapsto h(x,y)$ consists of all distributions supported on $S_{h}(x)=\arg\max_{y}h(x,y)$. Among these, the unique minimizer of $D_{f}(\,\cdot\,\|\,\pi_{\mathrm{ref}}(\cdot\mid x))$ is given by
$\pi_{\mathrm{greedy},h}(y\mid x)=\begin{cases}\dfrac{\pi_{\mathrm{ref}}(y\mid x)}{\omega_{h}(x)},&y\in S_{h}(x),\\[5.0pt] 0,&\text{otherwise},\end{cases}$
where $\omega_{h}(x)=\sum_{y\in S_{h}(x)}\pi_{\mathrm{ref}}(y\mid x)$. We have $D_{f}\bigl(\pi_{\mathrm{greedy},h}(\cdot\mid x)\,\big\|\,\pi_{\mathrm{ref}}(\cdot\mid x)\bigr)=\omega_{h}(x)\,f\!\bigl(1/\omega_{h}(x)\bigr).$ Thus the minimal divergence among $h$-greedy policies is $\Psi_{h}=\mathbb{E}_{x}\Bigl[\omega_{h}(x)\,f\!\bigl(1/\omega_{h}(x)\bigr)\Bigr].$
For each $\theta$, denote $\Psi_{\theta}=\Psi_{h_{\theta}}$. Under the assumptions on $f$, $\Phi_{\theta}(\beta)=\mathbb{E}_{x}D_{f}\bigl(\pi_{\beta,h_{\theta}}(\cdot\mid x)\,\big\|\,\pi_{\mathrm{ref}}(\cdot\mid x)\bigr)$ is continuous and strictly decreasing on $(0,\infty)$, with $\lim_{\beta\downarrow 0}\Phi_{\theta}(\beta)=\Psi_{\theta}>\kappa,\qquad\lim_{\beta\to\infty}\Phi_{\theta}(\beta)=0.$ So, there exists a unique $\beta_{\theta}\in(0,\infty)$ such that $\Phi_{\theta}(\beta_{\theta})=\kappa$. In particular, $\beta_{\star}=\beta_{\theta^{\star}}>0$.
Step 2: Relating policy discrepancy to canonical scores
Lemma 4.
Under the curvature lower bound $f^{\prime\prime}\geq m_{f}>0$ along the ratios $r_{\theta}$ and $r_{\theta^{\star}}$, $D(\theta)\ \leq\ \frac{1}{m_{f}}\,\|u_{\theta}-u_{\theta^{\star}}\|_{2}$ for all $\theta\in\Theta$.
Proof.
Fix $(x,y)$. Let $r_{1}=r_{\theta}(x,y)$, $r_{2}=r_{\theta^{\star}}(x,y)$ and $u_{j}=f^{\prime}(r_{j})$, $j=1,2$. By the mean value theorem there exists $\xi$ between $r_{1}$ and $r_{2}$ such that $u_{1}-u_{2}=f^{\prime\prime}(\xi)(r_{1}-r_{2}).$ Since $f^{\prime\prime}(\xi)\geq m_{f}$, we get $|r_{1}-r_{2}|\leq m_{f}^{-1}|u_{1}-u_{2}|$.
Define $\widetilde{D}(\theta)=\mathbb{E}_{x}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,\bigl|r_{\theta}(x,y)-r_{\theta^{\star}}(x,y)\bigr|.$ Then $\widetilde{D}(\theta)\leq\frac{1}{m_{f}}\mathbb{E}_{x}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,\bigl|u_{\theta}(x,y)-u_{\theta^{\star}}(x,y)\bigr|\leq\frac{1}{m_{f}}\,\|u_{\theta}-u_{\theta^{\star}}\|_{2},$ using Cauchy-Schwarz in the last step.
Since $\pi_{\theta}(y\mid x)=\pi_{\mathrm{ref}}(y\mid x)r_{\theta}(x,y)$ and $\pi_{\mathrm{ref}}\leq 1$, $D(\theta)=\mathbb{E}_{x}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)^{2}\bigl|r_{\theta}(x,y)-r_{\theta^{\star}}(x,y)\bigr|\leq\widetilde{D}(\theta),$ which yields the claim. ∎
Step 3: Lipschitz dependence of $u_{\beta,h}$ on $h$ for fixed $\beta$
Lemma 5.
Fix $\beta>0$. Under the bounds $m_{f}\leq f^{\prime\prime}\leq M_{f}$ along the relevant ratios, for any two bounded $h_{1},h_{2}$,
$\|u_{\beta,h_{1}}-u_{\beta,h_{2}}\|_{2}\ \leq\ \frac{1}{\beta}\sqrt{\frac{M_{f}}{m_{f}}}\,\|h_{1}-h_{2}\|_{2},$
where $u_{\beta,h}(x,y)=f^{\prime}(r_{\beta,h}(x,y))$ and $r_{\beta,h}=\pi_{\beta,h}/\pi_{\mathrm{ref}}$.
Proof.
Fix $x$ and for brevity omit it from arguments. For a given $h$, write $u(y)=u_{\beta,h}(y)=\beta^{-1}(h(y)-\lambda)$, where $\lambda=\lambda_{\beta,h}(x)$. Let $g=(f^{\prime})^{-1}$, so $r(y)=g(u(y))$ and the normalization constraint is $\sum_{y\in\mathcal{Y}}\pi_{\mathrm{ref}}(y\mid x)\,g(u(y))=1.$
Consider a perturbation $\Delta h$, with corresponding $\Delta u$ and $\Delta\lambda$ at fixed $\beta$. From $u(y)=\beta^{-1}(h(y)-\lambda)$ we get $\Delta u(y)=\beta^{-1}\bigl(\Delta h(y)-\Delta\lambda\bigr).$ Differentiating the normalization constraint yields $0=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,g^{\prime}(u(y))\,\Delta u(y)=\sum_{y}w(y)\Delta u(y),$ where $w(y)=\pi_{\mathrm{ref}}(y\mid x)g^{\prime}(u(y))$. Using the bounds $m_{f}\leq f^{\prime\prime}\leq M_{f}$ and $\pi_{\mathrm{ref}}\in[\varepsilon,1]$, we have $\frac{\varepsilon}{M_{f}}\ \leq\ w(y)\ \leq\ \frac{1}{m_{f}}.$
Substituting $\Delta u(y)=\beta^{-1}(\Delta h(y)-\Delta\lambda)$ gives $0=\frac{1}{\beta}\Bigl(\sum_{y}w(y)\Delta h(y)-\Delta\lambda\sum_{y}w(y)\Bigr),$ hence $\Delta\lambda=\frac{\sum_{y}w(y)\Delta h(y)}{\sum_{y}w(y)}.$ Thus $\Delta u(y)=\frac{1}{\beta}\Bigl(\Delta h(y)-\overline{\Delta h}\Bigr),\qquad\overline{\Delta h}=\frac{\sum_{y}w(y)\Delta h(y)}{\sum_{y}w(y)}.$
Define weighted norms $\|v\|_{w}^{2}=\sum_{y}w(y)v(y)^{2},~\|v\|_{\pi}^{2}=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)v(y)^{2}.$ The map $P_{w}:\Delta h\mapsto\overline{\Delta h}$ is the orthogonal projection in $\|\cdot\|_{w}$ onto the constants, so $\|\Delta u\|_{w}=\frac{1}{\beta}\|(I-P_{w})\Delta h\|_{w}\leq\frac{1}{\beta}\|\Delta h\|_{w}.$ Since $w$ and $\pi_{\mathrm{ref}}$ are comparable, for any $v$, $\|v\|_{\pi}^{2}\leq M_{f}\|v\|_{w}^{2},~\|v\|_{w}^{2}\leq\frac{1}{m_{f}}\|v\|_{\pi}^{2}.$ Therefore $\|\Delta u\|_{\pi}\leq\sqrt{M_{f}}\|\Delta u\|_{w}\leq\frac{\sqrt{M_{f}}}{\beta}\|\Delta h\|_{w}\leq\frac{1}{\beta}\sqrt{\frac{M_{f}}{m_{f}}}\|\Delta h\|_{\pi}.$ This bound holds pointwise in $x$. Squaring and integrating over $x$ yields the result. ∎
Step 4: Affine invariance and rescaling
We next use the affine invariance of the construction.
Lemma 6.
Let $h$, $a>0$, and $b:\mathcal{X}\to\mathbb{R}$ be given. Define $\tilde{h}(x,y)=a\,h(x,y)+b(x)$. Then $\pi_{\beta,\tilde{h}}=\pi_{a\beta,h}$ for all $\beta>0$. Consequently, $\beta_{\kappa,\tilde{h}}=\frac{\beta_{\kappa,h}}{a}$ and $\pi_{\kappa,\tilde{h}}=\pi_{\kappa,h}.$
Proof.
Set $\tilde{\beta}=a\beta$ and $\tilde{\lambda}_{\tilde{\beta},h}(x)=\frac{\lambda_{\beta,\tilde{h}}(x)-b(x)}{a}.$ Then $\tilde{\beta}^{-1}\bigl(h(x,y)-\tilde{\lambda}_{\tilde{\beta},h}(x)\bigr)=\beta^{-1}\bigl(ah(x,y)+b(x)-\lambda_{\beta,\tilde{h}}(x)\bigr),$ so $\pi_{\beta,\tilde{h}}(y\mid x)=\pi_{\mathrm{ref}}(y\mid x)\,(f^{\prime})^{-1}\!\Bigl(\tilde{\beta}^{-1}\bigl(h(x,y)-\tilde{\lambda}_{\tilde{\beta},h}(x)\bigr)\Bigr)=\pi_{\tilde{\beta},h}(y\mid x)=\pi_{a\beta,h}(y\mid x).$
For the divergence constraint, $\mathbb{E}_{x}D_{f}\bigl(\pi_{\beta,\tilde{h}}(\cdot\mid x)\,\big\|\,\pi_{\mathrm{ref}}(\cdot\mid x)\bigr)=\mathbb{E}_{x}D_{f}\bigl(\pi_{a\beta,h}(\cdot\mid x)\,\big\|\,\pi_{\mathrm{ref}}(\cdot\mid x)\bigr).$ If $\beta_{\kappa,h}$ is the unique $\beta$ such that the right-hand side equals $\kappa$, then the unique $\beta$ such that the left-hand side equals $\kappa$ is $\beta_{\kappa,\tilde{h}}=\beta_{\kappa,h}/a$, and the induced policies coincide at the constraint: $\pi_{\kappa,\tilde{h}}=\pi_{\beta_{\kappa,\tilde{h}},\tilde{h}}=\pi_{\beta_{\kappa,h},h}=\pi_{\kappa,h}.$
Step 5: Completing the proof of the theorem
Proof of Theorem 2.
Fix $\theta\in\Theta$ and recall $\beta_{\theta}=\beta_{\kappa,h_{\theta}}$ and $\beta_{\star}=\beta_{\theta^{\star}}$. By Lemma 6, for $a_{\theta}=\frac{\beta_{\theta}}{\beta_{\star}},$ we have that $\tilde{h}_{\theta}=a_{\theta}h_{\theta}$ satisfies $\beta_{\kappa,\tilde{h}_{\theta}}=\beta_{\star},~\pi_{\kappa,\tilde{h}_{\theta}}=\pi_{\kappa,h_{\theta}}=\pi_{\theta}.$ In particular, at $\beta_{\star}$ we have $u_{\theta}(x,y)=f^{\prime}\bigl(r_{\theta}(x,y)\bigr)=f^{\prime}\bigl(r_{\beta_{\star},\tilde{h}_{\theta}}(x,y)\bigr)=u_{\beta_{\star},\tilde{h}_{\theta}}(x,y),$ and similarly
$u_{\theta^{\star}}(x,y)=f^{\prime}\bigl(r_{\theta^{\star}}(x,y)\bigr)=f^{\prime}\bigl(r_{\beta_{\star},h_{\theta^{\star}}}(x,y)\bigr)=u_{\beta_{\star},h_{\theta^{\star}}}(x,y).$
Applying Lemma 5 with $\beta=\beta_{\star}$, $h_{1}=\tilde{h}_{\theta}$, $h_{2}=h_{\theta^{\star}}$ yields $\|u_{\theta}-u_{\theta^{\star}}\|_{2}=\|u_{\beta_{\star},\tilde{h}_{\theta}}-u_{\beta_{\star},h_{\theta^{\star}}}\|_{2}\ \leq\ \frac{1}{\beta_{\star}}\sqrt{\frac{M_{f}}{m_{f}}}\,\|\tilde{h}_{\theta}-h_{\theta^{\star}}\|_{2}.$
For each fixed scale $a>0$, $\inf_{b(x)}\Bigl\|ah_{\theta}(x,y)-b(x)-h_{\theta^{\star}}(x,y)\Bigr\|_{2}$ is achieved by centering $ah_{\theta}-h_{\theta^{\star}}$ in $y$ under $\pi_{\mathrm{ref}}(\cdot\mid x)$, and $\rho(h_{\theta})$ is the infimum of this quantity over all $a>0$. In particular, taking $a=a_{\theta}$ and $b=0$ gives $\rho(h_{\theta})\ \leq\ \|\tilde{h}_{\theta}-h_{\theta^{\star}}\|_{2}.$ Therefore, $\|u_{\theta}-u_{\theta^{\star}}\|_{2}\ \leq\ \frac{1}{\beta_{\star}}\sqrt{\frac{M_{f}}{m_{f}}}\,\rho(h_{\theta}).$
Combining this with Lemma 4 shows $D(\theta)\ \leq\ \frac{1}{m_{f}}\|u_{\theta}-u_{\theta^{\star}}\|_{2}\ \leq\ \frac{1}{m_{f}\beta_{\star}}\sqrt{\frac{M_{f}}{m_{f}}}\,\rho(h_{\theta})$. ∎
Appendix B Proof of Theorem 3
Proof.
Let the premise of Assumption 4 hold: we have $\theta\in\Theta$, a monotonic function $m$, and a set $S\subset\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}$ with $P(S)=1$ and $t_{\theta^{\star}}(x;y_{1},y_{0})=m(t_{\theta}(x;y_{1},y_{0}))\ \forall(x,y_{0},y_{1})\in S^{\prime}$. Let $M_{\theta}=\|t_{\theta}\|_{\infty}$. We will proceed to show that $t_{\theta^{\star}}(x;y_{1},y_{0})=at_{\theta}(x;y_{1},y_{0})$ $P$-almost surely for some $a\geq 0$. Then, for any $x$ of interest, fix some choice $y_{x}\in\mathrm{Support}(\pi_{\rm ref}(\cdot\mid x))$ and set $b(x)=h_{\theta^{\star}}(x,y_{x})-ah_{\theta}(x,y_{x})$. Then for any $y\in\mathrm{Support}(\pi_{\rm ref}(\cdot\mid x))$, by Assumption 3 we have that $y_{x},y\in\mathrm{Support}(P(y_{0},y_{1}\mid x))$, so that $t_{\theta^{\star}}(x;y,y_{x})=at_{\theta}(x;y,y_{x})$ gives exactly the desired result.
Since the support is the smallest measure-1 set and the monotonic mapping applies on $S$ with $P(S)=1$, applying the algebra in the display equations preceding the statement of the theorem imply that for every $u\in(-M_{\theta},M_{\theta})$ and $v\in(-\delta,\delta)$ with $u+v\in(-M_{\theta},M_{\theta})$, we have that $m(u+v)=m(u)+m(v)$.
Let any $u\in(-M_{\theta},M_{\theta})$ and $v\in(-M_{\theta},M_{\theta})$ with $u+v\in(-M_{\theta},M_{\theta})$. Write $v=k\epsilon+\iota$ where $k\in\mathbb{N}_{0}$ and $0<\left\lvert\iota\right\rvert<\left\lvert\epsilon\right\rvert<\delta$. Then $m(u+v)=m(u+\iota+(k-1)\epsilon+\epsilon)=m(u+\iota+(k-1)\epsilon)+m(\epsilon)=\cdots=m(u+\iota)+km(\epsilon)=m(u)+m(\iota)+km(\epsilon)=m(u)+m(\epsilon+\iota)+(k-1)m(\epsilon)=\cdots=m(u)+m(k\epsilon+\iota)=m(u)+m(v)$.
Thus we have a monotonic function satisfying Cauchy’s functional equation on an interval containing zero. Therefore, it must be linear on this interval. Call the coefficient $a$. Since $M_{\theta}=\|t_{\theta}\|_{\infty}$, the interval contains the relevant values so as to ensure that $m(t_{\theta^{\star}}(x;y_{1},y_{0}))=at_{\theta^{\star}}(x;y_{1},y_{0})$ $P$-almost surely. ∎
Appendix C Proof of Theorem 4
We first prove a uniform law of large numbers. Then, we prove upper semi-continuity. Then, we prove all maximizers are scalings of $t^{\star}$. Finally, we leverage compactness of the rescaled indices to establish the convergence.
In the following, define
$L_{n}^{\rm PSPO}(t)$ $=\sup_{\Psi\uparrow}\;\frac{1}{n}\sum_{i=1}^{n}\ell(t,\Psi;w_{i},z_{i}),$ $L^{\rm PSPO}(t)$ $=\sup_{\Psi\uparrow}\mathbb{E}\big[\ell(t,\Psi;w,z)\big],$
[…]
$\omega(p,q)$ $=p\log q+(1-p)\log(1-q),$ $\varpi(p)$ $=\omega(p,p).$
C.1 Uniform Law of Large Numbers
Lemma 7 (Isotonic profiling ULLN for $L_{n}^{\rm PSPO}$).
Suppose Assumption 3 holds so that $M=\sup_{t\in\mathcal{T}}\|t\|_{\infty}<\infty$. Moreover, suppose that the class $\{\mathbb{I}[t(w)\leq\tau]:t\in\mathcal{T},~\tau\in[-M,M]\}$ is $P$-Glivenko-Cantelli. Then,
$\sup_{t\in\mathcal{T}}\Big|L_{n}^{\rm PSPO}(t)-L^{\rm PSPO}(t)\Big|\ \xrightarrow{p}\ 0.$
Proof.
Let any monotone $\Psi:[-M,M]\to[0,1]$ be given. It has variation bounded by $1$ so it has a Lebesgue-Stieltjes representation as $\Psi(u)=\Psi(-M)+\int_{(-M,M]}\mathbb{I}\left[u>v\right]\mu_{\Psi}(dv)$ for a positive measure $\mu_{\Psi}$ on [M,M][-M,M] with mass at most 1. For any $t\in\mathcal{T}$, we can therefore write $\Psi(t(w))=\Psi(-M)+\int_{(-M,M]}\mathbb{I}\left[t(w)>v\right]\mu_{\Psi}(dv)$. Therefore,
$\sup_{\Psi\uparrow,t\in\mathcal{T}}\left\lvert(P_{n}-P)\Psi(t(w))\right\rvert$ $\leq\sup_{\Psi\uparrow,t\in\mathcal{T}}\left\lvert\int_{(-M,M]}(P_{n}-P)\mathbb{I}\left[t(w)>v\right]\mu_{\Psi}(dv)\right\rvert$
$\leq\sup_{t\in\mathcal{T},v\in[-M,M]}\left\lvert(P_{n}-P)\mathbb{I}\left[t(w)>v\right]\right\rvert$.
Since the class of thresholded indices is $P$-Glivenko-Cantelli, we conclude that $\{\Psi(t(w)):\Psi\uparrow,\,t\in\mathcal{T}\}$ is $P$-Glivenko-Cantelli.
Since $\Psi^{\star}$ is strictly increasing, $\mathbb{E}[z\mid w]$ only takes values in $\varepsilon,1-\varepsilon]$ for $\varepsilon=\min\{\Psi^{\star}(-M),1-\Psi^{\star}(M),0.25\}>0$. Note that $\mathbb{E}\ell(t,\Psi;w,z)=\mathbb{E}\ell(t,\Psi;w,\mathbb{E}[z\mid w])$. Since $q\mapsto p\log q+(1-p)\log(1-q)$ is strictly concave, strictly increasing until $p$, and strictly decreasing after $p$, we have that $\sup_{\Psi\uparrow\,[-M,M]\to[0,1]}\mathbb{E}\ell(t,\Psi;w,z)=\sup_{\Psi\uparrow\,[-M,M]\to[\varepsilon,1-\varepsilon]}\mathbb{E}\ell(t,\Psi;w,z)$. Since $q\mapsto p\log q+(1-p)\log(1-q)$ is Lipschitz on $\varepsilon,1-\varepsilon]$, preservation of $P$-Glivenko-Cantelli under transformation gives the result. ∎
Since both $L^{\rm PSPO}$ and $L_{n}^{\rm PSPO}$ are scale invariant, an immediate corollary of Lemma 7 is that
$\sup_{t\in\bar{\mathcal{T}}}\Big|L_{n}^{\rm PSPO}(t)-L^{\rm PSPO}(t)\Big|\ \xrightarrow{p}\ 0.$
C.2 Upper Semi-Continuity
Lemma 8.
[…]
Proof.
Let $t_{m}$ be a sequence with $\|t_{m}-t\|\to 0$. For each $m$, pick monotonic $\Psi_{m}:[-M,M]\to[0,1]$ such that $\mathbb{E}[\ell(t_{m},\Psi_{m};w,z)]\geq L^{\rm PSPO}(t_{m})-1/m$. Since $t_{m}\to t$ in $L_{2}$, along some subsequence we have $t_{m_{i}}\to t$ a.s. Since each $\Psi_{m_{i}}$ is nondecreasing and takes values in $[0,1]$, Helly’s selection theorem yields a further subsequence such that $\Psi_{m_{i_{j}}}(u)\to\Psi(u)$ for all continuity points $u$ of $\Psi$ for some nondecreasing $\Psi:[-M,M]\to[0,1]$. Fix $w_{0}$ such that $t_{m}(w_{0})\to t(w_{0})=u$ and $u$ is a continuity point of $\Psi$. Since the set of discontinuities of a monotone function is countable and hence has dense complement, choose continuity points $u_{k}\uparrow u$ and $u^{\prime}_{k}\downarrow u$ of $\Psi$. For each $k$, for all sufficiently large $j$, $u_{k}<t_{m_{i_{j}}}(w_{0})<u^{\prime}_{k}$, so by monotonicity $\Psi_{m_{i_{j}}}(u_{k})\leq\Psi_{m_{i_{j}}}(t_{m_{i_{j}}}(w_{0}))\leq\Psi_{m_{i_{j}}}(u^{\prime}_{k})$. Letting $j\to\infty$ gives $\Psi(u_{k})\leq\liminf_{j}\Psi_{m_{i_{j}}}(t_{m_{i_{j}}}(w_{0}))\leq\limsup_{j}\Psi_{m_{i_{j}}}(t_{m_{i_{j}}}(w_{0}))\leq\Psi(u^{\prime}_{j})$, and then taking $k\to\infty$ yields $\Psi_{m_{i_{j}}}(t_{m_{i_{j}}}(w_{0}))\to\Psi(u)$ by continuity of $\Psi$ at $u$. Because $\Psi$ has only countably many discontinuities and $t(w)$ has no atoms, $\mathbb{P}(t(w)\ \text{is a continuity point of }\Psi)=1$, hence $\Psi_{m_{i_{j}}}(t_{m_{i_{j}}}(w))\to\Psi(t(w))$ a.s. By dominated convergence, $\lim_{j\to\infty}\mathbb{E}[\ell(t_{m_{i_{j}}},\Psi_{m_{i_{j}}};w,z)]\;=\;\mathbb{E}[\ell(t,\Psi;w,z)]\;\leq\;L(t),$ which proves thats $\limsup_{m}L(t_{m})\leq L(t)$. ∎
C.3 Maximizers are scalings of $t^{\star}$
Lemma 9.
Under the conditions of Theorem 4, $t$ is a maximizer of $L^{\rm PSPO}(t)$ if and only if $t\neq 0$ has $t^{\star}(w)=at(w)$ almost surely for some $a\geq 0$.
Proof.
Let $\eta(w)=\mathbb{E}[z\mid w]=\Psi^{\star}(t^{\star}(w))$ and $g^{\star}_{t}(u)=\mathbb{E}[z\mid t(w)=u]=\mathbb{E}[\eta(w)\mid t(w)=u]$. Then, $\sup_{t\in\mathcal{T}}L^{\rm PSPO}(t)=L^{\rm PSPO}(t^{\star})=\mathbb{E}\omega(z,\eta(w))=\mathbb{E}\varpi(\eta(w))$. Let $J_{t}=\sup_{q:\mathbb{R}\to(0,1)}\mathbb{E}\omega(z,q(t(w)))$ where $q$ is unrestricted. Note that $\mathbb{E}[\omega(z,q)\mid t(w)]=\omega(g^{\star}_{t}(t(w)),q)$, which is strictly concave in $q$ with unique maximizer $q=g^{\star}_{t}(t(w))$. Therefore, $J_{t}=\mathbb{E}\omega(z,g^{\star}_{t}(t(w)))=\mathbb{E}\varpi(g^{\star}_{t}(t(w)))$, where the latter equality is by iterating expectations on $t(w)$. Since $L^{\rm PSPO}(t)$ is defined by restricting to monotone $\Psi$ while $J_{t}$ is defined with $q$ unrestricted, we have $L^{\rm PSPO}(t)\leq J_{t}$. By Jensen’s, $\mathbb{E}[\varpi(\eta(w))\mid t(w)]\geq\varpi(\mathbb{E}[\eta(w)\mid t(w)])=\varpi(g^{\star}_{t}(t(w)))$. Taking expectations, $L^{\rm PSPO}(t^{\star})\geq J_{t}$. In all, $L^{\rm PSPO}(t)\leq J_{t}\leq L^{\rm PSPO}(t^{\star})$.
Suppose $L^{\rm PSPO}(t)=L^{\rm PSPO}(t^{\star})$. Then $L^{\rm PSPO}(t)=J_{t}=L^{\rm PSPO}(t^{\star})$. We have $L^{\rm PSPO}(t)=\sup_{q\uparrow:\mathbb{R}\to(0,1)}\mathbb{E}\omega(z,q(t(w)))=\sup_{q:\mathbb{R}\to(0,1)}\mathbb{E}\omega(z,q(t(w)))=J_{t}$, while $q\mapsto\mathbb{E}[\omega(z,q)\mid t(w)=u]$ is strictly concave with unique maximizer $q=g^{\star}_{t}(u)$. Hence, any maximizer in the definition of $J_{t}$ must be almost everywhere equal to $g^{\star}_{t}$. If $g^{\star}_{t}$ were not equal to a monotonic function almost everywhere, we would have $L^{\rm PSPO}(t)<J_{t}$ because $L^{\rm PSPO}(t)$ is restricted to the closed convex subset of monotonic functions. We conclude that $t^{\star}(w)=m(t(w))$ almost surely for a monotonic function $m$ such that $m(t(w))={\Psi^{\star}}^{-1}(g^{\star}_{t}(t(w)))$ almost surely. Moreover, $\mathrm{Var}(\eta(w))\neq 0$, so $t^{\star}$ is not constant, and therefore $t(w)$ cannot be constant. Since $\mathbb{E}t(w)=0$, this means $t\neq 0$. Assumption 4 completes the proof. ∎
C.4 Completing the proof of the theorem
Proof of Theorem 4.
Let $\bar{t}^{\star}=t^{\star}/\|t^{\star}\|_{2}\in\bar{\mathcal{T}}$ and note that $L^{\rm PSPO}(\bar{t}^{\star})=L^{\rm PSPO}(t^{\star})$ by scale invariance. By Lemma 8, $L^{\rm PSPO}$ is upper semicontinuous on $\bar{\mathcal{T}}$. By Lemma 9, $\bar{t}^{\star}$ is the unique maximizer of $L^{\rm PSPO}$ over $\bar{\mathcal{T}}$: if $\bar{t}\in\bar{\mathcal{T}}$ and $L^{\rm PSPO}(\bar{t})=L^{\rm PSPO}(\bar{t}^{\star})$, write $\bar{t}=u/\|u\|_{2}$ with $u\in\mathcal{T}\setminus\{0\}$; then $L^{\rm PSPO}(u)=L^{\rm PSPO}(\bar{t})=L^{\rm PSPO}(\bar{t}^{\star})=L^{\rm PSPO}(t^{\star})$, hence $u=at^{\star}$ for some $a>0$, so $\bar{t}=\bar{t}^{\star}$.
Fix $\zeta>0$ and set $A_{\zeta}=\{\bar{t}\in\bar{\mathcal{T}}:\ \|\bar{t}-\bar{t}^{\star}\|_{2}\geq\zeta\}$, which is compact. Upper semicontinuity implies $m_{\zeta}=\max_{\bar{t}\in A_{\zeta}}L^{\rm PSPO}(\bar{t})$ is attained, and uniqueness gives $\Delta_{\zeta}=L^{\rm PSPO}(\bar{t}^{\star})-m_{\zeta}>0$. On the event $E_{n}=\{\sup_{\bar{t}\in\bar{\mathcal{T}}}|L^{\rm PSPO}_{n}(\bar{t})-L^{\rm PSPO}(\bar{t})|\leq\Delta_{\zeta}/3\}$ we have $L^{\rm PSPO}_{n}(\bar{t})\leq L^{\rm PSPO}(\bar{t})+\Delta_{\zeta}/3\leq m_{\zeta}+\Delta_{\zeta}/3=L^{\rm PSPO}(\bar{t}^{\star})-2\Delta_{\zeta}/3\leq L^{\rm PSPO}_{n}(\bar{t}^{\star})-\Delta_{\zeta}/3$ for all $\bar{t}\in A_{\zeta}$, hence every maximizer of $L^{\rm PSPO}_{n}$ over $\bar{\mathcal{T}}$ lies in $\bar{\mathcal{T}}\setminus A_{\zeta}$, i.e. within $\zeta$ of $\bar{t}^{\star}$.
If $0\notin\mathcal{T}$, then $t_{n}\neq 0$ a.s. and $\bar{t}_{n}=t_{n}/\|t_{n}\|_{2}\in\bar{\mathcal{T}}$ is well-defined; scale invariance yields $L^{\rm PSPO}_{n}(\bar{t}_{n})=L^{\rm PSPO}_{n}(t_{n})=\max_{t\in\mathcal{T}}L^{\rm PSPO}_{n}(t)=\max_{\bar{t}\in\bar{\mathcal{T}}}L^{\rm PSPO}_{n}(\bar{t})$, so $\bar{t}_{n}$ is a maximizer of $L^{\rm PSPO}_{n}$ over $\bar{\mathcal{T}}$. If $0\in\mathcal{T}$, let $\gamma=\tfrac{1}{2}(L^{\rm PSPO}(t^{\star})-L^{\rm PSPO}(0))>0$; on the event $\{|L^{\rm PSPO}_{n}(\bar{t}^{\star})-L^{\rm PSPO}(\bar{t}^{\star})|\leq\gamma\}\cap\{|L^{\rm PSPO}_{n}(0)-L^{\rm PSPO}(0)|\leq\gamma\}$ we have $L^{\rm PSPO}_{n}(\bar{t}^{\star})>L^{\rm PSPO}_{n}(0)$ and thus $t_{n}\neq 0$, so the same conclusion holds with probability tending to $1$. Therefore, for all large $n$ on an event of probability tending to $1$, $\bar{t}_{n}$ is a maximizer of $L^{\rm PSPO}_{n}$ over $\bar{\mathcal{T}}$ and hence $\|\bar{t}_{n}-\bar{t}^{\star}\|_{2}<\zeta$ on $E_{n}$. Thus $\mathbb{P}(\|\bar{t}_{n}-\bar{t}^{\star}\|_{2}\geq\zeta)\leq\mathbb{P}(E_{n}^{c})+o(1)\to 0$.
Finally, on $\{t_{n}\neq 0\}$,
$\inf_{a\geq 0}\|at_{n}-t^{\star}\|_{2}\ \leq\ \Big\|\frac{\|t^{\star}\|_{2}}{\|t_{n}\|_{2}}t_{n}-t^{\star}\Big\|_{2}=\|t^{\star}\|_{2}\,\|\bar{t}_{n}-\bar{t}^{\star}\|_{2}.$
Since $\zeta>0$ was arbitrary, $\|\bar{t}_{n}-\bar{t}^{\star}\|_{2}\to 0$ in probability implies $\inf_{a\geq 0}\|at_{n}-t^{\star}\|_{2}\to 0$ in probability, which dominates $\rho(h_{\theta})$ up to the coverage constant in Assumption 3. ∎
Appendix D Theoretical Properties of OSPO
In this section, we define
$g^{\star}_{t}(u)$ $=P(z=1\mid t(w)=u),$ $\eta(w)$ $=\mathbb{E}[z\mid w]=\Psi^{\star}(t^{\star}(w))=g^{\star}_{t^{\star}}(t^{\star}(w)),$
$L_{n}(t,g)$ $=\frac{1}{n}\sum_{i=1}^{n}\ell(t,g_{t};w_{i},z_{i}),$ $L(t,g)$ $=\mathbb{E}\ell(t,g_{t};w,z)$
[…]
$\omega(p,q)$ $=p\log q+(1-p)\log(1-q),$ $\varpi(p)$ $=\omega(p,p),$
$\varepsilon$ $=\min\{\Psi^{\star}(-M),1-\Psi^{\star}(M),0.25\}>0,$ $\tilde{\varepsilon}$ $=(1\wedge c_{\rm range})\varepsilon.$
D.1 Proof of Lemma 1
Proof.
Since $\Psi^{\star}$ is strictly increasing, $\mathbb{E}[z\mid w]$ only takes values in $\varepsilon,1-\varepsilon]$. Note that we also have $\varpi^{\prime\prime}(u)\geq\varpi^{\prime\prime}_{\min}=4$.
We have that $L^{\rm ideal}(t^{\star})-L^{\rm ideal}(t)=\mathbb{E}[\varpi(\mathbb{E}[z\mid t^{\star}])-\varpi(\mathbb{E}[\mathbb{E}[z\mid t^{\star}]\mid t])]=\mathbb{E}[\mathbb{E}[\varpi(\mathbb{E}[z\mid t^{\star}])-\varpi(\mathbb{E}[\mathbb{E}[z\mid t^{\star}]\mid t])\mid t]]$. By Taylor’s theorem $\mathbb{E}[\varpi(\mathbb{E}[z\mid t^{\star}])-\varpi(\mathbb{E}[\mathbb{E}[z\mid t^{\star}]\mid t])\mid t]\geq\frac{\varpi^{\prime\prime}_{\min}}{2}\mathrm{Var}(\mathbb{E}[z\mid t^{\star}]\mid t)$.
Since $\Psi^{\star}$ has derivative bounded below by some $c>0$ on [M,M][-M,M], mean value theorem implies that $\left\lvert\Psi^{\star}(u)-\Psi^{\star}(u^{\prime})\right\rvert\geq c\left\lvert u-u^{\prime}\right\rvert$ for $u,u^{\prime}\in[-M,M]$. Therefore, $\mathrm{Var}(\mathbb{E}[z\mid t^{\star}]\mid t)=\mathrm{Var}(\Psi^{\star}(t^{\star})\mid t)\geq c^{2}\mathrm{Var}(t^{\star}\mid t)$.
The proof is therefore completed by noting that $\mathbb{E}\mathrm{Var}(t^{\star}\mid t)=\inf_{f}\|t^{\star}-f(t_{\theta})\|^{2}$ and setting $\lambda_{0}=c^{2}_{\rm MSE}c^{2}\varpi^{\prime\prime}_{\min}/2$. ∎
D.2 Proof of Lemma 2
Proof.
Recall $\mathbb{E}[t]=\mathbb{E}[t^{\star}]=0$ since $P$ is symmetric in actions. By Lemma 1, $r=\inf_{a\in\mathbb{R}}\|t^{\star}-at\|\leq\frac{1}{c_{\rm MSE}}(\mathbb{E}L_{n}^{\rm ideal}(\theta^{\star})-\mathbb{E}L_{n}^{\rm ideal}(\theta))\leq\frac{1}{2}\mathbb{E}[t^{\star}(w)z]$.
Let $a_{0}=s/\mathbb{E}[t^{2}]$ be the unique minimizer of $\mathbb{E}[(t^{\star}-at)^{2}]$, and set $R=t^{\star}-a_{0}t$ so that $r^{2}=\mathbb{E}[R^{2}]$, $\mathbb{E}[tR]=0$, and $\mathbb{E}[(t^{\star})^{2}]=a_{0}^{2}\mathbb{E}[t^{2}]+r^{2}$. Since $z=\Psi^{\star}(t^{\star})$ with $\Psi^{\star}$ strictly increasing and $\mathbb{E}[t^{\star}]=0$, we have $\mathbb{E}[t^{\star}z]=\mathbb{E}[t^{\star}\Psi^{\star}(t^{\star})]>0$. Without loss of generality, $a_{0}>0$. Else, replace $t$ by t-t and all signs change accordingly.
Using $t=(t^{\star}-R)/a_{0}$ and $z=\Psi^{\star}(t^{\star})$, we have $\mathbb{E}[tz]=\frac{1}{a_{0}}\mathbb{E}[(t^{\star}-R)z]=\frac{1}{a_{0}}\bigl(\mathbb{E}[t^{\star}z]-\mathbb{E}[Rz]\bigr).$ Since $0\leq z\leq 1$, Cauchy-Schwarz gives $|\mathbb{E}[Rz]|\leq\|R\|_{2}\|z\|_{2}\leq r.$ We obtain $\mathbb{E}[t^{\star}z]-\mathbb{E}[Rz]\;\geq\;\mathbb{E}[t^{\star}z]-|\mathbb{E}[Rz]|\;\geq\;\mathbb{E}[t^{\star}z]-r\;\geq\;\tfrac{1}{2}\mathbb{E}[t^{\star}z].$ Hence $\mathbb{E}[tz]^{2}=\frac{(\mathbb{E}[t^{\star}z]-\mathbb{E}[Rz])^{2}}{a_{0}^{2}}\;\geq\;\frac{\mathbb{E}[t^{\star}z]^{2}}{4a_{0}^{2}}.$ From $\mathbb{E}[(t^{\star})^{2}]=a_{0}^{2}\mathbb{E}[t^{2}]+r^{2}$ we also have $\mathbb{E}[t^{2}]=\frac{\mathbb{E}[(t^{\star})^{2}]-r^{2}}{a_{0}^{2}}\;\leq\;\frac{\mathbb{E}[(t^{\star})^{2}]}{a_{0}^{2}}.$
By Chebyshev’s inequality, $\mathbb{P}(\hat{s}\leq 0)=\mathbb{P}\bigl(\hat{s}-\mathbb{E}[tz]\leq-\mathbb{E}[tz]\bigr)\leq\mathbb{P}\bigl(|\hat{s}-\mathbb{E}[tz]|\geq|\mathbb{E}[tz]|\bigr)\leq\frac{\mathrm{Var}(tz)}{\left\lvert\mathcal{D}\right\rvert\mathbb{E}[tz]^{2}}\leq\frac{\mathbb{E}[(tz)^{2}]}{\left\lvert\mathcal{D}\right\rvert\mathbb{E}[tz]^{2}}.$ Since $t^{2}z^{2}\leq t^{2}$, we get $\mathbb{P}(\hat{s}\leq 0)\leq\frac{\mathbb{E}[t^{2}]}{\left\lvert\mathcal{D}\right\rvert\mathbb{E}[tz]^{2}}\;\leq\;\frac{\mathbb{E}[(t^{\star})^{2}]/a_{0}^{2}}{\left\lvert\mathcal{D}\right\rvert\,(\mathbb{E}[t^{\star}z]^{2}/(4a_{0}^{2}))}=\frac{4}{\left\lvert\mathcal{D}\right\rvert}\,\frac{\mathbb{E}[(t^{\star})^{2}]}{\mathbb{E}[t^{\star}z]^{2}}.$
Alternatively, under the additional smoothness assumption, by Hoeffding’s inequality, we have that $\mathbb{P}(\hat{s}\leq 0)=\mathbb{P}\bigl(\hat{s}-\mathbb{E}[tz]\leq-\mathbb{E}[tz]\bigr)\leq\exp(-2\left\lvert\mathcal{D}\right\rvert\mathbb{E}[tz]^{2}/(2\|t\|_{\infty})^{2})\leq\exp(-\left\lvert\mathcal{D}\right\rvert\mathbb{E}[tz]^{2}/(2c_{\rm smooth}\|t\|_{2}^{2}))\leq\exp(-\left\lvert\mathcal{D}\right\rvert\mathbb{E}[t^{\star}z]^{2}/(8c_{\rm smooth}\mathbb{E}[(t^{\star})^{2}]))$. ∎
D.3 Proof of Theorem 5
Define the $\hat{g}-$quasi-likelihood difference, scale-equivalence error, and logistic log-loss, respectively:
[…]
$\mathcal{E}(t)=\inf_{a\in\mathbb{R}}\left\lVert t^{\star}-at\right\rVert_{2},$
$\phi_{z}(p)=-z\log p-(1-z)\log(1-p)=-\omega(z,p),$
so that $\ell(t,g;w,z)=-\phi_{z}\big(g_{t}(t(w))\big)$.
Further, we let $\hat{g}_{t}(u)=\hat{g}_{\theta}(u)$ when $t=t_{\theta}$. We also use the shorthand $q_{g}(t)(w)=g_{t}(t(w))$, $\hat{q}(t)(w)=\hat{g}(t(w))$, $q^{\star}(t)(w)=g^{\star}_{t}(t(w))$. We then define the errors
$\Delta q(t)=\hat{q}(t)-q^{\star}(t),\qquad\zeta=\sup_{t\in\mathcal{T}}\left\lVert\Delta q(t)\right\rVert_{4}.$
Lastly, let $\hat{t}=t_{\hat{\theta}^{\rm OSPO}}$.
D.3.1 Supporting lemmas
Lemma 10 (Derivative bounds for the log-loss).
For $p\in[{\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$] and $z\in\{0,1\}$,
$|\phi_{z}^{\prime}(p)|\leq\frac{1}{{\tilde{\varepsilon}}},\qquad|\phi_{z}^{\prime\prime}(p)|\leq\frac{1}{{\tilde{\varepsilon}}^{2}},\qquad|\phi_{z}^{\prime\prime\prime}(p)|\leq\frac{2}{{\tilde{\varepsilon}}^{3}}.$
In particular, $\phi_{z}^{\prime\prime}$ is $(2/{\tilde{\varepsilon}}^{3})$-Lipschitz on ${\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$].
Proof.
Compute $\phi_{z}^{\prime}(p)=-(z/p)+(1-z)/(1-p)$, $\phi_{z}^{\prime\prime}(p)=z/p^{2}+(1-z)/(1-p)^{2}$, and $\phi_{z}^{\prime\prime\prime}(p)=-2z/p^{3}+2(1-z)/(1-p)^{3}$. The bounds follow since $p,(1-p)\geq{\tilde{\varepsilon}}$. ∎
Lemma 11 (Score-zero identity for each fixed $t$).
For any index $t$ and any measurable $a:\mathbb{R}\to\mathbb{R}$ with $\mathbb{E}[a(t(w))^{2}]<\infty$,
$\mathbb{E}\Big[\phi_{z}^{\prime}\big(q^{\star}(t)(w)\big)\,a\big(t(w)\big)\Big]=0.$
Proof.
Write $p^{\star}(w)=q^{\star}(t)(w)=\mathbb{E}[z\mid t(w)]$. Condition on $t(w)$:
$\mathbb{E}\big[\phi_{z}^{\prime}(p^{\star}(w))\,a(t(w))\big]=\mathbb{E}\Big[a(t(w))\cdot\mathbb{E}\big[\phi_{z}^{\prime}(p^{\star}(w))\mid t(w)\big]\Big].$
Since $z\mid t(w)$ is Bernoulli with mean $p^{\star}(w)$, direct substitution gives $\mathbb{E}[\phi_{z}^{\prime}(p^{\star}(w))\mid t(w)]=0$. ∎
Lemma 12 (Oracle link minimizes log-loss conditional on $t(w)$).
For any index $t$ and any measurable $p(t(w))\in[{\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$],
$\mathbb{E}\big[\phi_{z}(p(t(w)))\big]\ \geq\ \mathbb{E}\big[\phi_{z}(q^{\star}(t)(w))\big].$
In particular, $L(t,\hat{g})\leq L(t,g^{\star})$ for every $t\in\mathcal{T}$.
Proof.
Fix $t$. For any measurable $p(t(w))\in[{\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$], $\mathbb{E}\big[\phi_{z}(p(t(w)))\mid t(w)\big]$ is the Bernoulli cross-entropy between the true conditional mean $\mathbb{E}[z\mid t(w)]=g_{t}^{\star}(t(w))$ and prediction $p(t(w))$. For each realized $t(w)$, this is minimized at $p(t(w))=g_{t}^{\star}(t(w))$, i.e. at $p(t(w))=q^{\star}(t)(w)$. Integrate over $t(w)$ to conclude. ∎
Lemma 13 (Quadratic nuisance effect at a fixed $t$).
For any $t$ such that $q^{\star}(t)(w),\hat{q}(t)(w)\in[{\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$] almost surely,
$0\ \leq\ L(t,g^{\star})-L(t,\hat{g})\ \leq\ \frac{1}{2{\tilde{\varepsilon}}^{2}}\,\mathbb{E}\big[(\Delta q(t)(w))^{2}\big]\ \leq\ \frac{1}{2{\tilde{\varepsilon}}^{2}}\,\left\lVert\Delta q(t)\right\rVert_{4}^{2}.$
Proof.
Fix $t$ and write $p^{\star}(w)=q^{\star}(t)(w)$ and $\hat{p}(w)=\hat{q}(t)(w)$, so $\Delta q(t)=\hat{p}-p^{\star}$. By Lemma 10, for all $p,p^{\prime}\in[{\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$],
$\phi_{z}(p)\leq\phi_{z}(p^{\prime})+\phi_{z}^{\prime}(p^{\prime})(p-p^{\prime})+\frac{1}{2{\tilde{\varepsilon}}^{2}}(p-p^{\prime})^{2}.$
Apply with $p=\hat{p}$ and $p^{\prime}=p^{\star}$ and take conditional expectation given $t(w)$. The linear term vanishes since $\mathbb{E}[\phi_{z}^{\prime}(p^{\star})\mid t(w)]=0$. Taking expectations yields the upper bound. The lower bound is Lemma 12. Finally, $\mathbb{E}[(\Delta q)^{2}]\leq\|\Delta q\|_{4}^{2}$. ∎
Lemma 14 (Entry into the local region in $\mathcal{E}$).
Under Assumption 7 and the conditions of Lemma 1, if
$\Delta L(\hat{g})+\frac{1}{2{\tilde{\varepsilon}}^{2}}\zeta^{2}\ <\ \lambda_{0}\,\iota^{2},$
then $\mathcal{E}(\hat{t})<\iota$.
Proof.
[…]
$).L(t^{\star},g^{\star})-L(\hat{t},g^{\star})\leq\Delta L(\hat{g})+\big(L(t^{\star},g^{\star})-L(t^{\star},\hat{g}$$\big)$.
By Lemma 13 at $t=t^{\star}$ and $\|\Delta q(t^{\star})\|_{4}\leq\zeta$, the last term is at most $(2{\tilde{\varepsilon}}^{2})^{-1}\zeta^{2}$. Therefore $L(t^{\star},g^{\star})-L(\hat{t},g^{\star})<\lambda_{0}\iota^{2}$. Lemma 1 completes the argument. ∎
Lemma 15 (Lipschitzness of likelihood to nuisance difference).
Assume Assumption 7. Define $F(t)=L(t,\hat{g})-L(t,g^{\star})$. The for all $t$ with $\mathcal{E}(t)\leq\iota$,
$|F(t^{\star})-F(t)|\ \leq\ \Big(\frac{c_{\rm Lip}^{\prime}}{{\tilde{\varepsilon}}^{2}}+\frac{c_{\rm Lip}+c_{\rm Lip}^{\prime}}{{\tilde{\varepsilon}}^{3}}\Big)\,\zeta^{2}\,\mathcal{E}(t).$
Proof.
Write $\Delta=\Delta q(t)$ and $p=q^{\star}(t)$. By the integral remainder form of Taylor’s theorem,
$\phi_{z}(p+\Delta)-\phi_{z}(p)=\phi_{z}^{\prime}(p)\Delta+\int_{0}^{1}(1-s)\phi_{z}^{\prime\prime}(p+s\Delta)\Delta^{2}\,ds.$
Taking expectation and using Lemma 11 with $a(\cdot)=\Delta(\cdot)$ gives $\mathbb{E}[\phi_{z}^{\prime}(p)\Delta]=0$, hence
$F(t)=L(t,\hat{g})-L(t,g^{\star})=-\int_{0}^{1}(1-s)\,\mathbb{E}\big[\phi_{z}^{\prime\prime}(p+s\Delta)\Delta^{2}\big]\,ds.$
Similarly for $t^{\star}$ with $p^{\prime}=q^{\star}(t^{\prime})$ and $\Delta^{\prime}=\Delta q(t^{\star})$. Thus, by triangle inequality,
$|F(t)-F(t^{\star})|$ $\leq\int_{0}^{1}(1-s)\,\mathbb{E}\Big|\phi_{z}^{\prime\prime}(p+s\Delta)\Delta^{2}-\phi_{z}^{\prime\prime}(p^{\prime}+s\Delta^{\prime})(\Delta^{\prime})^{2}\Big|\,ds$
$\leq\int_{0}^{1}(1-s)\Big[\mathbb{E}\big(|\phi_{z}^{\prime\prime}(p+s\Delta)|\,|\Delta^{2}-(\Delta^{\prime})^{2}|\big)+\mathbb{E}\big(|\phi_{z}^{\prime\prime}(p+s\Delta)-\phi_{z}^{\prime\prime}(p^{\prime}+s\Delta^{\prime})|\,(\Delta^{\prime})^{2}\big)\Big]ds.$
For the first term, Lemma 10 gives $|\phi_{z}^{\prime\prime}|\leq{\tilde{\varepsilon}}^{-2}$, hence
$\mathbb{E}\big(|\phi_{z}^{\prime\prime}(p+s\Delta)|\,|\Delta^{2}-(\Delta^{\prime})^{2}|\big)\leq\frac{1}{{\tilde{\varepsilon}}^{2}}\mathbb{E}\big(|\Delta-\Delta^{\prime}|\cdot|\Delta+\Delta^{\prime}|\big)\leq\frac{1}{{\tilde{\varepsilon}}^{2}}\|\Delta-\Delta^{\prime}\|_{4}\,\|\Delta+\Delta^{\prime}\|_{4}.$
Since $\|\Delta\|_{4},\|\Delta^{\prime}\|_{4}\leq\zeta$, we have $\|\Delta+\Delta^{\prime}\|_{4}\leq 2\zeta$. Also Assumption 7 implies $\|\Delta-\Delta^{\prime}\|_{4}\leq c_{\rm Lip}^{\prime}\zeta\,\mathcal{E}(t)$. Therefore the first term is bounded by $(2c_{\rm Lip}^{\prime}/{\tilde{\varepsilon}}^{2})\zeta^{2}\mathcal{E}(t)$.
For the second term, Lemma 10 implies $\phi_{z}^{\prime\prime}$ is $(2/{\tilde{\varepsilon}}^{3})$-Lipschitz, so
$|\phi_{z}^{\prime\prime}(p+s\Delta)-\phi_{z}^{\prime\prime}(p^{\prime}+s\Delta^{\prime})|\leq\frac{2}{{\tilde{\varepsilon}}^{3}}\,|(p-p^{\prime})+s(\Delta-\Delta^{\prime})|.$
Hence, by Cauchy-Schwarz and $\|(\Delta^{\prime})^{2}\|_{2}=\|\Delta^{\prime}\|_{4}^{2}\leq\zeta^{2}$,
$\mathbb{E}\big(|\phi_{z}^{\prime\prime}(p+s\Delta)-\phi_{z}^{\prime\prime}(p^{\prime}+s\Delta^{\prime})|\,(\Delta^{\prime})^{2}\big)\leq\frac{2}{{\tilde{\varepsilon}}^{3}}\,\|(p-p^{\prime})+s(\Delta-\Delta^{\prime})\|_{2}\,\|(\Delta^{\prime})^{2}\|_{2}\leq\frac{2\zeta^{2}}{{\tilde{\varepsilon}}^{3}}\,\|(p-p^{\prime})+s(\Delta-\Delta^{\prime})\|_{4}.$
Now $\|p-p^{\prime}\|_{4}\leq c_{\rm Lip}\mathcal{E}(t)$ by Assumption 7, and $\|\Delta-\Delta^{\prime}\|_{4}\leq c_{\rm Lip}^{\prime}\zeta\mathcal{E}(t)$ as above. Thus
$\|(p-p^{\prime})+s(\Delta-\Delta^{\prime})\|_{4}\leq(c_{\rm Lip}+c_{\rm Lip}^{\prime}\zeta)\,\mathcal{E}(t).$
Since $|\Delta q(t)(w)|\leq 1$ a.s. by the range assumption (difference of two ${\tilde{\varepsilon}},1-{\tilde{\varepsilon}}$] quantities), we have $\zeta\leq 1$, hence $c_{\rm Lip}+c_{\rm Lip}^{\prime}\zeta\leq c_{\rm Lip}+c_{\rm Lip}^{\prime}$. Therefore the second term is bounded by $(2/{\tilde{\varepsilon}}^{3})(c_{\rm Lip}+c_{\rm Lip}^{\prime})\zeta^{2}\|t-t^{\prime}\|_{2}$. Finally, integrate over $s\in[0,1]$ with weight (1s)(1-s), noting $\int_{0}^{1}(1-s)\,ds=1/2$, and obtain the desired bound. ∎
D.3.2 Proof of the theorem
Proposition 3 (Localized convergence of projected norm).
Suppose Assumptions 1, 2, 3, 5, 6 and 7 hold. Define the constant
$c=\frac{c_{\rm Lip}^{\prime}}{{\tilde{\varepsilon}}^{2}}+\frac{c_{\rm Lip}+c_{\rm Lip}^{\prime}}{{\tilde{\varepsilon}}^{3}}.$
Then, if
$\Delta L(\hat{g})+\frac{1}{2{\tilde{\varepsilon}}^{2}}\zeta^{2}\ <\ \lambda_{0}\,\iota^{2},$
then
$\mathcal{E}(\hat{t})\leq\iota,\qquad\text{and}\qquad\mathcal{E}(\hat{t})^{2}\ \leq\ \frac{2}{\lambda_{0}}\,\Delta L(\hat{g})\ +\ \frac{c^{2}}{\lambda^{2}}\,\zeta^{4}.$
Proof.
Under the premise, $\mathcal{E}(\hat{t})<\iota$ by Lemma 14.
[…]
$L(t^{\star},g^{\star})-L(\hat{t},g^{\star})$ $)\displaystyle=\big(L(t^{\star},\hat{g})-L(\hat{t},\hat{g}$$\big)$
$\quad+\Big(\big[L(\hat{t},\hat{g})-L(\hat{t},g^{\star})\big]-\big[L(t^{\star},\hat{g})-L(t^{\star},g^{\star})\big]\Big)$
$\leq\Delta L(\hat{g})+\big|F(\hat{t})-F(t^{\star})\big|.$
By Lemma 15,
[…]
Hence
[…]
Lemma 1 gives
$\lambda_{0}\,\mathcal{E}(\hat{t})^{2}\leq L(t^{\star},g^{\star})-L(\hat{t},g^{\star}).$
Combine and apply $ab\leq a^{2}/2+b^{2}/2$ with $a=\mathcal{E}(\hat{t})\sqrt{\lambda_{0}}$ and $b=c\zeta^{2}/\sqrt{\lambda_{0}}$:
$\frac{\lambda_{0}}{2}\,\mathcal{E}(\hat{t})^{2}\leq\Delta L(\hat{g})+\frac{c^{2}}{2\lambda_{0}}\,\zeta^{4}.$
Rearranging yields the stated bound. ∎
Proof of Theorem 5.
Define the local Rademacher complexity
$\mathcal{R}_{n}(\mathcal{F},\varrho)=\frac{1}{2^{n}}\sum_{\epsilon\in\{-1,1\}^{n}}\mathbb{E}_{w_{1},\dots,w_{n}\sim P}\sup_{f\in\mathcal{F}:\|f\|\leq\varrho}\left\lvert\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f(w_{i})\right\rvert$.
Consider the star hull $\mathcal{T}_{\star}=\{\beta(t_{\theta}-t^{\star}):\theta\in\Theta,\beta\in[0,1]\}$. We will first seek to bound the critical radius $\bar{\varrho}_{n}$, being the smallest solution to inequality $\mathcal{R}_{n}(\mathcal{T}_{\star},\varrho)\leq\varrho^{2}/M$. Toward that end, invoke the “refined Dudley integral” [Srebro and Sridharan, 2010, Bartlett et al., 2005]: $\mathcal{R}_{n}(\mathcal{T}_{\star},\varrho)\lesssim\inf_{\epsilon\in(0,\varrho]}\left(\epsilon+\frac{1}{\sqrt{n}}\int_{\epsilon}^{\varrho}\sqrt{\log N_{2}(2\epsilon,\mathcal{T}_{\star})}du\right)$. Then by lemma 4.5 of Mendelson [2002], we have $\log N_{2}(2\epsilon,\mathcal{T}_{\star})\lesssim\log\frac{2}{\epsilon}+\log N_{2}(\epsilon,\mathcal{T}-t^{\star})=\log\frac{2}{\epsilon}+\log N_{2}(\epsilon,\mathcal{T})$. Altogether, we find that $\bar{\varrho}_{n}\lesssim\varrho_{n}$ for $\varrho_{n}$ as given in the statement of the theorem.
[…]
$\mathcal{E}(\hat{t})\;\leq\;c_{0}\left(\varrho_{n}+\sqrt{\varrho_{n}\mathcal{E}(\hat{t})}+\sqrt{\frac{\log\log n}{n}}+\sqrt{\frac{\log(1/\delta)}{n}}+\sup_{\theta\in\Theta}\|\hat{g}_{\theta}-g^{\star}_{\theta}\|_{4}^{2}\right).$
[…]
D.4 Proof of Theorem 6
We first make some definitions local to this subsection under Assumptions 9 and 10. For the kernel, define $\mu_{1}(K)=\int\left\lvert u\right\rvert K(u)du$ and $R(K)=\int K(u)^{2}du$, let $L_{K}$ be its Lipschitz constant, and $K_{\max}$ its upper bound. Let $L_{m}$ be the Lipschitz constant of $g^{\star}_{t}(u)$. Let $f_{t}$ be the density of $t(w)$ with $f_{\min}\leq f_{t}(u)\leq f_{\max}$ for $u\in[-M,M]$. The density bounds imply the existence of $0<\sigma_{\min}\leq\sigma_{\max}<\infty$ depending only on $f_{\min},f_{\max},M$ such that $\sigma(t)=\mathrm{Var}^{1/2}(t(w))\in[\sigma_{\min},\sigma_{\max}]$ for $t\in\mathcal{T}$. We further the define the empirical mean and variance as $\bar{t}=\frac{1}{m}\sum_{i=1}^{m}t(w_{i})$, $\hat{\sigma}(t)^{2}=\frac{1}{m}\sum_{i=1}^{m}\Big(t(w_{i})-\bar{t}\Big)^{2}$.
For $b>0$ define
$\hat{f}_{t,b}(u)=\frac{1}{m}\sum_{i=1}^{m}K\!\left(\frac{u-t(w_{i})}{b}\right),\qquad\hat{F}_{t,b}(u)=\frac{1}{m}\sum_{i=1}^{m}K\!\left(\frac{u-t(w_{i})}{b}\right)z_{i},$
and $f_{t,b}(u)=\mathbb{E}[\hat{f}_{t,b}(u)]$, $F_{t,b}(u)=\mathbb{E}[\hat{F}_{t,b}(u)]$. Further, define $b_{-}=h\sigma_{\min}/2$, $b_{+}=2h\sigma_{\max}$, and $\Xi=\mathcal{T}\times[-M,M]\times[b_{-},b_{+}]$.
D.4.1 Supporting lemmas
Lemma 16 (Uniform concentration of $\hat{\sigma}(t)$).
Under Assumption 10, for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sup_{t\in\mathcal{T}}|\hat{\sigma}(t)-\sigma(t)|\ \leq\ c_{\sigma}\,\sqrt{\frac{(v+1)\log(c_{\sigma}m)+\log(1/\delta)}{m}},$
for a constant $c_{\sigma}<\infty$ depending only on $(M,\sigma_{\min},\sigma_{\max},A,v)$. Consequently, for all sufficiently large $m$,
$\hat{\sigma}(t)\in\big[\sigma_{\min}/2,\ 2\sigma_{\max}\big]\quad\text{for all }t\in\mathcal{T}.$ (14)
Proof.
The proof is a standard $\varepsilon$-net plus Hoeffding argument. Fix $\varepsilon\in(0,1]$ and let $\{t_{j}\}_{j=1}^{N}$ be an $\varepsilon$-net of $\mathcal{T}$ in $\|\cdot\|_{\infty}$, with $N\leq(A/\varepsilon)^{v}$. For any $t\in\mathcal{T}$, pick $j(t)$ with $\|t-t_{j(t)}\|_{\infty}\leq\varepsilon$. Then $|t(w_{i})-t_{j(t)}(w_{i})|\leq\varepsilon$ and $|t(w_{i})^{2}-t_{j(t)}(w_{i})^{2}|\leq 2M\varepsilon$.
Hoeffding’s inequality and a union bound over $\{t_{j}\}_{j=1}^{N}$ give that with probability at least $1-\delta/2$,
$\max_{1\leq j\leq N}\left|\frac{1}{m}\sum_{i=1}^{m}t_{j}(w_{i})-\mathbb{E}[t_{j}(w)]\right|\leq M\sqrt{\frac{2\log(2N/\delta)}{m}}.$
Thus, for any $t$,
$\left|\bar{t}-\mathbb{E}[t(w)]\right|\leq\left|\frac{1}{m}\sum_{i}t_{j(t)}(w_{i})-\mathbb{E}[t_{j(t)}(w)]\right|+2\varepsilon\leq M\sqrt{\frac{2\log(2N/\delta)}{m}}+2\varepsilon.$
Similarly, applying the same bound to $\mathcal{T}_{2}=\{t^{2}:t\in\mathcal{T}\}$ (using $|t^{2}-s^{2}|\leq 2M|t-s|$) gives
$\sup_{t\in\mathcal{T}}\left|\frac{1}{m}\sum_{i}t(w_{i})^{2}-\mathbb{E}[t(w)^{2}]\right|\leq M^{2}\sqrt{\frac{2\log(2N/\delta)}{m}}+4M\varepsilon.$
Since $\hat{\sigma}(t)^{2}=\frac{1}{m}\sum_{i}t(w_{i})^{2}-\bar{t}^{2}$ and $\sigma(t)^{2}=\mathbb{E}[t(w)^{2}]-\mathbb{E}[t(w)]^{2}$, we obtain
$\sup_{t}|\hat{\sigma}(t)^{2}-\sigma(t)^{2}|\leq c\sqrt{\frac{\log(2N/\delta)}{m}}+c\varepsilon$
for a constant $c$ depending only on $M$. Choose $\varepsilon=m^{-1}$ and use $N\leq(Am)^{v}$. Finally, $|\hat{\sigma}(t)-\sigma(t)|\leq|\hat{\sigma}(t)^{2}-\sigma(t)^{2}|/(\hat{\sigma}(t)+\sigma(t))\leq|\hat{\sigma}(t)^{2}-\sigma(t)^{2}|/\sigma_{\min}$, and absorb $1/\sigma_{\min}$ into the constant. ∎
Lemma 17 (Covering of the kernel class).
Let
$\mathcal{F}_{0}=\Big\{(w,z)\mapsto K\!\left(\frac{u-t(w)}{b}\right):\ (t,u,b)\in\Xi\Big\},\qquad\mathcal{F}=\Big\{(w,z)\mapsto K\!\left(\frac{u-t(w)}{b}\right)z:\ (t,u,b)\in\Xi\Big\}.$
Then for all $\eta\in(0,1]$,
$\log N\big(\eta,\mathcal{F}_{0},\|\cdot\|_{\infty}\big)\leq(v+2)\log\Big(\frac{c_{\mathrm{cov}}}{\eta b_{-}}\Big),\qquad\log N\big(\eta,\mathcal{F},\|\cdot\|_{\infty}\big)\leq(v+2)\log\Big(\frac{c_{\mathrm{cov}}}{\eta b_{-}}\Big),$
for a constant $c_{\mathrm{cov}}<\infty$ depending only on $(A,M,L_{K},\sigma_{\min},\sigma_{\max},v)$.
Proof.
For $(t,u,b)$ and $(t^{\prime},u^{\prime},b^{\prime})$ in $\Xi$, by Lipschitzness of $K$ and $b,b^{\prime}\geq b_{-}$,
$\sup_{w}\left|K\!\left(\frac{u-t(w)}{b}\right)-K\!\left(\frac{u^{\prime}-t^{\prime}(w)}{b^{\prime}}\right)\right|\leq L_{K}\left(\frac{|u-u^{\prime}|+\|t-t^{\prime}\|_{\infty}}{b_{-}}+\frac{2M}{b_{-}^{2}}|b-b^{\prime}|\right).$
Thus an $\eta$-cover in $\|\cdot\|_{\infty}$ is obtained by taking: (i) a $\|\cdot\|_{\infty}$ cover of $\mathcal{T}$ with mesh $\varepsilon_{t}=\eta b_{-}/(3L_{K})$, (ii) a grid of [M,M][-M,M] with mesh $\varepsilon_{u}=\eta b_{-}/(3L_{K})$, (iii) a grid of $b_{-},b_{+}]$ with mesh $\varepsilon_{b}=\eta b_{-}^{2}/(6ML_{K})$. The net sizes satisfy $N_{t}\leq(A/\varepsilon_{t})^{v}$, $N_{u}\leq 1+2M/\varepsilon_{u}$, and $N_{b}\leq 1+(b_{+}-b_{-})/\varepsilon_{b}$. Since $b_{+}-b_{-}\lesssim b_{-}$, $N_{b}\lesssim 1/\eta$ up to fixed constants. Multiplying and taking logs yields the stated bound. Since $0\leq z\leq 1$, multiplying by $z$ does not change $\|\cdot\|_{\infty}$ covering numbers. ∎
Lemma 18 (Uniform Bernstein bound).
Assume Eq. 14 holds. Then there exists $c_{\mathrm{ep}}<\infty$ depending only on $(K_{\max},R(K),f_{\max},A,v,M,L_{K},\sigma_{\min},\sigma_{\max})$ such that for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sup_{(t,u,b)\in\Xi}\big|\hat{f}_{t,b}(u)-f_{t,b}(u)\big|$ $\leq c_{\mathrm{ep}}\Bigg[\sqrt{\frac{b_{+}\Big((v+2)\log(\tfrac{c_{\mathrm{ep}}m}{b_{-}})+\log(\tfrac{1}{\delta})\Big)}{m}}+\frac{(v+2)\log(\tfrac{c_{\mathrm{ep}}m}{b_{-}})+\log(\tfrac{1}{\delta})}{m}\Bigg],$ (15)
$\sup_{(t,u,b)\in\Xi}\big|\hat{F}_{t,b}(u)-F_{t,b}(u)\big|$ $\leq c_{\mathrm{ep}}\Bigg[\sqrt{\frac{b_{+}\Big((v+2)\log(\tfrac{c_{\mathrm{ep}}m}{b_{-}})+\log(\tfrac{1}{\delta})\Big)}{m}}+\frac{(v+2)\log(\tfrac{c_{\mathrm{ep}}m}{b_{-}})+\log(\tfrac{1}{\delta})}{m}\Bigg].$ (16)
Proof.
We prove Eq. 15; Eq. 16 is identical since $0\leq z\leq 1$. Fix $(t,u,b)\in\Xi$ and define $Z_{i}=K((u-t(w_{i}))/b)\in[0,K_{\max}]$. Using Assumption 10 and the change of variables v=ubsv=u-bs,
$\mathbb{E}[Z_{i}^{2}]=\int K\!\left(\frac{u-v}{b}\right)^{2}f_{t}(v)\,dv=b\int K(s)^{2}f_{t}(u-bs)\,ds\leq bf_{\max}R(K)\leq b_{+}f_{\max}R(K)=V_{0}b_{+},$
so $\mathrm{Var}(Z_{i})\leq V_{0}b_{+}$. Bernstein’s inequality gives, for any $\alpha>0$,
$\mathbb{P}\!\left(\left|\frac{1}{m}\sum_{i=1}^{m}(Z_{i}-\mathbb{E}Z_{i})\right|\geq\sqrt{\frac{2V_{0}b_{+}\alpha}{m}}+\frac{2K_{\max}\alpha}{3m}\right)\leq 2e^{-\alpha}.$
Let $\mathcal{N}$ be an $\eta$-net of $\mathcal{F}_{0}$ in $\|\cdot\|_{\infty}$ and use Lemma 17: $|\mathcal{N}|\leq\exp((v+2)\log(\frac{c_{\mathrm{cov}}}{\eta b_{-}}))$. Set $\alpha=\log(2|\mathcal{N}|/\delta)$ and union bound over $f\in\mathcal{N}$ to get that with probability at least $1-\delta/2$,
$\sup_{f\in\mathcal{N}}\left|\frac{1}{m}\sum_{i=1}^{m}\big(f(w_{i})-\mathbb{E}f(w)\big)\right|\leq\sqrt{\frac{2V_{0}b_{+}\log(2|\mathcal{N}|/\delta)}{m}}+\frac{2K_{\max}\log(2|\mathcal{N}|/\delta)}{3m}.$
Extend from $\mathcal{N}$ to $\mathcal{F}_{0}$ by approximation: for any $f\in\mathcal{F}_{0}$, pick $\tilde{f}\in\mathcal{N}$ with $\|f-\tilde{f}\|_{\infty}\leq\eta$. Then the empirical and expected means differ by at most $2\eta$. Choose $\eta=m^{-2}$ so $2\eta\leq 1/m$ and absorb into the second term. Since $\log|\mathcal{N}|\lesssim(v+2)\log(\tfrac{m}{b_{-}})$, we obtain Eq. 15 after absorbing constants into $c_{\mathrm{ep}}$. ∎
Lemma 19 (Lower bound for the expected denominator).
Under Assumptions 3, 9 and 10, for all $(t,u,b)\in\Xi$,
$f_{t,b}(u)\geq\frac{bf_{\min}}{2}$
provided $b\leq M$.
Proof.
Fix $(t,u,b)\in\Xi$. Write $U_{t}=t(w)$ with density $f_{t}$ and change variables v=ubsv=u-bs:
$f_{t,b}(u)=\mathbb{E}\!\left[K\!\left(\frac{u-U_{t}}{b}\right)\right]=b\int K(s)f_{t}(u-bs)\,ds.$
Since $K$ is supported on [1,1][-1,1], the integral reduces to $s\in[-1,1]$. If $u\geq 0$, then for every $s\in[0,1]$ we have $u-bs\in[u-b,u]\subseteq[-M,M]$ because $u\leq M$ and $b\leq M$. If $u\leq 0$, then for every $s\in[-1,0]$ we have $u-bs\in[u,u+b]\subseteq[-M,M]$. In either case, on an interval of $s$-values of length $1$ (namely $[0,1]$ or [1,0][-1,0]) we have $u-bs\in[-M,M]$ and thus $f_{t}(u-bs)\geq f_{\min}$. By symmetry of $K$ and $\int K=1$, we have $\int_{0}^{1}K(s)\,ds=\int_{-1}^{0}K(s)\,ds=1/2$. Therefore
$f_{t,b}(u)\geq b\int_{0}^{1}K(s)f_{\min}\,ds=bf_{\min}/2\quad\text{or}\quad f_{t,b}(u)\geq b\int_{-1}^{0}K(s)f_{\min}\,ds=bf_{\min}/2,$
and the result follows. ∎
Lemma 20 (Uniform bias bound (order $b$)).
Under Assumptions 3, 9 and 10, there exists $c_{\mathrm{bias}}<\infty$ such that for all $(t,u,b)\in\Xi$,
$\left|\frac{F_{t,b}(u)}{f_{t,b}(u)}-g^{\star}_{t}(u)\right|\leq c_{\mathrm{bias}}\,b.$
Proof.
Fix $(t,u,b)\in\Xi$. Since $g^{\star}_{t}(u)=\mathbb{E}[z\mid U_{t}=u]$ and $U_{t}$ has density $f_{t}$,
$F_{t,b}(u)=\mathbb{E}\!\left[K\!\left(\frac{u-U_{t}}{b}\right)z\right]=\mathbb{E}\!\left[K\!\left(\frac{u-U_{t}}{b}\right)g^{\star}_{t}(U_{t})\right]=b\int K(s)\,g^{\star}_{t}(u-bs)\,f_{t}(u-bs)\,ds.$
Also $f_{t,b}(u)=b\int K(s)f_{t}(u-bs)\,ds$. Therefore
$F_{t,b}(u)-g^{\star}_{t}(u)f_{t,b}(u)=b\int K(s)\,\big(g^{\star}_{t}(u-bs)-g^{\star}_{t}(u)\big)\,f_{t}(u-bs)\,ds.$
Taking absolute values and using that $g^{\star}_{t}$ is $L_{m}$-Lipschitz on [M,M][-M,M] and $f_{t}\leq f_{\max}$ on [M,M][-M,M],
$|F_{t,b}(u)-g^{\star}_{t}(u)f_{t,b}(u)|\leq b\int|K(s)|\,L_{m}|bs|\,f_{\max}\,ds=L_{m}f_{\max}\mu_{1}(K)\,b^{2}.$
Divide by the denominator lower bound from Lemma 19, $f_{t,b}(u)\geq bf_{\min}/2$, to obtain
$\left|\frac{F_{t,b}(u)}{f_{t,b}(u)}-g^{\star}_{t}(u)\right|=\frac{|F_{t,b}(u)-g^{\star}_{t}(u)f_{t,b}(u)|}{f_{t,b}(u)}\leq\frac{2L_{m}f_{\max}\mu_{1}(K)}{f_{\min}}\,b.$
This is the claimed bound with $c_{\mathrm{bias}}=2L_{m}f_{\max}\mu_{1}(K)/f_{\min}$. ∎
Lemma 21 (Uniform ratio bound).
Assume Eq. 14 holds and suppose that
$\sup_{(t,u,b)\in\Xi}\big|\hat{f}_{t,b}(u)-f_{t,b}(u)\big|\ \leq\ \frac{b_{-}f_{\min}}{4}.$ (17)
Then
$\sup_{t\in\mathcal{T}}\sup_{u\in[-M,M]}\big|\hat{g}_{t}(u)-g_{t}^{\star}(u)\big|\leq c_{\mathrm{bias}}b_{+}+\frac{c_{\mathrm{rat}}}{b_{-}}\Big(\sup_{(t,u,b)\in\Xi}|\hat{F}_{t,b}(u)-F_{t,b}(u)|+\sup_{(t,u,b)\in\Xi}|\hat{f}_{t,b}(u)-f_{t,b}(u)|\Big),$
where $c_{\mathrm{rat}}<\infty$ depends only on $(f_{\min},f_{\max},\sigma_{\min},\sigma_{\max})$.
Proof.
Fix $t\in\mathcal{T}$ and $u\in[-M,M]$, and set $b_{t}=h\hat{\sigma}(t)\in[b_{-},b_{+}]$. Write $\hat{g}_{t}(u)=\hat{F}_{t,b_{t}}(u)/\hat{f}_{t,b_{t}}(u)$ and $g_{t}^{\star}(u)=g^{\star}_{t}(u)$. Add and subtract $F_{t,b_{t}}(u)/f_{t,b_{t}}(u)$:
$|\hat{g}_{t}(u)-g^{\star}_{t}(u)|\leq\left|\frac{\hat{F}_{t,b_{t}}(u)}{\hat{f}_{t,b_{t}}(u)}-\frac{F_{t,b_{t}}(u)}{f_{t,b_{t}}(u)}\right|+\left|\frac{F_{t,b_{t}}(u)}{f_{t,b_{t}}(u)}-g^{\star}_{t}(u)\right|.$
The second term is at most $c_{\mathrm{bias}}b_{t}\leq c_{\mathrm{bias}}b_{+}$ by Lemma 20. For the first term,
[…]
On Eq. 17 and Lemma 19, $f_{t,b_{t}}(u)\geq b_{t}f_{\min}/2\geq b_{-}f_{\min}/2$ and $\hat{f}_{t,b_{t}}(u)\geq f_{t,b_{t}}(u)-b_{-}f_{\min}/4\geq b_{-}f_{\min}/4$. Also $0\leq F_{t,b_{t}}(u)\leq f_{t,b_{t}}(u)\leq b_{+}f_{\max}$. Therefore
$\left|\frac{\hat{g}-g}{\hat{f}}\right|\leq\frac{4}{b_{-}f_{\min}}|\hat{g}-g|,\qquad\left|\frac{g}{\hat{f}f}(f-\hat{f})\right|\leq\frac{b_{+}f_{\max}}{(b_{-}f_{\min}/4)(b_{-}f_{\min}/2)}|\hat{f}-f|\leq\frac{8b_{+}f_{\max}}{b_{-}^{2}f_{\min}^{2}}|\hat{f}-f|.$
Taking suprema and using $b_{+}/b_{-}\leq 4\sigma_{\max}/\sigma_{\min}$ yields the claim with $c_{\mathrm{rat}}$ absorbing constants. ∎
D.4.2 Proof of the theorem
Proof of Theorem 6.
Let $\delta\in(0,0.5)$. Intersect the events of Lemma 16 (at confidence $\delta/3$) and Lemma 18 (at confidence $\delta/3$). On this intersection, Eq. 14 holds and the bounds Eqs. 15 and 16 hold.
The condition $mh\geq c_{0}((v+2)\log(c_{3}m/h)+\log(1/\delta))$ implies that the right-hand side of Eq. 15 is at most $b_{-}f_{\min}/4$ (after absorbing constants), hence Eq. 17 holds. Thus, by a union bound, with probability at least $1-\delta$ all conditions of Lemma 21 hold.
Apply Lemma 21, then substitute Eqs. 15 and 16 and $b_{+}\asymp h$, $b_{-}\asymp h$ from Eq. 14. This yields
$\sup_{t\in\mathcal{T}}\sup_{u\in[-M,M]}|\hat{g}_{t}(u)-g_{t}^{\star}(u)|\leq c_{1}h+c_{2}\left[\sqrt{\frac{(v+2)\log(\tfrac{c_{3}m}{h})+\log(1/\delta)}{mh}}+\frac{(v+2)\log(\tfrac{c_{3}m}{h})+\log(1/\delta)}{mh}\right].$
We get the stated result by adjusting the constants so that we can replace $v+2$ with $v$, noting that $v\geq 1$ so $3v\geq v+2\geq v$. ∎
Appendix E Theoretical Properties of RSPO
E.1 Proof of Proposition 2
E.1.1 Supporting lemmas
Lemma 22 (AUC as a pairwise identification success probability).
Let $(w_{0},z_{0}),(w_{1},z_{1})$ be i.i.d. and set $D=\{z_{0}\neq z_{1}\}$. Define the success indicator
$S_{t}=\mathbb{I}\left[t(w_{1})>t(w_{0}),\,z_{1}=1,z_{0}=0\right]+\mathbb{I}\left[t(w_{0})>t(w_{1}),\,z_{0}=1,z_{1}=0\right]+\frac{1}{2}\,\mathbb{I}\left[t(w_{1})=t(w_{0}),\,D\right].$
Then $\mathrm{AUC}(t)=\mathbb{E}[S_{t}\mid D]$.
Proof.
By exchangeability, $\mathbb{P}(z_{0}=0,z_{1}=1\mid D)=\mathbb{P}(z_{0}=1,z_{1}=0\mid D)=1/2$. On $\{z_{0}=0,z_{1}=1\}$ we have $S_{t}=\phi(t(w_{1})-t(w_{0}))$, and on $\{z_{0}=1,z_{1}=0\}$ we have $S_{t}=\phi(t(w_{0})-t(w_{1}))$. Moreover, the conditional law of $(w_{0},w_{1})$ given $\{z_{0}=1,z_{1}=0\}$ equals the law of $(w_{1},w_{0})$ given $\{z_{0}=0,z_{1}=1\}$, hence
$\mathbb{E}\big[\phi(t(w_{0})-t(w_{1}))\mid z_{0}=1,z_{1}=0\big]=\mathbb{E}\big[\phi(t(w_{1})-t(w_{0}))\mid z_{0}=0,z_{1}=1\big]=\mathrm{AUC}(t).$
Therefore $\mathbb{E}[S_{t}\mid D]=\mathrm{AUC}(t)$. ∎
Lemma 23 (Monotone version lemma).
Let $X$ be real-valued and $Y=g_{0}(X)$ a.s. for some Borel $g_{0}:\mathbb{R}\to\mathbb{R}$. Let $(X,Y)$ and $(X^{\prime},Y^{\prime})$ be i.i.d. copies. If $\mathbb{P}(X<X^{\prime},\,Y>Y^{\prime})=0$, then there exists a nondecreasing Borel function $g:\mathbb{R}\to\mathbb{R}$ such that $Y=g(X)$ a.s.
Proof.
Let $\mu$ be the law of $X$ and define the inversion set
$B=\{(x,x^{\prime})\in\mathbb{R}^{2}:\ x<x^{\prime},\ g_{0}(x)>g_{0}(x^{\prime})\}.$
Since $(X,Y)=(X,g_{0}(X))$ a.s. and $(X^{\prime},Y^{\prime})=(X^{\prime},g_{0}(X^{\prime}))$ a.s.,
$0=\mathbb{P}(X<X^{\prime},\,Y>Y^{\prime})=\int\!\!\int\mathbb{I}_{B}(x,x^{\prime})\,\mu(dx)\mu(dx^{\prime})=(\mu\times\mu)(B).$
By Fubini, there exists a Borel set $G\subset\mathbb{R}$ with $\mu(G)=1$ such that for every $x^{\prime}\in G$,
$\mu\big(\{x<x^{\prime}:g_{0}(x)>g_{0}(x^{\prime})\}\big)=0.$
Define
$g(s)=\sup\{g_{0}(x):x\leq s,\ x\in G\},\qquad s\in\mathbb{R},$
with the convention $\sup\emptyset=$-$\infty$ (this case occurs only for $s<\inf G$ and will be irrelevant). Then $g$ is nondecreasing by construction. For $s\in G$, the defining property implies that for $\mu$-a.e. $x<s$ with $x\in G$, $g_{0}(x)\leq g_{0}(s)$, hence $g(s)\leq g_{0}(s)$; also $g(s)\geq g_{0}(s)$ since $s$ itself is feasible in the supremum. Therefore $g(s)=g_{0}(s)$ for all $s\in G$, and since $\mu(G)=1$ we obtain $g(X)=g_{0}(X)=Y$ a.s. ∎
E.1.2 Proof of the proposition
Proof.
Let $(w_{0},z_{0}),(w_{1},z_{1})$ be i.i.d. copies and $D=\{z_{0}\neq z_{1}\}$. Fix a measurable $t$.
Pointwise Bayes optimal comparison. Condition on $(w_{0},w_{1})=(u,v)$. Write $\eta_{u}=\eta(u)$ and $\eta_{v}=\eta(v)$. Since $z_{0}\mid w_{0}=u\sim\mathrm{Bernoulli}(\eta_{u})$ and $z_{1}\mid w_{1}=v\sim\mathrm{Bernoulli}(\eta_{v})$ are conditionally independent,
$\mathbb{P}(z_{1}=1,z_{0}=0\mid w_{0}=u,w_{1}=v)=\eta_{v}(1-\eta_{u}),\qquad\mathbb{P}(z_{1}=0,z_{0}=1\mid w_{0}=u,w_{1}=v)=(1-\eta_{v})\eta_{u}.$
On $D$, define (when the denominator is 0, set $p(u,v)=1/2$; this case has $\mathbb{P}(D\mid u,v)=0$ and is irrelevant):
$p(u,v)=\mathbb{P}(z_{1}=1,z_{0}=0\mid w_{0}=u,w_{1}=v,D)=\frac{\eta_{v}(1-\eta_{u})}{\eta_{v}(1-\eta_{u})+(1-\eta_{v})\eta_{u}}.$
Therefore,
$p(u,v)>\tfrac{1}{2}\iff\eta_{v}>\eta_{u},\qquad p(u,v)=\tfrac{1}{2}\iff\eta_{v}=\eta_{u},\qquad p(u,v)<\tfrac{1}{2}\iff\eta_{v}<\eta_{u}.$
Given $(u,v,D)$, the success probability of the rule induced by $t$ (pick the larger score, break ties uniformly) equals
$\mathbb{E}[S_{t}\mid w_{0}=u,w_{1}=v,D]=\begin{cases}p(u,v),&t(v)>t(u),\\ 1-p(u,v),&t(v)<t(u),\\ \tfrac{1}{2},&t(v)=t(u).\end{cases}$
Hence, pointwise,
$\mathbb{E}[S_{t}\mid w_{0}=u,w_{1}=v,D]\ \leq\ \max\{p(u,v),1-p(u,v)\},$
with equality if and only if:
  • if $\eta_{v}>\eta_{u}$, then $t(v)>t(u)$;
  • if $\eta_{v}<\eta_{u}$, then $t(v)<t(u)$;
  • if $\eta_{v}\neq\eta_{u}$, then $t(v)\neq t(u)$ (no tie across unequal $\eta$).
Global optimality and equality conditions. Combining with Lemma 22 gives
$\mathrm{AUC}(t)=\mathbb{E}[S_{t}\mid D]\ \leq\ \mathbb{E}\!\left[\max\{p(w_{0},w_{1}),1-p(w_{0},w_{1})\}\,\big|\,D\right].$
The right-hand side is achieved by any score that ranks $w$ by $\eta(w)$ (ties allowed only when $\eta$ ties), e.g. by $\eta$ itself. Therefore,
$\sup_{t}\mathrm{AUC}(t)=\mathrm{AUC}(\eta).$
Since $\eta=\Psi^{\star}(t^{\star})$ and $\Psi^{\star}$ is strictly increasing, $\mathrm{AUC}(\eta)=\mathrm{AUC}(t^{\star})$, so $t^{\star}$ is a maximizer. Moreover, a score $t$ is a maximizer if and only if it attains equality in the pointwise bound above for $(w_{0},w_{1})$ distributed as $(w_{0},w_{1})\mid D$. Because $\eta(w)\in(0,1)$ a.s. (since $\Psi^{\star}:\mathbb{R}\to(0,1)$), $\mathbb{P}(D\mid w_{0},w_{1})>0$ a.s., so equality holds for $(w_{0},w_{1})$ under $(w_{0},w_{1})\mid D$ if and only if it holds for i.i.d. $(w_{0},w_{1})$ under $P\otimes P$. Thus, for a maximizer $t$, with i.i.d. $w,w^{\prime}$,
$\mathbb{P}\big(\eta(w)>\eta(w^{\prime}),\ t(w)\leq t(w^{\prime})\big)=0,\qquad\mathbb{P}\big(t(w)=t(w^{\prime}),\ \eta(w)\neq\eta(w^{\prime})\big)=0.$ (18)
Deduce $\eta(w)=g(t(w))$ with $g$ nondecreasing, hence $t^{\star}(w)=m(t(w))$. From the second part of Eq. 18, $\eta(w)$ is a.s. constant on level sets of $t(w)$, i.e. $\eta(w)$ is $\sigma(t(w))$-measurable. By the Doob-Dynkin lemma, there exists a Borel function $g_{0}:\mathbb{R}\to(0,1)$ such that
$\eta(w)=g_{0}(t(w))\qquad\text{a.s.}$
Using the first part of Eq. 18 and i.i.d. copies $(w,w^{\prime})$, we have
$0=\mathbb{P}\big(\eta(w)>\eta(w^{\prime}),\,t(w)<t(w^{\prime})\big)=\mathbb{P}\big(g_{0}(t(w))>g_{0}(t(w^{\prime})),\,t(w)<t(w^{\prime})\big).$
Applying Lemma 23 with $X=t(w)$ and $Y=g_{0}(t(w))=\eta(w)$ yields a nondecreasing Borel $g$ with
$\eta(w)=g(t(w))\qquad\text{a.s.}$
Finally, since $\Psi^{\star}$ is strictly increasing, it has an inverse on its range; define the (Borel) inverse
$(\Psi^{\star})^{-1}(u)=\inf\{x\in\mathbb{R}:\Psi^{\star}(x)\geq u\},\qquad u\in(0,1).$
Set $m=(\Psi^{\star})^{-1}\circ g$. Then $m$ is nondecreasing and
$t^{\star}(w)=(\Psi^{\star})^{-1}(\eta(w))=(\Psi^{\star})^{-1}(g(t(w)))=m(t(w))\qquad\text{a.s.}$
This shows that every AUC maximizer lies in the stated class.
Converse. Conversely, suppose $t^{\star}(w)=m(t(w))$ a.s. for some nondecreasing $m$. Then $\eta(w)=\Psi^{\star}(t^{\star}(w))=(\Psi^{\star}\circ m)(t(w))$, where $\Psi^{\star}\circ m$ is nondecreasing. In particular, for any $u,v$,
$\eta(u)>\eta(v)\ \Rightarrow\ t^{\star}(u)>t^{\star}(v)\ \Rightarrow\ m(t(u))>m(t(v))\ \Rightarrow\ t(u)>t(v),$
and $t(u)=t(v)\Rightarrow t^{\star}(u)=t^{\star}(v)\Rightarrow\eta(u)=\eta(v)$. Thus $t$ satisfies the pointwise equality conditions of Step 1, so $\mathrm{AUC}(t)=\mathrm{AUC}(\eta)=\mathrm{AUC}(t^{\star})$, i.e. $t$ is an AUC maximizer.
Combining both directions proves the theorem. ∎
E.2 Proof of Theorem 7
Proof.
First we make some definitions based on the assumptions. Let $I=[\underline{t^{\star}},\overline{t^{\star}}]\subseteq[-M,M]$ be the support of $t^{\star}$. Let $p$ its density satisfying $0<\underline{p}\leq p(u)\leq\overline{p}<\infty$ for all $u\in I$. Let $\eta(w)=\mathbb{P}(z=1\mid w)=\mathbb{P}(z=1\mid t^{\star}(w))=\Psi^{\star}(t^{\star}(w))\in[\varepsilon,1-\varepsilon]$ for $\varepsilon=\min\{\Psi^{\star}(-M),1-\Psi^{\star}(M),0.25\}>0$. The derivative $(\Psi^{\star})^{\prime}$ satisfies $0<\underline{\psi}\leq(\Psi^{\star})^{\prime}(u)\leq\overline{\psi}<\infty$ for all $u\in[-M,M]$. Finally, define $\underline{q}=\underline{p}/\overline{\psi}$, $\lambda_{0}=c_{\rm MSE}\,\underline{q}\,\underline{\psi}^{2}=c_{\rm MSE}\,\frac{\underline{p}}{\overline{\psi}}\,\underline{\psi}^{2}$, $B_{0}=\frac{1-\varepsilon}{\varepsilon}\,\overline{p}$, $C_{H}=6\,\varepsilon^{-1/3}\,B_{0}^{2/3}=6\,\frac{(1-\varepsilon)^{2/3}}{\varepsilon}\,\overline{p}^{2/3}$, $\gamma_{0}=\frac{1}{16}\,\underline{q}^{2}\left(\frac{\underline{\psi}}{\overline{p}}\right)^{3}=\frac{\underline{p}^{2}\,\underline{\psi}^{3}}{16\,\overline{\psi}^{2}\,\overline{p}^{3}}$, and $\epsilon_{0}=\min\Big\{\frac{\gamma_{0}}{2},\ \lambda_{0}\Big(\frac{\gamma_{0}}{2C_{H}}\Big)^{3}\Big\}$. Let $\mathrm{sgn}_{0}(u)=2\phi(u)-1\in\{-1,0,1\}$. And let $\pi_{1}=\mathbb{P}(z=1)$ and $\pi_{0}=1-\pi_{1}$ (in our setting $\pi_{0}=\pi_{1}=1/2$ by the symmetry of $P$ but the proof actually does not use this and works more generally so we keep these as symbols).
A symmetrized representation of AUC and the Bayes score. Let $(w_{0},z_{0})$ and $(w_{1},z_{1})$ be i.i.d. copies of $(w,z)$ and set $\eta_{i}=\eta(w_{i})$ and $t_{i}=t(w_{i})$. A direct conditioning on $(w_{0},w_{1})$ gives
$\pi_{0}\pi_{1}\,\mathrm{AUC}(t)=\mathbb{E}\!\left[(1-\eta_{0})\eta_{1}\,\phi(t_{1}-t_{0})\right].$ (19)
Swapping indices $0\leftrightarrow 1$ and using $\phi(u)+\phi(-u)=1$ yields
$\pi_{0}\pi_{1}\,\mathrm{AUC}(t)=\mathbb{E}\!\left[(1-\eta_{1})\eta_{0}\,\phi(t_{0}-t_{1})\right]=\pi_{0}\pi_{1}-\mathbb{E}\!\left[(1-\eta_{1})\eta_{0}\,\phi(t_{1}-t_{0})\right].$
Adding this identity to Eq. 19 gives
$2\pi_{0}\pi_{1}\,\mathrm{AUC}(t)=\pi_{0}\pi_{1}+\mathbb{E}\!\left[(\eta_{1}-\eta_{0})\,\phi(t_{1}-t_{0})\right].$
Since $\mathbb{E}[\eta_{1}-\eta_{0}]=0$ and $\phi(u)=\tfrac{1}{2}(1+\mathrm{sgn}_{0}(u))$, we obtain
$\mathrm{AUC}(t)=\frac{1}{2}+\frac{1}{4\pi_{0}\pi_{1}}\,\mathbb{E}\!\left[(\eta_{1}-\eta_{0})\,\mathrm{sgn}_{0}(t_{1}-t_{0})\right].$ (20)
Because $\Psi^{\star}$ is strictly increasing, $\mathrm{sgn}_{0}(t^{\star}_{1}-t^{\star}_{0})=\mathrm{sgn}_{0}(\eta_{1}-\eta_{0})$, hence
$\mathrm{AUC}(t^{\star})=\frac{1}{2}+\frac{1}{4\pi_{0}\pi_{1}}\,\mathbb{E}\!\left[|\eta_{1}-\eta_{0}|\right].$ (21)
AUC-regret controls the best nonlinear $L^{2}$ approximation error. Fix $t\in\mathcal{T}$ and define the posterior given the score
$\bar{\eta}(w)=\mathbb{E}[\eta(w)\mid t(w)].$
Since $(w_{0},w_{1})$ are independent and $\bar{\eta}_{i}$ is measurable w.r.t. $t_{i}$, we have $\mathbb{E}[\eta_{1}-\eta_{0}\mid t_{1},t_{0}]=\bar{\eta}_{1}-\bar{\eta}_{0}$. Applying Eq. 20 to the score $\bar{\eta}$ and conditioning on $(t_{1},t_{0})$ gives
$\mathrm{AUC}(\bar{\eta})=\frac{1}{2}+\frac{1}{4\pi_{0}\pi_{1}}\,\mathbb{E}\!\left[(\eta_{1}-\eta_{0})\,\mathrm{sgn}_{0}(\bar{\eta}_{1}-\bar{\eta}_{0})\right]=\frac{1}{2}+\frac{1}{4\pi_{0}\pi_{1}}\,\mathbb{E}\!\left[|\bar{\eta}_{1}-\bar{\eta}_{0}|\right].$ (22)
Furthermore, for any measurable $g:\mathbb{R}\to\mathbb{R}$, the score $g(t)$ is measurable w.r.t. $\sigma(t)$, and by Eq. 20, the conditional expectation given $(t_{1},t_{0})$ is maximized by choosing the sign $\mathrm{sgn}_{0}(g(t_{1})-g(t_{0}))$ equal to $\mathrm{sgn}_{0}(\bar{\eta}_{1}-\bar{\eta}_{0})$. Hence $\mathrm{AUC}(g(t))\leq\mathrm{AUC}(\bar{\eta})$ for all $g$, in particular $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$. Combining Eqs. 21 and 22 and $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$ yields
$\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\ \geq\ \mathrm{AUC}(t^{\star})-\mathrm{AUC}(\bar{\eta})=\frac{1}{4\pi_{0}\pi_{1}}\Big(\mathbb{E}|\eta_{1}-\eta_{0}|-\mathbb{E}|\bar{\eta}_{1}-\bar{\eta}_{0}|\Big).$ (23)
Since $\pi_{0}\pi_{1}\leq 1/4$, we have $(4\pi_{0}\pi_{1})^{-1}\geq 1$, so
$\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\ \geq\ \mathbb{E}|\eta_{1}-\eta_{0}|-\mathbb{E}|\bar{\eta}_{1}-\bar{\eta}_{0}|.$ (24)
We now lower bound the right-hand side by an $L^{2}$ projection error.
Lemma 24 (Gini gap controls $L^{2}$ error).
Let $U$ be a real-valued random variable with support an interval $J=[a,b]$ and density $q$ satisfying $q(u)\geq\underline{q}>0$ for all $u\in J$. Let $\mathcal{F}$ be any sub-$\sigma$-field and set $V=\mathbb{E}[U\mid\mathcal{F}]$. Let $(U,V)$ and $(U^{\prime},V^{\prime})$ be i.i.d. copies. Then
$\mathbb{E}|U-U^{\prime}|-\mathbb{E}|V-V^{\prime}|\ \geq\ \underline{q}\,\mathbb{E}[(U-V)^{2}].$
Proof.
Define $\varphi(r)=\mathbb{E}|U-r|$. Since $U$ has density $q$ supported on $J=[a,b]$, for all $r\in(a,b)$,
$\varphi^{\prime}(r)=\mathbb{E}[-\mathrm{sgn}(U-r)]=2F_{U}(r)-1,\qquad\varphi^{\prime\prime}(r)=2q(r)\geq 2\underline{q},$
so $\varphi$ is $(2\underline{q})$-strongly convex on $J$. Equivalently, for all $u,v\in J$,
$\varphi(u)\ \geq\ \varphi(v)+\varphi^{\prime}(v)(u-v)+\underline{q}\,(u-v)^{2}.$ (25)
Next, since $V=\mathbb{E}[U\mid\mathcal{F}]$, conditional Jensen gives that for every convex $\psi$, $\mathbb{E}[\psi(U)\mid\mathcal{F}]\geq\psi(\mathbb{E}[U\mid\mathcal{F}])=\psi(V)$, hence $\mathbb{E}\psi(U)\geq\mathbb{E}\psi(V)$. Applying this with $\psi(\cdot)=|v-\cdot|$ (convex) shows that for every fixed $v\in J$,
$\mathbb{E}|v-U^{\prime}|\ \geq\ \mathbb{E}|v-V^{\prime}|.$
Taking $v=V$ (which is independent of $(U^{\prime},V^{\prime})$) and then expectation yields
$\mathbb{E}|V-U^{\prime}|\ \geq\ \mathbb{E}|V-V^{\prime}|.$ (26)
Finally, note that $\mathbb{E}|U-U^{\prime}|=\mathbb{E}[\varphi(U)]$ and $\mathbb{E}|V-U^{\prime}|=\mathbb{E}[\varphi(V)]$. Therefore, by Eq. 26,
$\mathbb{E}|U-U^{\prime}|-\mathbb{E}|V-V^{\prime}|\ \geq\ \mathbb{E}|U-U^{\prime}|-\mathbb{E}|V-U^{\prime}|=\mathbb{E}[\varphi(U)-\varphi(V)].$
Apply Eq. 25 with $(u,v)=(U,V)$ and take conditional expectation given $V$. Since $\mathbb{E}[U-V\mid V]=0$, the linear term vanishes and we get $\mathbb{E}[\varphi(U)-\varphi(V)]\geq\underline{q}\,\mathbb{E}[(U-V)^{2}]$, which proves the claim. ∎
Under (H2)–(H5), $\eta(w)=\Psi^{\star}(t^{\star}(w))$ has support an interval $J=[\Psi^{\star}(\underline{t^{\star}}),\Psi^{\star}(\overline{t^{\star}})]$ and admits a density $q$ given by the change-of-variables formula $q(s)=p(r)/(\Psi^{\star})^{\prime}(r)$ where $s=\Psi^{\star}(r)$. Thus $q(s)\geq\underline{p}/\overline{\psi}=\underline{q}$ on $J$. Applying Lemma 24 with $U=\eta(w)$ and $\mathcal{F}=\sigma(t(w))$ gives
$\mathbb{E}|\eta_{1}-\eta_{0}|-\mathbb{E}|\bar{\eta}_{1}-\bar{\eta}_{0}|\ \geq\ \underline{q}\,\mathbb{E}[(\eta-\bar{\eta})^{2}]=\underline{q}\,\inf_{h:\mathbb{R}\to\mathbb{R}}\ \|\eta-h(t)\|^{2}.$
Combining with Eq. 24 yields
$\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\ \geq\ \underline{q}\,\inf_{h:\mathbb{R}\to\mathbb{R}}\ \|\eta-h(t)\|^{2}.$ (27)
We now pass from $\eta$ back to $t^{\star}$. Since $\eta=\Psi^{\star}(t^{\star})$ and $(\Psi^{\star})^{\prime}\geq\underline{\psi}$ on [M,M][-M,M], the inverse $(\Psi^{\star})^{-1}$ exists on $J$ and satisfies
$|(\Psi^{\star})^{-1}(s)-(\Psi^{\star})^{-1}(s^{\prime})|\ \leq\ \underline{\psi}^{-1}\,|s-s^{\prime}|\quad(s,s^{\prime}\in J).$
For any measurable $h:\mathbb{R}\to\mathbb{R}$, define the clipped map $\Pi(h)=\min\{\max\{h,\,\Psi^{\star}(\underline{t^{\star}})\},\,\Psi^{\star}(\overline{t^{\star}})\}$. Then $\|\eta-\Pi(h)(t)\|\leq\|\eta-h(t)\|$, so we may restrict to $h$ with range contained in $J$. For such $h$, set $f=(\Psi^{\star})^{-1}\circ h$. Then pointwise,
$|\eta-h(t)|=|\Psi^{\star}(t^{\star})-\Psi^{\star}(f(t))|\ \geq\ \underline{\psi}\,|t^{\star}-f(t)|,$
hence $\|\eta-h(t)\|^{2}\geq\underline{\psi}^{2}\,\|t^{\star}-f(t)\|^{2}$. Taking infima implies
$\inf_{h:\mathbb{R}\to\mathbb{R}}\ \|\eta-h(t)\|^{2}\ \geq\ \underline{\psi}^{2}\ \inf_{f:\mathbb{R}\to\mathbb{R}}\ \|t^{\star}-f(t)\|^{2}.$
Plugging into Eq. 27 and using (H6) yields the global inequality
$\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\ \geq\ \lambda_{0}\ \inf_{a\in\mathbb{R}}\ \|t^{\star}-at\|^{2},\qquad\lambda_{0}=c_{\rm MSE}\,\underline{q}\,\underline{\psi}^{2}.$ (28)
Small AUC-regret forces the optimal scaling to be nonnegative. Write
$d(t)^{2}=\inf_{a\in\mathbb{R}}\ \|t^{\star}-at\|^{2}.$
Let $\hat{a}$ be a minimizer (it exists because $a\mapsto\|t^{\star}-at\|^{2}$ is a quadratic polynomial). Assume for contradiction that $\hat{a}<0$ and set $s=(-\hat{a})\,t$. Then $s$ is a positive multiple of $t$, so $\mathrm{AUC}(s)=\mathrm{AUC}(t)$, and $d(t)^{2}=\|t^{\star}+s\|^{2}$.
We need a continuity bound comparing $\mathrm{AUC}(s)$ and $\mathrm{AUC}(-t^{\star})$. First, under (H2)–(H4), the conditional density of $r=t^{\star}(w)$ given $z=1$ is $p_{1}(r)=\eta(r)p(r)/\pi_{1}$, hence
$\|p_{1}\|_{\infty}\leq\frac{1-\varepsilon}{\varepsilon}\,\overline{p}=B_{0}.$
Similarly, the conditional density $p_{0}$ of $t^{\star}(w)$ given $z=0$ satisfies $\|p_{0}\|_{\infty}\leq B_{0}$. Let $r^{+}\sim p_{1}$ and $r^{-}\sim p_{0}$ be independent; then $r^{+}-r^{-}$ has a density bounded by $B_{0}$, hence
$\mathbb{P}\big(|r^{+}-r^{-}|\leq u\big)\ \leq\ 2B_{0}\,u,\qquad u\geq 0.$ (29)
Lemma 25 (Hölder continuity of AUC in $L^{2}$).
Assume Eq. 29 holds for $r^{+}-r^{-}$ with constant $B_{0}$. Then for any measurable scores $r,s:\mathcal{W}\to[-M,M]$,
$|\mathrm{AUC}(r)-\mathrm{AUC}(s)|\ \leq\ C_{H}\ \|r-s\|^{2/3},\qquad C_{H}=6\,\varepsilon^{-1/3}\,B_{0}^{2/3}.$
Proof.
Fix $u>0$ and let $w^{+},w^{-}$ be independent draws from $w\mid z=1$ and $w\mid z=0$, respectively. If $|r(w^{+})-r(w^{-})|>2u$ and $|s(w^{\pm})-r(w^{\pm})|\leq u$ for both $\pm$, then
$|(s(w^{+})-s(w^{-}))-(r(w^{+})-r(w^{-}))|\leq|s(w^{+})-r(w^{+})|+|s(w^{-})-r(w^{-})|\leq 2u,$
so $\mathrm{sgn}_{0}(s(w^{+})-s(w^{-}))=\mathrm{sgn}_{0}(r(w^{+})-r(w^{-}))$ and hence $\phi(s(w^{+})-s(w^{-}))=\phi(r(w^{+})-r(w^{-}))$. Therefore,
$|\mathrm{AUC}(s)-\mathrm{AUC}(r)|\leq\mathbb{P}\big(|r(w^{+})-r(w^{-})|\leq 2u\big)+\mathbb{P}\big(|(s-r)(w^{+})|>u\big)+\mathbb{P}\big(|(s-r)(w^{-})|>u\big).$
The first term is bounded by $4B_{0}u$ using Eq. 29. For the second term, Markov’s inequality and $\mathbb{P}(z=1)=\pi_{1}\geq\varepsilon$ give
$\mathbb{P}\big(|(s-r)(w^{+})|>u\big)\leq\frac{\mathbb{E}[(s-r)(w^{+})^{2}]}{u^{2}}=\frac{\mathbb{E}[(s-r)(w)^{2}\,\mathbb{I}\left[z=1\right]]}{\pi_{1}u^{2}}\leq\frac{\|s-r\|^{2}}{\varepsilon u^{2}},$
and similarly $\mathbb{P}(|(s-r)(w^{-})|>u)\leq\|s-r\|^{2}/(\varepsilon u^{2})$. Thus
$|\mathrm{AUC}(s)-\mathrm{AUC}(r)|\leq 4B_{0}u+\frac{2}{\varepsilon}\frac{\|s-r\|^{2}}{u^{2}}.$
Optimizing over $u>0$ yields $|\mathrm{AUC}(s)-\mathrm{AUC}(r)|\leq 6\,\varepsilon^{-1/3}B_{0}^{2/3}\,\|s-r\|^{2/3}$. ∎
Apply Lemma 25 with $r=-t^{\star}$ and $s$ as above. Since $t^{\star}(w^{+})-(-t^{\star}(w^{-}))=-(r^{+}-r^{-})$, the margin bound Eq. 29 applies, hence
$\mathrm{AUC}(t)=\mathrm{AUC}(s)\ \leq\ \mathrm{AUC}(-t^{\star})+C_{H}\,\|s-(-t^{\star})\|^{2/3}=\mathrm{AUC}(-t^{\star})+C_{H}\,d(t)^{2/3}.$ (30)
By Eq. 28, $d(t)^{2}\leq(\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t))/\lambda_{0}$. Let $\Delta=\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)$; then $d(t)^{2/3}\leq(\Delta/\lambda_{0})^{1/3}$. Using $\mathrm{AUC}(-t^{\star})=1-\mathrm{AUC}(t^{\star})$ (order reversal) in Eq. 30 gives
$\mathrm{AUC}(t)\leq 1-\mathrm{AUC}(t^{\star})+C_{H}\Big(\frac{\Delta}{\lambda_{0}}\Big)^{1/3}.$
Equivalently,
$\mathrm{AUC}(t^{\star})-\Delta\ \leq\ 1-\mathrm{AUC}(t^{\star})+C_{H}\Big(\frac{\Delta}{\lambda_{0}}\Big)^{1/3}.$ (31)
We now lower bound the separation $\mathrm{AUC}(t^{\star})-\tfrac{1}{2}$ using only the constants in the assumptions. Let $s=\eta(w)=\Psi^{\star}(t^{\star}(w))$. Since $p(r)\leq\overline{p}$ on $I$ and integrates to 11, we have $|I|=\overline{t^{\star}}-\underline{t^{\star}}\geq 1/\overline{p}$. Since $(\Psi^{\star})^{\prime}\geq\underline{\psi}$, the support $J=[\Psi^{\star}(\underline{t^{\star}}),\Psi^{\star}(\overline{t^{\star}})]$ has length at least
$|J|\geq\underline{\psi}\,|I|\ \geq\ \frac{\underline{\psi}}{\overline{p}}.$
Also $s$ has density $q\geq\underline{q}$ on $J$. Let $J_{L}$ and $J_{R}$ be the leftmost and rightmost subintervals of $J$ of length $|J|/4$. Then $\mathbb{P}(s\in J_{L})\geq\underline{q}\,|J|/4$ and similarly for $J_{R}$. Hence, with $s_{0},s_{1}$ i.i.d. copies of $s$,
$\mathbb{E}|s_{1}-s_{0}|\ \geq\ \frac{|J|}{2}\,\mathbb{P}(s_{0}\in J_{L},s_{1}\in J_{R})+\frac{|J|}{2}\,\mathbb{P}(s_{0}\in J_{R},s_{1}\in J_{L})\ \geq\ \frac{|J|}{2}\cdot 2\Big(\frac{\underline{q}\,|J|}{4}\Big)^{2}\ =\ \frac{1}{16}\,\underline{q}^{2}\,|J|^{3}\ \geq\ \gamma_{0}.$
By Eq. 21 and $(4\pi_{0}\pi_{1})^{-1}\geq 1$, we get
$\mathrm{AUC}(t^{\star})-\frac{1}{2}=\frac{1}{4\pi_{0}\pi_{1}}\mathbb{E}|\eta_{1}-\eta_{0}|\ \geq\ \mathbb{E}|\eta_{1}-\eta_{0}|\ \geq\ \gamma_{0}.$
Therefore $\mathrm{AUC}(t^{\star})\geq\tfrac{1}{2}+\gamma_{0}$ and $\mathrm{AUC}(-t^{\star})=1-\mathrm{AUC}(t^{\star})\leq\tfrac{1}{2}-\gamma_{0}$.
Now assume $\Delta\leq\epsilon_{0}$. Then $\mathrm{AUC}(t)=\mathrm{AUC}(t^{\star})-\Delta\geq\tfrac{1}{2}+\gamma_{0}-\epsilon_{0}\geq\tfrac{1}{2}+\gamma_{0}/2$. On the other hand, Eq. 31 and $\Delta\leq\epsilon_{0}\leq\lambda_{0}(\gamma_{0}/(2C_{H}))^{3}$ give
$\mathrm{AUC}(t)\leq\mathrm{AUC}(-t^{\star})+C_{H}\Big(\frac{\Delta}{\lambda_{0}}\Big)^{1/3}\leq\Big(\tfrac{1}{2}-\gamma_{0}\Big)+\frac{\gamma_{0}}{2}=\tfrac{1}{2}-\frac{\gamma_{0}}{2},$
a contradiction. Hence $\hat{a}\geq 0$ whenever $\Delta\leq\epsilon_{0}$, and thus
$\inf_{a\in\mathbb{R}}\ \|t^{\star}-at\|^{2}=\inf_{a\geq 0}\ \|t^{\star}-at\|^{2}\qquad\text{for all }t\in\mathcal{T}\text{ with }\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\leq\epsilon_{0}.$
To complete the theorem, combine the above with Eq. 28. ∎
E.3 Proof of Lemma 3
Proof.
Define $R_{t}=\{(w,w^{\prime}):\ t(w)-t(w^{\prime})\geq 0\}$, $S\triangle S^{\prime}=(S\backslash S^{\prime})\cup(S^{\prime}\backslash S)$, and $d_{\triangle}(S,S^{\prime})=P^{2}(S\triangle S^{\prime})$ for $S,S^{\prime}\subset(\mathcal{W})^{2}$. We will show that for all $t,s\in\mathcal{T}$, we have $d_{\triangle}(R_{t},R_{s})\leq c_{q}\,\|t-s\|_{q}^{\beta}$ for $c_{q}=2^{\,1+\beta}\,c_{\rm margin}^{\,q/(\alpha+q)}$.
Fix $t,s\in\mathcal{T}$ and write $\delta=t-s$. Let $D_{t}=t(w)-t(w^{\prime})$ and $D_{s}=s(w)-s(w^{\prime})$. On the symmetric difference $R_{t}\triangle R_{s}$, the quantities $D_{t}$ and $D_{s}$ have opposite negativity/nonnegativity, hence $|D_{t}|\leq|D_{t}-D_{s}|$. Since
$D_{t}-D_{s}\;=\;(t-s)(w)-(t-s)(w^{\prime})\;=\;\delta(w)-\delta(w^{\prime}),$
we have the inclusion
$R_{t}\triangle R_{s}\ \subseteq\ \bigl\{|D_{t}|\leq|\delta(w)-\delta(w^{\prime})|\bigr\}.$
Therefore, for any $\lambda>0$,
$R_{t}\triangle R_{s}\subseteq\bigl\{|D_{t}|\leq\lambda\bigr\}\ \cup\ \bigl\{|\delta(w)-\delta(w^{\prime})|>\lambda\bigr\},$
and thus
$d_{\triangle}(R_{t},R_{s})\leq P^{2}(|D_{t}|\leq\lambda)\;+\;P^{2}(|\delta(w)-\delta(w^{\prime})|>\lambda).$ (32)
The assumed margin condition gives $P^{2}(|D_{t}|\leq\lambda)\leq c_{\rm margin}\lambda^{\alpha}$. For the second term, Markov’s inequality and $|a-b|^{q}\leq 2^{q-1}(|a|^{q}+|b|^{q})$ yield
$P^{2}(|\delta(w)-\delta(w^{\prime})|>\lambda)\leq\frac{\mathbb{E}|\delta(w)-\delta(w^{\prime})|^{q}}{\lambda^{q}}\leq\frac{2^{q}\,\mathbb{E}|\delta(w)|^{q}}{\lambda^{q}}=\frac{2^{q}\,\|t-s\|_{q}^{q}}{\lambda^{q}}.$
Plugging into Eq. 32 gives
$d_{\triangle}(R_{t},R_{s})\leq c_{\rm margin}\lambda^{\alpha}+\frac{2^{q}\|t-s\|_{q}^{q}}{\lambda^{q}}.$
Choose $\lambda=\bigl(2^{q}\|t-s\|_{q}^{q}/c_{\rm margin}\bigr)^{1/(\alpha+q)}$, which equalizes the two terms, to obtain $d_{\triangle}(R_{t},R_{s})\leq 2\,c_{\rm margin}^{q/(\alpha+q)}\,(2^{q}\|t-s\|_{q}^{q})^{\alpha/(\alpha+q)}.$
E.4 Proof of Theorem 8
To prove Theorem 8, we first consider the implication of theorem 5 of Clémençon et al. [2008] under symmetric difference entropy conditions in order obtain rates beyond the case of VC classes. Then we show that a Tsybakov-style margin condition implies their generic noise condition. Then we use these building blocks to prove Theorem 8.
E.5 Guarantees for ranking empirical risk optimization under symmetric-difference entropy conditions
For this subsection, we abandon all previously defined notation and conventions and adopt the notation of Clémençon et al. [2008].
E.5.1 Setup and notation (following Clémençon et al., 2008)
Let $(X,Y)$ be a random pair taking values in $\mathcal{X}\times\mathbb{R}$, and let $(X^{\prime},Y^{\prime})$ be an independent copy. Write $Z=(Y-Y^{\prime})/2$. A (pairwise) ranking rule is a measurable function $r:\mathcal{X}\times\mathcal{X}\to\{-1,1\}$. Throughout we consider symmetric ranking rules (as in the paper): $r(x,x^{\prime})=-r(x^{\prime},x)$, and extend the co-domain just for identical inputs in order to set $r(x,x)=0$.
The ranking risk is
$L(r)=\mathbb{P}\{Z\,r(X,X^{\prime})<0\}.$
Let $r^{\star}$ denote a risk minimizer over all measurable ranking rules and $L^{\star}=L(r^{\star})$.
Given i.i.d. training data $D_{n}=(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$, define the empirical ranking risk (a $U$-statistic)
$L_{n}(r)=\frac{1}{n(n-1)}\sum_{i\neq j}\mathbb{I}\left[Z_{i,j}\,r(X_{i},X_{j})<0\right],\qquad Z_{i,j}=(Y_{i}-Y_{j})/2,$
and an empirical risk minimizer over $\mathcal{R}$ by
$r_{n}\in\arg\min_{r\in\mathcal{R}}L_{n}(r).$
In Section 4 of the paper the excess risk $\Lambda(r)=L(r)-L^{\star}$ is written as a $U$-statistic with kernel
$q_{r}\big((x,y),(x^{\prime},y^{\prime})\big)=\mathbb{I}\left[(y-y^{\prime})\,r(x,x^{\prime})<0\right]-\mathbb{I}\left[(y-y^{\prime})\,r^{\star}(x,x^{\prime})<0\right],$
so that $\Lambda(r)=\mathbb{E}\,q_{r}\big((X,Y),(X^{\prime},Y^{\prime})\big)$ and its empirical estimate $\Lambda_{n}(r)$ admits the Hoeffding decomposition
$\Lambda_{n}(r)-\Lambda(r)=2T_{n}(r)+W_{n}(r),$
where $T_{n}(r)=(1/n)\sum_{i=1}^{n}h_{r}(X_{i},Y_{i})$ and $W_{n}(r)$ is a degenerate $U$-statistic with kernel $\widetilde{h}_{r}$ (see the paper for the exact definitions).
Assumption 13 (Assumption 4 of Clémençon et al. [2008]).
There exist constants $c>0$ and $\alpha\in[0,1]$ such that for all $r\in\mathcal{R}$,
$\mathrm{Var}\big(h_{r}(X,Y)\big)\leq c\,\Lambda(r)^{\alpha}.$
We next restate theorem 5 of Clémençon et al. [2008] for reference. Let $\varepsilon_{1},\dots,\varepsilon_{n}$ be i.i.d. Rademacher signs independent of the data. Define the three (random) complexity quantities
$Z_{\varepsilon}=\sup_{r\in\mathcal{R}}\sum_{i,j}\varepsilon_{i}\varepsilon_{j}\,\widetilde{h}_{r}\big((X_{i},Y_{i}),(X_{j},Y_{j})\big),$
$U_{\varepsilon}=\sup_{r\in\mathcal{R}}\sup_{\|\beta\|_{2}\leq 1}\sum_{i,j}\varepsilon_{i}\beta_{j}\,\widetilde{h}_{r}\big((X_{i},Y_{i}),(X_{j},Y_{j})\big),$
$M=\sup_{r\in\mathcal{R}}\sup_{1\leq k\leq n}\sum_{i=1}^{n}\varepsilon_{i}\,\widetilde{h}_{r}\big((X_{i},Y_{i}),(X_{k},Y_{k})\big).$
Also define the centered empirical process $\nu_{n}$ based on the “loss” $\ell$ in the paper:
$\ell(r,(x,y))=2\,\mathbb{E}\big[\mathbb{I}\left[(y-Y)\,r(x,X)<0\right]\big]-L(r),\qquad\nu_{n}(r)=\frac{1}{n}\sum_{i=1}^{n}\ell(r,(X_{i},Y_{i}))-L(r).$
Finally define the pseudo-distance
$d(r,r^{\prime})=\Big(\mathbb{E}\big[\mathbb{E}\big[\mathbb{I}\left[r(X,X^{\prime})\neq r^{\prime}(X,X^{\prime})\right]\mid X\big]^{2}\big]\Big)^{1/2}.$
Let $\phi$ be any nondecreasing function such that $\phi(x)/x$ is nonincreasing, $\phi(1)\geq 1$, and
$\sqrt{n}\,\mathbb{E}\sup_{r^{\prime}\in\mathcal{R}:\,d(r,r^{\prime})\leq\sigma}\big|\nu_{n}(r)-\nu_{n}(r^{\prime})\big|\leq\phi(\sigma)\quad\text{for all }r\in\mathcal{R},\;\sigma>0.$
Let $\rho>0$ be the unique solution of $\sqrt{n}\,\rho^{2}=\phi(\rho^{\alpha})$.
Theorem 10 (Theorem 5 of Clémençon et al. [2008]).
Assume Assumption 13 holds. Then there exists a universal constant $C$ such that for all $\delta\in(0,1)$, with probability at least $1-\delta$,
$L(r_{n})-L^{\star}\leq 2\inf_{r\in\mathcal{R}}\big(L(r)-L^{\star}\big)+C\left(\frac{\mathbb{E}Z_{\varepsilon}}{n^{2}}+\frac{\mathbb{E}U_{\varepsilon}}{n^{2}}\sqrt{\log\frac{1}{\delta}}+\frac{\mathbb{E}M}{n^{2}}\log\frac{1}{\delta}+\frac{1}{n}\log\frac{1}{\delta}+\rho^{2}\log\frac{1}{\delta}\right).$
E.5.2 Symmetric-difference covering numbers
For a ranker $r\in\mathcal{R}$, define the set $A_{r}=\{(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}:\ r(x,x^{\prime})=1\}$. Define the symmetric-difference pseudo-metric
$\Delta(r,r^{\prime})=\mathbb{P}\big(r(X,X^{\prime})\neq r^{\prime}(X,X^{\prime})\big)=\mathbb{P}\big((X,X^{\prime})\in A_{r}\triangle A_{r^{\prime}}\big).$
For $\varepsilon\in(0,1]$, define the symmetric-difference covering number $N_{\triangle}(\varepsilon,\mathcal{R})$ as the smallest $m$ such that there exist sets $S_{1},\dots,S_{m}\subseteq\mathcal{X}\times\mathcal{X}$ with the property that for each $r\in\mathcal{R}$ there exists $j$ such that $\mathbb{P}(A_{r}\triangle S_{j})\leq\varepsilon$. (Notice that the slightly different notation for $N_{\triangle}$ compared to the main text, since here $\mathcal{R}$ is a class of indicators.) Equivalently (up to the identification $r\leftrightarrow A_{r}$), $N_{\triangle}(\varepsilon,\mathcal{R})$ is the $\varepsilon$-covering number of $\mathcal{R}$ under the pseudo-metric $\Delta$.
We will consider either of the following entropy assumptions.
E.5.3 From symmetric-difference entropy to the quantities in Theorem 10
We need to control the expectations $\mathbb{E}Z_{\varepsilon}$, $\mathbb{E}U_{\varepsilon}$, $\mathbb{E}M$ and the fixed point $\rho$. The following reductions are elementary and isolate the role of the entropy.
Lemma 26 (Comparing $\Delta$ and $d$).
For all $r,r^{\prime}\in\mathcal{R}$ we have $d(r,r^{\prime})^{2}\leq\Delta(r,r^{\prime})$. Consequently, for all $\sigma\in(0,1]$,
$N\big(\sigma,\mathcal{R},d\big)\leq N_{\triangle}(\sigma^{2},\mathcal{R}),\qquad\log N(\sigma,\mathcal{R},d)\leq\log N_{\triangle}(\sigma^{2},\mathcal{R}),$
where $N(\sigma,\mathcal{R},d)$ is the $\sigma$-covering number with respect to the pseudo-metric $d$.
Proof.
Let $\Delta_{X}(r,r^{\prime})=\mathbb{P}\big(r(X,X^{\prime})\neq r^{\prime}(X,X^{\prime})\mid X\big)$. Then $d(r,r^{\prime})^{2}=\mathbb{E}[\Delta_{X}(r,r^{\prime})^{2}]\leq\mathbb{E}[\Delta_{X}(r,r^{\prime})]=\Delta(r,r^{\prime})$ by Jensen’s inequality. The covering-number inequality follows: if $d(r,r^{\prime})\leq\sigma$ then $\Delta(r,r^{\prime})\leq\sigma^{2}$. ∎
Lemma 27 (Lipschitz control of the degenerate kernels).
There is a universal constant $B_{0}<\infty$ such that for all $r,r^{\prime}\in\mathcal{R}$:
  1. (i)
    $\|\widetilde{h}_{r}\|_{\infty}\leq B_{0}$,
  2. (ii)
    $\|\widetilde{h}_{r}-\widetilde{h}_{r^{\prime}}\|_{L_{2}(\mathbb{P}_{(X,Y)}^{2})}\leq B_{0}\,\Delta(r,r^{\prime})^{1/2}$.
Consequently, the $L_{2}(\mathbb{P}_{(X,Y)}^{2})$ covering numbers of the kernel class $\widetilde{\mathcal{H}}=\{\widetilde{h}_{r}:\ r\in\mathcal{R}\}$ satisfy
$\log N_{2}(u,\widetilde{\mathcal{H}})\;\leq\;\log N_{\triangle}\!\left(\frac{u^{2}}{B_{0}^{2}},\mathcal{R}\right),\qquad u\in(0,B_{0}],$
where $N_{2}$ denotes covering numbers in $L_{2}(\mathbb{P}_{(X,Y)}^{2})$.
Proof.
We use only the definitions in Section 4 of Clémençon et al. [2008]. Since $q_{r}$ is a difference of indicators, $\|q_{r}\|_{\infty}\leq 1$. By definition,
$\widetilde{h}_{r}=q_{r}-\Lambda(r)-h_{r}(\cdot)-h_{r}(\cdot),$
so $\|\widetilde{h}_{r}\|_{\infty}\leq 1+|\Lambda(r)|+2\|h_{r}\|_{\infty}\leq 4$ because $|\Lambda(r)|\leq 1$ and $\|h_{r}\|_{\infty}\leq 1$. Thus (i) holds with (say) $B_{0}=4$.
For (ii), observe that for any two rankers $r,r^{\prime}$,
$\big|q_{r}\big((x,y),(x^{\prime},y^{\prime})\big)-q_{r^{\prime}}\big((x,y),(x^{\prime},y^{\prime})\big)\big|\leq\mathbb{I}\left[r(x,x^{\prime})\neq r^{\prime}(x,x^{\prime})\right].$
Therefore,
$\|q_{r}-q_{r^{\prime}}\|_{L_{2}(\mathbb{P}_{(X,Y)}^{2})}^{2}\leq\mathbb{P}\big(r(X,X^{\prime})\neq r^{\prime}(X,X^{\prime})\big)=\Delta(r,r^{\prime}).$
Next, $h_{r}(x,y)=\mathbb{E}[q_{r}((x,y),(X^{\prime},Y^{\prime}))]-\Lambda(r)$, hence
$|h_{r}(x,y)-h_{r^{\prime}}(x,y)|\leq\mathbb{E}\big|q_{r}((x,y),(X^{\prime},Y^{\prime}))-q_{r^{\prime}}((x,y),(X^{\prime},Y^{\prime}))\big|+|\Lambda(r)-\Lambda(r^{\prime})|\leq\Delta_{x}(r,r^{\prime})+\Delta(r,r^{\prime}),$
where $\Delta_{x}(r,r^{\prime})=\mathbb{P}(r(x,X^{\prime})\neq r^{\prime}(x,X^{\prime}))$. Taking $L_{2}(\mathbb{P}_{(X,Y)})$ norms and using $\Delta(r,r^{\prime})\leq\Delta(r,r^{\prime})^{1/2}$ (since $\Delta\leq 1$) gives
$\|h_{r}-h_{r^{\prime}}\|_{L_{2}(\mathbb{P}_{(X,Y)})}\leq\|\Delta_{X}(r,r^{\prime})\|_{L_{2}(\mathbb{P}_{X})}+\Delta(r,r^{\prime})^{1/2}=d(r,r^{\prime})+\Delta(r,r^{\prime})^{1/2}\leq 2\Delta(r,r^{\prime})^{1/2},$
where we used Lemma 26 in the last step. Finally, by the triangle inequality,
$\|\widetilde{h}_{r}-\widetilde{h}_{r^{\prime}}\|_{2}\leq\|q_{r}-q_{r^{\prime}}\|_{2}+|\Lambda(r)-\Lambda(r^{\prime})|+2\|h_{r}-h_{r^{\prime}}\|_{2}\leq\Delta^{1/2}+\Delta^{1/2}+4\Delta^{1/2}\leq 6\Delta^{1/2}.$
Thus (ii) holds for $B_{0}=6$, and we may keep $B_{0}=6$ for the covering-number statement. ∎
E.5.4 A complexity proxy
For $n\geq 2$, define the entropy integral proxy
$\mathcal{J}_{n}=\inf_{0<\eta\leq 1}\left\{\eta+\frac{1}{\sqrt{n}}\int_{\eta}^{1}\sqrt{\log N_{\triangle}(u^{2},\mathcal{R})}\,du\right\}.$ (33)
(Any other fixed upper limit in $(0,\infty)$ is equivalent up to constants; the choice $1$ is only for notational convenience.)
Lemma 28 (Controlling $\mathbb{E}Z_{\varepsilon},\mathbb{E}U_{\varepsilon},\mathbb{E}M$ via entropy).
There exists a universal constant $C_{0}<\infty$ such that
$\mathbb{E}Z_{\varepsilon}\leq C_{0}\,n^{2}\,\mathcal{J}_{n}^{2},\qquad\mathbb{E}U_{\varepsilon}\leq C_{0}\,n^{3/2}\,\mathcal{J}_{n},\qquad\mathbb{E}M\leq C_{0}\,n\,\mathcal{J}_{n}.$
Proof.
[…]
$I\;=\;\int_{0}^{B_{0}}\sqrt{\log N_{2}(u)}\,du.$
By Lemma 27, $\log N_{2}(u)\leq\log N_{\triangle}(u^{2}/B_{0}^{2},\mathcal{R})$, and therefore $I\lesssim\sqrt{n}\,\mathcal{J}_{n}$ (up to enlarging the universal constants).
Bounding $\mathbb{E}Z_{\varepsilon}$. The random variable $Z_{\varepsilon}$ is an order-$2$ Rademacher chaos indexed by the class of bounded kernels $\widetilde{\mathcal{H}}$. A maximal inequality for such chaoses (see, e.g., de la Peña and Giné, 1999, Ch. 3 together with the entropy-integral bound for chaoses in Arcones and Giné, 1993, Prop. 2.2 and Prop. 2.6) yields a universal constant $C$ such that
$\mathbb{E}Z_{\varepsilon}\;\leq\;C\,n\,I^{2}.$
Using $I\lesssim\sqrt{n}\,\mathcal{J}_{n}$ gives $\mathbb{E}Z_{\varepsilon}\lesssim n^{2}\mathcal{J}_{n}^{2}$.
Bounding $\mathbb{E}U_{\varepsilon}$. Likewise, $U_{\varepsilon}$ is the (Hilbertian) Rademacher functional associated to $\widetilde{\mathcal{H}}$. The corresponding entropy-integral maximal inequality (again, e.g. Arcones and Giné, 1993, Prop. 2.6; see also de la Peña and Giné, 1999, Ch. 3) gives
$\mathbb{E}U_{\varepsilon}\;\leq\;C\,n\,I\;\lesssim\;n^{3/2}\,\mathcal{J}_{n}.$
Bounding $\mathbb{E}M$. For $M$ one uses the (ordinary) Dudley entropy-integral bound for Rademacher averages [Boucheron et al., 2013, Sec. 13.3]. Together with Lemma 27 (to transfer symmetric-difference entropy of $\mathcal{R}$ to $L_{2}$ entropy of the sections of $\widetilde{\mathcal{H}}$), this yields
$\mathbb{E}M\;\leq\;C\,\sqrt{n}\,I\;\lesssim\;n\,\mathcal{J}_{n}.$
Collecting the three estimates proves the lemma. ∎
E.5.5 A usable bound for the modulus function $\phi$ and the fixed point $\rho$
We will use the following symmetrization + chaining estimate for the ordinary empirical process $\nu_{n}$. Let $\mathcal{L}=\{\ell(r,\cdot):r\in\mathcal{R}\}$ be the loss class in the notation of Theorem 10.
Lemma 29 (A choice of $\phi$ from metric entropy).
There exists a universal constant $C_{1}<\infty$ such that for every $n\geq 2$, one may take
$\phi(\sigma)=C_{1}\inf_{0<\eta\leq\sigma}\left\{\eta\sqrt{\log N(\eta,\mathcal{R},d)}+\int_{\eta}^{\sigma}\sqrt{\log N(u,\mathcal{R},d)}\,du\right\},\qquad\sigma\in(0,1],$
so that the condition in Theorem 10 holds. In particular, using Lemma 26, one has the crude upper bound
$\phi(\sigma)\leq C_{1}\inf_{0<\eta\leq\sigma}\left\{\eta\sqrt{\log N_{\triangle}(\eta^{2},\mathcal{R})}+\int_{\eta}^{\sigma}\sqrt{\log N_{\triangle}(u^{2},\mathcal{R})}\,du\right\}.$
Proof.
Fix $r\in\mathcal{R}$ and consider the increment class $\mathcal{F}_{r,\sigma}=\{\ell(r,\cdot)-\ell(r^{\prime},\cdot):\ r^{\prime}\in\mathcal{R},\ d(r,r^{\prime})\leq\sigma\}$. Then
$\nu_{n}(r)-\nu_{n}(r^{\prime})=(\mathbb{P}_{n}-\mathbb{P})\big(\ell(r,\cdot)-\ell(r^{\prime},\cdot)\big),$
so by symmetrization,
$\mathbb{E}\sup_{d(r,r^{\prime})\leq\sigma}|\nu_{n}(r)-\nu_{n}(r^{\prime})|\leq 2\mathbb{E}\sup_{f\in\mathcal{F}_{r,\sigma}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i},Y_{i})\right|.$
By standard chaining/Dudley bounds for Rademacher averages indexed by a class of bounded functions (e.g. [Boucheron et al., 2013, Theorem 13.5]), for a universal constant $C_{1}$,
$\mathbb{E}\sup_{f\in\mathcal{F}_{r,\sigma}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i},Y_{i})\right|\leq\frac{C_{1}}{\sqrt{n}}\inf_{0<\eta\leq\sigma}\left\{\eta\sqrt{\log N_{2}(\eta,\mathcal{F}_{r,\sigma})}+\int_{\eta}^{\sigma}\sqrt{\log N_{2}(u,\mathcal{F}_{r,\sigma})}\,du\right\}.$
Finally, the map $r^{\prime}\mapsto\ell(r,\cdot)-\ell(r^{\prime},\cdot)$ is Lipschitz in $d$ (in $L_{2}(\mathbb{P})$) up to a universal constant, so $N_{2}(u,\mathcal{F}_{r,\sigma})\leq N(u,\mathcal{R},d)$ for all $u\leq\sigma$, again up to constants. Absorbing constants into $C_{1}$ yields the displayed formula for $\phi$. The last displayed inequality follows from Lemma 26. ∎
E.5.6 Ranking suboptimality bound under symmetric-difference entropy
We now state and prove the instantiation of theorem 5 of Assumption 13.
Theorem 11 (Entropy-based instantiation of Theorem 10).
Assume Assumption 13 holds and let $r_{n}$ be an empirical minimizer of $L_{n}(r)$ over $\mathcal{R}$. There exists a constant $C<\infty$ depending only on the universal constants in Theorem 10 and on the constant $c$ from Assumption 13 such that, for every $\delta\in(0,1)$, with probability at least $1-\delta$,
$L(r_{n})-L^{\star}\leq 2\inf_{r\in\mathcal{R}}\big(L(r)-L^{\star}\big)+C\,\mathfrak{R}_{n,\delta},$
where $\mathfrak{R}_{n,\delta}$ can be taken as follows.
Case 1: Assume $\log N_{\triangle}(\varepsilon,\mathcal{R})\leq A\varepsilon^{-p}$ for some $A>0$ and $p>0$. Then:
  1. $(1a)$
    If $0<p<1$, then
    $\mathfrak{R}_{n,\delta}=\left(\frac{A}{n}\right)^{\!\frac{1}{2-\alpha(1-p)}}\log\frac{1}{\delta}+\frac{A+\log(1/\delta)}{n}.$
  2. $(1b)$
    If $p=1$, then
    $\mathfrak{R}_{n,\delta}=\frac{\sqrt{A}\,\log n}{\sqrt{n}}\,\log\frac{1}{\delta}+\frac{A\log^{2}n+\log(1/\delta)}{n}.$
  3. $(1c)$
    If $p>1$, then
    $\mathfrak{R}_{n,\delta}=A^{\frac{1}{2p}}\,n^{-\frac{1}{2p}}\log\frac{1}{\delta}+A^{\frac{1}{p}}\,n^{-\frac{1}{p}}+\frac{\log(1/\delta)}{n}.$
Case 2: Assume $N_{\triangle}(\varepsilon,\mathcal{R})\leq(A/\varepsilon)^{v}$ for some $A\geq 1$ and $v\geq 1$. Then
$\mathfrak{R}_{n,\delta}=\left(\frac{v\,\log(An)}{n}\right)^{\!\frac{1}{2-\alpha}}\,\log\frac{1}{\delta}+\frac{v+\log(1/\delta)}{n}.$
Proof.
We apply Theorem 10. By Lemma 28,
$\frac{\mathbb{E}Z_{\varepsilon}}{n^{2}}\leq C_{0}\mathcal{J}_{n}^{2},\qquad\frac{\mathbb{E}U_{\varepsilon}}{n^{2}}\sqrt{\log\frac{1}{\delta}}\leq C_{0}\frac{\mathcal{J}_{n}}{\sqrt{n}}\sqrt{\log\frac{1}{\delta}},\qquad\frac{\mathbb{E}M}{n^{2}}\log\frac{1}{\delta}\leq C_{0}\frac{\mathcal{J}_{n}}{n}\log\frac{1}{\delta}.$
Since $\mathcal{J}_{n}\leq 1$ (take $\eta=1$ in Eq. 33), the $M$-term is always bounded by $C_{0}(\log(1/\delta))/n$ and is absorbed by the explicit $(\log(1/\delta))/n$ term in Theorem 10. Likewise, by $ab\leq(a^{2}+b^{2})/2$,
$\frac{\mathcal{J}_{n}}{\sqrt{n}}\sqrt{\log\frac{1}{\delta}}\leq\frac{1}{2}\mathcal{J}_{n}^{2}+\frac{1}{2}\frac{\log(1/\delta)}{n}.$
Thus Theorem 10 implies, with probability at least $1-\delta$,
$L(r_{n})-L^{\star}\leq 2\inf_{r\in\mathcal{R}}\big(L(r)-L^{\star}\big)+C\left(\mathcal{J}_{n}^{2}+\frac{\log(1/\delta)}{n}+\rho^{2}\log\frac{1}{\delta}\right),$ (34)
for a possibly enlarged universal constant $C$.
It remains to control $\mathcal{J}_{n}$ and $\rho$ in each entropy regime.
Bounding $\mathcal{J}_{n}$.
Under Case 1, we have $\log N_{\triangle}(u^{2},\mathcal{R})\leq Au^{-2p}$ for $u\in(0,1]$, hence
$\mathcal{J}_{n}\leq\inf_{0<\eta\leq 1}\left\{\eta+\frac{\sqrt{A}}{\sqrt{n}}\int_{\eta}^{1}u^{-p}\,du\right\}.$
If $0<p<1$, the integral is bounded by $(1-p)^{-1}$, so $\mathcal{J}_{n}^{2}\lesssim A/n$. If $p=1$, the integral equals $\log(1/\eta)$ and optimizing at $\eta\asymp\sqrt{A}/\sqrt{n}$ yields $\mathcal{J}_{n}\lesssim(\sqrt{A}/\sqrt{n})\log n$ and thus $\mathcal{J}_{n}^{2}\lesssim A\log^{2}n/n$. If $p>1$, the integral behaves like $\eta^{1-p}$ and optimizing at $\eta\asymp(\sqrt{A}/\sqrt{n})^{1/p}$ yields $\mathcal{J}_{n}\lesssim A^{1/(2p)}n^{-1/(2p)}$ and thus $\mathcal{J}_{n}^{2}\lesssim A^{1/p}n^{-1/p}$.
Under Case 2, we have $\log N_{\triangle}(u^{2},\mathcal{R})\leq v\log(A/u^{2})$ and the integral $\int_{0}^{1}\sqrt{\log(A/u^{2})}\,du$ is finite (as $\int_{0}^{1}\sqrt{\log(1/u)}\,du<\infty$), so $\mathcal{J}_{n}\lesssim\sqrt{v/n}$ and hence $\mathcal{J}_{n}^{2}\lesssim v/n$.
Bounding $\rho$.
We use Lemma 29 to select an admissible modulus $\phi$ and then solve (or upper bound) the fixed point equation $\sqrt{n}\,\rho^{2}=\phi(\rho^{\alpha})$.
Case 1 with $0<p<1$. Lemma 29 and Lemma 26 imply that we may take $\phi(\sigma)\lesssim\int_{0}^{\sigma}\sqrt{\log N_{\triangle}(u^{2},\mathcal{R})}\,du$. Under Case 1, this gives
$\phi(\sigma)\lesssim\sqrt{A}\int_{0}^{\sigma}u^{-p}\,du\;\asymp\;\sqrt{A}\,\sigma^{1-p}.$
Thus the fixed point condition $\sqrt{n}\,\rho^{2}=\phi(\rho^{\alpha})$ yields
$\sqrt{n}\,\rho^{2}\;\lesssim\;\sqrt{A}\,(\rho^{\alpha})^{1-p}=\sqrt{A}\,\rho^{\alpha(1-p)},$
equivalently $\rho^{2-\alpha(1-p)}\lesssim\sqrt{A}\,n^{-1/2}$, hence
$\rho^{2}\lesssim\left(\frac{A}{n}\right)^{\frac{1}{2-\alpha(1-p)}}.$
Case 1 with $p=1$. When $p=1$ the local entropy integral diverges at 0; a convenient admissible choice is to upper bound local increments by global ones:
$\sup_{d(r,r^{\prime})\leq\sigma}|\nu_{n}(r)-\nu_{n}(r^{\prime})|\leq 2\sup_{r\in\mathcal{R}}|\nu_{n}(r)|.$
By symmetrization and a truncated entropy-integral bound for Rademacher averages (as in the definition Eq. 33), $\sqrt{n}\,\mathbb{E}\sup_{r}|\nu_{n}(r)|\lesssim\sqrt{A}\log n$ under $p=1$. Thus we may take $\phi(\sigma)\equiv C\sqrt{A}\log n$ (constant in $\sigma$), which is admissible. The fixed point equation then gives $\sqrt{n}\,\rho^{2}\lesssim\sqrt{A}\log n$, i.e.
$\rho^{2}\lesssim\frac{\sqrt{A}\,\log n}{\sqrt{n}}.$
Case 1 with $p>1$. Similarly, for $p>1$, the global (truncated) entropy integral implies $\sqrt{n}\,\mathbb{E}\sup_{r}|\nu_{n}(r)|\lesssim A^{1/(2p)}\,n^{1/2-1/(2p)}$. Thus we may take $\phi(\sigma)\equiv CA^{1/(2p)}\,n^{1/2-1/(2p)}$, yielding
$\rho^{2}\lesssim A^{1/(2p)}\,n^{-1/(2p)}.$
Case 2. Under Case 2 and Lemma 29, we may bound
$\phi(\sigma)\lesssim\int_{0}^{\sigma}\sqrt{v\log(A/u^{2})}\,du\;\lesssim\;\sigma\sqrt{v\log(A/\sigma^{2})}\qquad(\sigma\in(0,1]),$
using the standard estimate $\int_{0}^{\sigma}\sqrt{\log(1/u)}\,du\lesssim\sigma\sqrt{\log(1/\sigma)}$. Hence the fixed point condition implies
$\sqrt{n}\,\rho^{2}\lesssim\rho^{\alpha}\sqrt{v\log\!\left(\frac{A}{\rho^{2\alpha}}\right)}.$
Set $\beta=1/(2-\alpha)$ and define $\bar{\rho}^{2}=D\,(v\log(An)/n)^{\beta}$ with $D$ large enough. A direct substitution shows $\sqrt{n}\,\bar{\rho}^{2}\gtrsim\bar{\rho}^{\alpha}\sqrt{v\log(A/\bar{\rho}^{2\alpha})}$, since $\log(A/\bar{\rho}^{2\alpha})\lesssim\log(An)$ for $n\geq 2$ (and fixed $A,v,\alpha$). By monotonicity of $\sigma\mapsto\sqrt{n}\,\sigma^{2}-\phi(\sigma^{\alpha})$, the unique solution $\rho$ satisfies $\rho\leq\bar{\rho}$, that is,
$\rho^{2}\lesssim\left(\frac{v\log(An)}{n}\right)^{\frac{1}{2-\alpha}}.$
Assemble the bounds.
Plugging the bounds for $\mathcal{J}_{n}^{2}$ and $\rho^{2}$ into Eq. 34 yields the four displayed forms of $\mathfrak{R}_{n,\delta}$. ∎
E.5.7 A bipartite margin condition is sufficient for Assumption 13
We next show that the following margin condition is sufficient for Assumption 13 in the case of bipartite ranking. Recall that in the bipartite ranking setting, we define $\eta(X)=\mathbb{P}(Y=1\mid X)$.
Assumption 14 (Margin condition).
There exist constants $c_{\rm margin}>0$ and $\alpha\in[0,1]$ such that for all $x\in\mathcal{X}$ and $t>0$, we have $\mathbb{P}(\left\lvert\eta(X)-\eta(x)\right\rvert\leq t)\leq ct^{\alpha}$.
This is similar to the standard classification margin condition [Tsybakov, 2004, Tsybakov and Audibert, 2007], except it holds around every feasible threshold $\eta(x)$ of the conditional probability $\eta(X)$ instead of just around 0, which is the only threshold that matters for classification.
For a scoring function $s:\mathcal{X}\to\mathbb{R}$, denote by $L(s)$ the ranking risk induced by $r_{s}(x,x^{\prime})=2\mathbb{I}\left[s(x)\geq s(x^{\prime})\right]-1$ and by $L^{*}$ the Bayes ranking risk. As in example 1 of Clémençon et al. [2008], the excess ranking risk is
$\mathcal{E}(s)=L(s)-L^{*}=\mathbb{E}\Big[\,|\eta(X)-\eta(X^{\prime})|\,\mathbb{I}\left[(s(X)-s(X^{\prime}))(\eta(X)-\eta(X^{\prime}))<0\right]\Big].$
Let $h_{s}$ be the first-order term in the Hoeffding decomposition of the excess-risk U-statistic $\mathcal{E}_{n}(s)-\mathcal{E}(s)$ as in section 4 of Clémençon et al. [2008].
Lemma 30.
Under Assumption 14, for every scoring function $s$,
$\mathrm{Var}(h_{s}(X,Y))\;\leq\;(c_{\rm margin}+1)^{2}\,\mathcal{E}(s)^{\alpha}.$
In particular, Assumption 13 holds with the same exponent $\alpha$.
Proof.
Define the misranking event
$\mathcal{A}_{s}(x,x^{\prime})=\{(s(x)-s(x^{\prime}))(\eta(x)-\eta(x^{\prime}))<0\}.$
The first line in the proof of proposition 7 in Clémençon et al. [2008] yields the generic bound
$\mathrm{Var}(h_{s}(X,Y))\;\leq\;\mathbb{E}\Big[\big(\mathbb{E}_{X^{\prime}}\,\mathbb{I}\left[\mathcal{A}_{s}(X,X^{\prime})\right]\big)^{2}\Big].$ (35)
Fix $x\in\mathcal{X}$ and set
$g(x)=\mathbb{E}_{X^{\prime}}\,\mathbb{I}\left[\mathcal{A}_{s}(x,X^{\prime})\right],\qquad e(x)=\mathbb{E}_{X^{\prime}}\Big[|\eta(x)-\eta(X^{\prime})|\,\mathbb{I}\left[\mathcal{A}_{s}(x,X^{\prime})\right]\Big].$
Note that $0\leq e(x)\leq 1$ since $|\eta(\cdot)-\eta(\cdot)|\leq 1$. For any $t>0$, with $\Delta=|\eta(x)-\eta(X^{\prime})|$,
$g(x)\leq\mathbb{P}(\Delta\leq t)+\mathbb{E}_{X^{\prime}}\big[\mathbb{I}\left[\mathcal{A}_{s}(x,X^{\prime})\right]\mathbb{I}\left[\Delta>t\right]\big]\leq c_{\rm margin}t^{\alpha}+\frac{1}{t}\,e(x),$
because on $\{\Delta>t\}$ we have $\mathbb{I}\left[\mathcal{A}_{s}\right]\leq\Delta\,\mathbb{I}\left[\mathcal{A}_{s}\right]/t$. Choose $t=e(x)^{1/(1+\alpha)}$ (if $e(x)=0$, the bound is trivial). Then
$g(x)\leq(c_{\rm margin}+1)\,e(x)^{\alpha/(1+\alpha)}\quad\Longrightarrow\quad g(x)^{2}\leq(c_{\rm margin}+1)^{2}\,e(x)^{2\alpha/(1+\alpha)}.$
Since $\alpha\in[0,1]$, one has $2\alpha/(1+\alpha)\geq\alpha$, and since $e(x)\in[0,1]$, $e(x)^{2\alpha/(1+\alpha)}\leq e(x)^{\alpha}$. Therefore
$g(x)^{2}\leq(c_{\rm margin}+1)^{2}\,e(x)^{\alpha}.$
Taking expectation over $X$ and using Eq. 35 gives
$\mathrm{Var}(h_{s}(X,Y))\leq(c_{\rm margin}+1)^{2}\,\mathbb{E}\big[e(X)^{\alpha}\big].$
Because $u\mapsto u^{\alpha}$ is concave on $0,\infty)$ for $\alpha\in[0,1]$, Jensen’s inequality yields $\mathbb{E}[e(X)^{\alpha}]\leq(\mathbb{E}e(X))^{\alpha}$. Finally, by the definition of $\mathcal{E}(s)$, $\mathbb{E}e(X)=\mathcal{E}(s)$, so $\mathrm{Var}(h_{s}(X,Y))\leq(c_{\rm margin}+1)^{2}\,\mathcal{E}(s)^{\alpha}$. ∎
E.6 Proof of Theorem 8
Proof.
First note that the density conditions in Assumption 11, we satisfy Assumption 14 with $\alpha=1$. Hence, by Lemma 30, we have that Assumption 13 holds with $\alpha=1$. Next apply Theorem 11 to obtain a bound on $\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t_{\hat{\theta}_{\rm RSPO}})$ (there $L(r_{n})-L^{\star}$”). Note that the first term in the bound ($\inf_{r\in\mathcal{R}}\big(L(r)-L^{\star}\big)$”) is zero since $t^{\star}\in\mathcal{T}$ by Assumptions 1 and 2. Simplify the bound to involve only the leading terms by inflating the constant, noting that $\delta\leq 0.5$. Next, invoke Theorem 7 to translate the bound to a bound on $\inf_{a\geq 0}\|t^{\star}-at_{\hat{\theta}_{\rm RSPO}}\|^{2}$ for large enough $n$, noting that that the bound on AUC suboptimality is vanishing in $n$ so with sufficient probability and large enough $n$ it is less than $\iota$. Conclude by noting that $\inf_{a\geq 0}\|t^{\star}-at_{\hat{\theta}_{\rm RSPO}}\|$ latter dominates $\rho(h_{\hat{\theta}_{\rm RSPO}})$ under the coverage in Assumption 3. ∎
Appendix F Proof of Theorem 9
Proof.
We proceed to prove this for each $\theta$ satisfying the conditions with $\theta$-independent constant. Fix $\theta$. Write $h=h_{\theta}$, $\beta_{\star}=\beta_{\kappa,\theta}$, and suppress the $\theta$ dependence of $\hat{\Phi}$ and $\hat{\beta}$.
Without loss of generality assume also $f^{\prime}(1)=0$ (replace $f(t)$ by $f(t)-f^{\prime}(1)(t-1)$, which leaves $D_{f}$ unchanged). Define
$Z_{\beta}(x)=D_{f}\!\bigl(\pi_{\beta,h}(\cdot\mid x)\,\big\|\,\pi_{\mathrm{ref}}(\cdot\mid x)\bigr),\qquad\Phi(\beta)=\mathbb{E}_{x}Z_{\beta}(x).$
Set $g=(f^{\prime})^{-1}:\mathbb{R}\to(0,\infty)$, which under our assumptions is well-defined, $C^{1}$, and increasing. Define
$c_{-}=g\!\left(-\frac{2M}{\underline{\beta}}\right),\qquad c_{+}=g\!\left(\frac{2M}{\underline{\beta}}\right),$
and
$m_{f}=\inf_{t\in[c_{-},c_{+}]}f^{\prime\prime}(t),\qquad M_{f}=\sup_{t\in[c_{-},c_{+}]}f^{\prime\prime}(t),\qquad D_{\max}=\sup_{t\in[c_{-},c_{+}]}f(t).$
Under our assumptions, these are finite and satisfy $0<m_{f}\leq M_{f}<\infty$.
Existence/uniqueness of $\lambda_{\beta}(x)$ and ratio bounds. Fix $x$ and $\beta$. Let $u_{\beta}(y)=\beta^{-1}(h(x,y)-\lambda)$ and define
$F(\lambda)=\sum_{y\in\mathcal{Y}}\pi_{\mathrm{ref}}(y\mid x)\,g(u_{\beta}(y))-1.$
Since $g$ is continuous and increasing, $F$ is continuous and strictly decreasing in $\lambda$. Moreover, as $\lambda\to+\infty$, $u_{\beta}(y)\to$-$\infty$ so $g(u_{\beta}(y))\to 0$ and $F(\lambda)\to-1$; as $\lambda\to$-$\infty$, $u_{\beta}(y)\to+\infty$ so $g(u_{\beta}(y))\to+\infty$ and $F(\lambda)\to+\infty$. Thus there exists a unique $\lambda_{\beta}(x)$ with $F(\lambda_{\beta}(x))=0$.
Let $h_{\max}(x)=\max_{y}h(x,y)$ and $h_{\min}(x)=\min_{y}h(x,y)$. Because $g(0)=1$ (since $f^{\prime}(1)=0$), if $\lambda=h_{\max}(x)$ then $u_{\beta}(y)\leq 0$ for all $y$, hence $g(u_{\beta}(y))\leq 1$ and $\sum_{y}\pi_{\mathrm{ref}}g(u_{\beta}(y))\leq 1$; similarly if $\lambda=h_{\min}(x)$ then $\sum_{y}\pi_{\mathrm{ref}}g(u_{\beta}(y))\geq 1$. By monotonicity of $F$, we obtain
$h_{\min}(x)\ \leq\ \lambda_{\beta}(x)\ \leq\ h_{\max}(x).$
Since $\|h\|_{\infty}\leq M$, this implies $|\lambda_{\beta}(x)|\leq M$ and therefore
$|h(x,y)-\lambda_{\beta}(x)|\leq 2M\quad\Rightarrow\quad u_{\beta}(x,y)\in\Bigl[-\frac{2M}{\beta},\frac{2M}{\beta}\Bigr]\subseteq\Bigl[-\frac{2M}{\underline{\beta}},\frac{2M}{\underline{\beta}}\Bigr]\quad\text{for all }\beta\in\mathcal{B}.$
Hence the ratio
$r_{\beta}(x,y)=\frac{\pi_{\beta,h}(y\mid x)}{\pi_{\mathrm{ref}}(y\mid x)}=g(u_{\beta}(x,y))$
lies in $c_{-},c_{+}]$ for all $(x,y)$ and all $\beta\in\mathcal{B}$.
Boundedness of $Z_{\beta}(x)$. Since $r_{\beta}(x,y)\in[c_{-},c_{+}]$ and $Z_{\beta}(x)=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)f(r_{\beta}(x,y))$ with $\sum_{y}\pi_{\mathrm{ref}}(y\mid x)=1$,
$0\leq Z_{\beta}(x)\leq D_{\max}\qquad\text{for all $x$ and $\beta\in\mathcal{B}$}.$
Differentiability and bounds for $\Phi^{\prime}(\beta)$ and Lipschitzness of $Z_{\beta}(\cdot)$. Fix $x$ and abbreviate $\lambda_{\beta}=\lambda_{\beta}(x)$, $u(y)=\beta^{-1}(h(x,y)-\lambda_{\beta})$, $r(y)=g(u(y))$, and
$D_{x}(\beta)=Z_{\beta}(x)=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,f(r(y)).$
The constraint $\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,r(y)=1$ can be written as
$G(\beta,\lambda)=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,g\!\left(\beta^{-1}(h(x,y)-\lambda)\right)-1=0.$
Since $g$ is $C^{1}$ and $g^{\prime}>0$, we have
$\partial_{\lambda}G(\beta,\lambda)=-\frac{1}{\beta}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,g^{\prime}(u(y))<0,$
so by the implicit function theorem $\beta\mapsto\lambda_{\beta}$ is $C^{1}$ on $(0,\infty)$.
Differentiate $D_{x}(\beta)$ using $f^{\prime}(r(y))=u(y)$ and $r^{\prime}(y)=g^{\prime}(u(y))u^{\prime}(y)$:
$D_{x}^{\prime}(\beta)=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,f^{\prime}(r(y))\,r^{\prime}(y)=\sum_{y}w(y)\,u(y)\,u^{\prime}(y),\qquad w(y)=\pi_{\mathrm{ref}}(y\mid x)\,g^{\prime}(u(y)).$
Differentiating $u(y)=\beta^{-1}(h(x,y)-\lambda_{\beta})$ gives $u^{\prime}(y)=-\beta^{-2}(h(x,y)-\lambda_{\beta})-\beta^{-1}\lambda_{\beta}^{\prime}$. Differentiating the normalization $\sum_{y}\pi_{\mathrm{ref}}r(y)=1$ yields $\sum_{y}w(y)u^{\prime}(y)=0$, which implies
$\lambda_{\beta}^{\prime}=-\frac{1}{\beta}\frac{\sum_{y}w(y)\,(h(x,y)-\lambda_{\beta})}{\sum_{y}w(y)}.$
Substituting this identity and simplifying (a standard variance algebra) gives
$D_{x}^{\prime}(\beta)=-\frac{A(\beta)}{\beta^{3}}\,\mathrm{Var}_{p_{\beta}}\bigl(h(x,\cdot)\bigr),\qquad A(\beta)=\sum_{y}w(y),\quad p_{\beta}(y)=\frac{w(y)}{A(\beta)}.$ (36)
On $c_{-},c_{+}]$ we have $m_{f}\leq f^{\prime\prime}\leq M_{f}$, hence
$g^{\prime}(u(y))=\frac{1}{f^{\prime\prime}(r(y))}\in\Bigl[\frac{1}{M_{f}},\frac{1}{m_{f}}\Bigr],$
so $A(\beta)\in[1/M_{f},\,1/m_{f}]$ and thus from Eq. 36 and $|h|\leq M$,
$|D_{x}^{\prime}(\beta)|\leq\frac{1}{m_{f}}\frac{1}{\underline{\beta}^{\,3}}\cdot\mathrm{Var}_{p_{\beta}}(h(x,\cdot))\leq\frac{1}{m_{f}}\frac{1}{\underline{\beta}^{\,3}}\cdot\mathbb{E}_{p_{\beta}}[h(x,\cdot)^{2}]\leq\frac{M^{2}}{m_{f}\,\underline{\beta}^{\,3}}.$
Therefore $\beta\mapsto Z_{\beta}(x)$ is Lipschitz on $\mathcal{B}$ with constant
$L_{Z}=\frac{M^{2}}{m_{f}\,\underline{\beta}^{\,3}}.$
Averaging Eq. 36 over $x$ yields $\Phi^{\prime}(\beta)=\mathbb{E}_{x}D_{x}^{\prime}(\beta)$.
To lower bound $\Phi^{\prime}$, note that $p_{\beta}(y)=\pi_{\mathrm{ref}}(y\mid x)\,g^{\prime}(u(y))/A(\beta)$ and $g^{\prime}(u(y))\geq 1/M_{f}$ while $A(\beta)\leq 1/m_{f}$, hence
$p_{\beta}(y)\ \geq\ \frac{m_{f}}{M_{f}}\,\pi_{\mathrm{ref}}(y\mid x)\qquad\text{for all $y$}.$
Thus for any $c\in\mathbb{R}$, $\sum_{y}p_{\beta}(y)(h(x,y)-c)^{2}\geq\frac{m_{f}}{M_{f}}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)(h(x,y)-c)^{2}$, and minimizing over $c$ gives
$\mathrm{Var}_{p_{\beta}}\bigl(h(x,\cdot)\bigr)\geq\frac{m_{f}}{M_{f}}\,\mathrm{Var}_{\pi_{\mathrm{ref}}(\cdot\mid x)}\bigl(h(x,\cdot)\bigr).$
Using $A(\beta)\geq 1/M_{f}$ in Eq. 36, for all $\beta\in\mathcal{B}$,
$\Phi^{\prime}(\beta)=\mathbb{E}_{x}\frac{A(\beta)}{\beta^{3}}\mathrm{Var}_{p_{\beta}}(h(x,\cdot))\geq\frac{1}{M_{f}}\frac{1}{\overline{\beta}^{\,3}}\cdot\frac{m_{f}}{M_{f}}\,\mathbb{E}_{x}\mathrm{Var}_{\pi_{\mathrm{ref}}(\cdot\mid x)}(h(x,\cdot))=\ c_{\Phi},$
where
$c_{\Phi}=\frac{m_{f}}{M_{f}^{2}\,\overline{\beta}^{\,3}}\,v\ >\ 0.$
Existence and uniqueness of $\beta_{\star}$. By Step 3, $\Phi$ is $C^{1}$ on $(0,\infty)$ with $\Phi^{\prime}(\beta)\leq-c_{\Phi}<0$ on $\mathcal{B}$, hence strictly decreasing. As $\beta\to\infty$, Step 1 implies $u_{\beta}(x,y)\to 0$ uniformly and thus $r_{\beta}(x,y)\to g(0)=1$, whence $Z_{\beta}(x)\to f(1)=0$ and $\Phi(\beta)\to 0$ by dominated convergence. As $\beta\downarrow 0$, a standard vanishing-regularization argument shows $\Phi(\beta)\to\Psi$ (the unique minimum-divergence greedy limit), hence $\lim_{\beta\downarrow 0}\Phi(\beta)=\Psi>\kappa$. Therefore, by continuity and strict monotonicity, there is a unique $\beta_{\star}\in(0,\infty)$ with $\Phi(\beta_{\star})=\kappa$. (And by construction $\beta_{\star}\in(\underline{\beta},\overline{\beta})$.)
Lipschitzness of the policy map in $\beta$. Fix $x$ and $y$ and write $\pi_{\beta}=\pi_{\beta,h}$. Since $\pi_{\beta}(y\mid x)=\pi_{\mathrm{ref}}(y\mid x)\,r_{\beta}(x,y)$ with $r_{\beta}=g(u_{\beta})$, and $g^{\prime}(u)\leq 1/m_{f}$ on $\mathcal{B}$, it suffices to bound $\partial_{\beta}u_{\beta}$. From $u_{\beta}(x,y)=\beta^{-1}(h(x,y)-\lambda_{\beta}(x))$ and $|h-\lambda_{\beta}|\leq 2M$ (Step 1),
$\left|\partial_{\beta}u_{\beta}(x,y)\right|\leq\frac{|h(x,y)-\lambda_{\beta}(x)|}{\beta^{2}}+\frac{|\lambda_{\beta}^{\prime}(x)|}{\beta}\leq\frac{2M}{\underline{\beta}^{\,2}}+\frac{|\lambda_{\beta}^{\prime}(x)|}{\underline{\beta}}.$
Using the explicit formula for $\lambda_{\beta}^{\prime}$ above and $|h-\lambda_{\beta}|\leq 2M$,
$|\lambda_{\beta}^{\prime}(x)|=\frac{1}{\beta}\left|\frac{\sum_{y}w(y)(h(x,y)-\lambda_{\beta}(x))}{\sum_{y}w(y)}\right|\leq\frac{2M}{\underline{\beta}}.$
Thus $\sup_{\beta\in\mathcal{B},x,y}|\partial_{\beta}u_{\beta}(x,y)|\leq 4M/\underline{\beta}^{\,2}$ and hence
$\sup_{\beta\in\mathcal{B},x,y}|\partial_{\beta}r_{\beta}(x,y)|\leq\frac{1}{m_{f}}\cdot\frac{4M}{\underline{\beta}^{\,2}}=\frac{4M}{m_{f}\,\underline{\beta}^{\,2}}.$
Therefore for any $\beta,\beta^{\prime}\in\mathcal{B}$ and any $x$,
$\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,|\pi_{\beta}(y\mid x)-\pi_{\beta^{\prime}}(y\mid x)|=\sum_{y}\pi_{\mathrm{ref}}(y\mid x)^{2}\,|r_{\beta}-r_{\beta^{\prime}}|\leq\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,|r_{\beta}-r_{\beta^{\prime}}|\leq\frac{4M}{m_{f}\,\underline{\beta}^{\,2}}|\beta-\beta^{\prime}|.$
Averaging over $x$ gives
$\mathbb{E}_{x}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,|\pi_{\beta,h}(y\mid x)-\pi_{\beta^{\prime},h}(y\mid x)|\leq L_{\pi}|\beta-\beta^{\prime}|,\qquad L_{\pi}=\frac{4M}{m_{f}\,\underline{\beta}^{\,2}}.$ (37)
Uniform concentration of $\hat{\Phi}_{m}$ on $\mathcal{B}$. Let $N=m$ and $\Delta=(\overline{\beta}-\underline{\beta})/N$, and consider the grid $\beta_{k}=\underline{\beta}+k\Delta$ for $k=0,\dots,N$. Since $Z_{\beta}(x)$ is $L_{Z}$-Lipschitz in $\beta$ (Step 3), the standard discretization argument gives
$\sup_{\beta\in\mathcal{B}}|\hat{\Phi}_{m}(\beta)-\Phi(\beta)|\leq\max_{0\leq k\leq N}|\hat{\Phi}_{m}(\beta_{k})-\Phi(\beta_{k})|+2L_{Z}\Delta.$
By Step 2, $0\leq Z_{\beta_{k}}(x)\leq D_{\max}$, so Hoeffding’s inequality and a union bound yield that with probability at least $1-\delta$,
$\max_{0\leq k\leq N}|\hat{\Phi}_{m}(\beta_{k})-\Phi(\beta_{k})|\leq D_{\max}\sqrt{\frac{\log(2(N+1)/\delta)}{2m}}=D_{\max}\sqrt{\frac{\log(2(m+1)/\delta)}{2m}}.$
Hence on this event,
$\varepsilon=\sup_{\beta\in\mathcal{B}}|\hat{\Phi}_{m}(\beta)-\Phi(\beta)|\leq D_{\max}\sqrt{\frac{\log(2(m+1)/\delta)}{2m}}+2L_{Z}\frac{\overline{\beta}-\underline{\beta}}{m}.$
Convert uniform calibration error to $\beta$-error to policy error. Because $\hat{\beta}$ minimizes $|\hat{\Phi}_{m}(\beta)-\kappa|$ over $\mathcal{B}$ and $\Phi(\beta_{\star})=\kappa$,
$|\hat{\Phi}_{m}(\hat{\beta})-\kappa|\leq|\hat{\Phi}_{m}(\beta_{\star})-\kappa|=|\hat{\Phi}_{m}(\beta_{\star})-\Phi(\beta_{\star})|\leq\varepsilon.$
Thus $|\Phi(\hat{\beta})-\kappa|\leq 2\varepsilon$, hence $|\Phi(\hat{\beta})-\Phi(\beta_{\star})|\leq 2\varepsilon$. By the mean value theorem and $\Phi^{\prime}(\beta)\geq c_{\Phi}$ on $\mathcal{B}$ (Step 3),
$|\hat{\beta}-\beta_{\star}|\leq\frac{2\varepsilon}{c_{\Phi}}.$
Finally, applying Eq. 37 with $\beta=\hat{\beta}$ and $\beta^{\prime}=\beta_{\star}$ yields
$\mathbb{E}_{x}\sum_{y}\pi_{\mathrm{ref}}(y\mid x)\,|\pi_{\hat{\beta},h}(y\mid x)-\pi_{\beta_{\star},h}(y\mid x)|\leq L_{\pi}|\hat{\beta}-\beta_{\star}|\leq\frac{2L_{\pi}}{c_{\Phi}}\,\varepsilon.$
Substituting $L_{\pi}=\frac{4M}{m_{f}\underline{\beta}^{\,2}}$, $c_{\Phi}=\frac{m_{f}}{M_{f}^{2}\overline{\beta}^{\,3}}v$, and $L_{Z}=\frac{M^{2}}{m_{f}\underline{\beta}^{\,3}}$ gives the bound
$\mathbb{E}_{x}\sum_{y\in\mathcal{Y}}\pi_{\mathrm{ref}}(y\mid x)\,\bigl|\pi_{\kappa,h}(y\mid x)-\pi_{\hat{\beta},h}(y\mid x)\bigr|\ \leq\ \frac{8MM_{f}^{2}\,\overline{\beta}^{\,3}}{m_{f}^{\,2}\,\underline{\beta}^{\,2}\,v}\left(D_{\max}\sqrt{\frac{\log(2(m+1)/\delta)}{2m}}\;+\;\frac{2M^{2}(\overline{\beta}-\underline{\beta})}{m_{f}\,\underline{\beta}^{\,3}\,m}\right).$
Noting that $\delta\leq 0.5$ means the statement is true for some $c>0$. ∎

### v2

@@
 Abstract
-Aligning large language models to preference data is commonly implemented by assuming a known link function between the distribution of observed preferences and the unobserved rewards (e.g., a logistic link as in Bradley-Terry).
-If the link is wrong, however, inferred rewards can be biased and policies be misaligned.
-We study policy alignment to preferences under an unknown and unrestricted link.
-We consider an $f$-divergence-constrained reward maximization problem and show that realizability of the solution in a policy class implies a semiparametric single-index binary choice model, where a scalar-valued index determined by a policy captures the dependence on demonstrations and the rest of the preference distribution is an unrestricted function thereof.
-Rather than focus on estimation of identifiable finite-dimensional structural parameters in the index as in econometrics, we focus on policy learning, focusing on error to the optimal policy and allowing unidentifiable and nonparametric indices.
-We develop a variety of policy learners based on profiling the link function, orthogonalizing the link function, and using link-agnostic bipartite ranking objectives.
-We analyze these and provide finite-sample policy error bounds that depend on generic functional complexity measures of the index class.
-We further consider practical implementations using first-order optimization suited to neural networks and batched data.
-The resulting methods are robust to unknown preference noise distribution and scale, while preserving the direct optimization of policies without explicitly fitting rewards.
+Aligning large language models (LLMs) to preference data typically assumes a known link function between observed preferences and latent rewards (e.g., a logistic Bradley-Terry link).
+Misspecification of this link can bias inferred rewards and misalign learned policies.
+We study preference alignment under an unknown and unrestricted link function.
+We show that realizability of $f$-divergence-constrained reward maximization in a policy class induces a semiparametric single-index binary choice model, where a scalar policy-dependent index captures all dependence on demonstrations and the remaining preference distribution is unrestricted.
+Rather than assuming this model has identifiable finite-dimensional structural parameters and estimating them, as in econometrics, we focus on policy learning with the reward function implicit, analyzing error to the optimal policy and allowing for unidentifiable nonparametric indices.
+We develop preference optimization algorithms robust to the unknown link and prove convergence guarantees in terms of generic function complexity measures.
+We demonstrate this empirically on LLM alignment.
+Code is available at https://github.com/causalml/spo/.
 1 Introduction
@@
 The policy class induces a class of indices (corresponding to reward or utility functions), which can be an arbitrary nonparametric class, and moreover the true index (corresponding to the true reward function) is generally not identifiable.
-We study policy learning methods and analyze bounds on the policy difference to the optimal policy (rather than convergence of structural estimates).
-Finally, we give practical gradient-based optimization algorithms that can be used to align neural-net-based models such as transformer LLMs.
-We call the resulting general approach Semiparametric Preference Optimization (SPO).
-The primary algorithm we study is Orthogonalized SPO (OSPO), which performs empirical risk minimization with a loss function that depends on a family of univariate regression functions related to the unknown link.
-The analysis shows we can simply plug in estimates for these regression (for example, kernel regression estimates, since the regression is just univariate) and the method is locally invariant to the estimation: the error bounds scale quadratically in the error of the plug-in.
-This shows how we can perform almost as well as if we had known the link exactly (as, for example, DPO assumes we do) while being robust to not knowing the link at all (e.g., even when it is not logistic).
-In summary, SPO keeps the divergence-constrained policy search without an explicit reward model of DPO, while removing unverifiable assumptions on the reward-to-preference link and still provably optimally aligning policies to preferences.
+We study policy learning methods and analyze bounds on the policy difference to the optimal policy.
+We call the resulting approach Semiparametric Preference Optimization (SPO).
 Summary of contributions
-  • From RLHF to single-index model.
-We formulate realizability of the $f$-divergence-constrained optimal policy and preferences observed with unknown link function as a single-index model of binary choice, with an index given by our policy class.
-  • Semiparametric policy learners for RLHF.
-We develop the policy learners Profiled SPO (PSPO), Orthogonalized SPO (OSPO), Ranking SPO (RSPO) as three ways to search over policies that would maximize the likelihood or rank of their implied index in this model, without explicitly learning a reward model.
+  • From RLHF to a single-index model.
+We show that realizability of the $f$-divergence-constrained optimal policy with unknown link induces a semiparametric single-index choice model, with an index determined by the policy class.
+We show the optimal policy is generally learnable even thought the index is not.
+  • Semiparametric policy learners.
+We develop the policy learners Profiled SPO (PSPO), Orthogonalized SPO (OSPO), and Ranking SPO (RSPO) as three ways to optimize over policies without specifying the link or explicitly fitting a reward model.
+PSPO maximizes the likelihood also over the unknown link and provides a proof of concept that specifying the link can be avoided for policy learning.
+OSPO leverages a plug-in estimate for a nuisance regression related to the link function and to first order behaves as though the link function were known.
+RSPO avoids nuisance estimation and, while it can be less statistically efficient, it can be instantiated as a simple and practical “pairs-of-pairs” variant of DPO (PoP-DPO) that can provide additional robustness to the link function.
   • Policy-learning theory.
-We formalize an affine-invariant metric on indices and prove it controls policy difference to the optimal policy.
-We analyze guarantees of our policy learners in terms of this metric, showing consistency of PSPO (without a rate) and establishing rates for OSPO and RSPO that naturally depend on the functional complexity of the given classes.
-  • Practical algorithms.
-We provide practical first-order algorithms suited to neural networks and batched data and study a simple empirical recalibration step at the end that enforces the divergence budget.
+We establish policy convergence guarantees for our methods under functional complexity conditions that admit generic policy classes (and implied reward index classes).
+These leverage analyses of the statistics of each empirical optimization problem, how near optimality translates to index error up to its non-identifiable aspects, how that in turn translates to policy error, and how plug-in nuisance error can have vanishing impact on performance.
+  • Algorithms and experiments.
+We provide practical first-order algorithms suited to neural networks and batched data and an empirical recalibration step to enforce the divergence budget.
+We demonstrate these methods numerically first in an illustrative synthetic example and then in LLM alignment using the Qwen3 model and UltraFeedback dataset.
 Roadmap
@@
 Section 6 analyzes empirical calibration to the divergence budget.
+Section 7 provides an empirical investigation of the algorithms, including in LLM alignment.
 Section 8 discusses possible extensions.
@@
 $z\mid x,y_{0},y_{1}$ $\sim\mathrm{Bernoulli}\left(\Psi\left(t_{\theta}(x,y_{0},y_{1})\right)\right),$ (4)
-$h_{\theta}(x,y)$ $=f^{\prime}\left(\frac{\pi_{\theta}(y\mid x)}{\pi_{\rm ref}(y\mid x)}\right),\qquad t_{\theta}(x,y_{0},y_{1})=h_{\theta}(x,y_{1})-h_{\theta}(x,y_{0}).$ (5)
+$h_{\theta}(x,y)$ $=f^{\prime}\left(\frac{\pi_{\theta}(y\mid x)}{\pi_{\rm ref}(y\mid x)}\right),$ (5)
+$t_{\theta}(x,y_{0},y_{1})$ $=h_{\theta}(x,y_{1})-h_{\theta}(x,y_{0}).$ (6)
 Combining Theorem 1 with Eq. 3 and rearranging establishes that this is a valid model under policy realizability.
@@
 may be parametric or nonparametric and is in general as rich as the policy class $\{\pi_{\theta}:\theta\in\Theta\}$.
-As such, we will characterize $\mathcal{T}$ in terms of its own functional complexity, permitting rich and growing function classes.
+As such, we will characterize $\mathcal{T}$ in terms of its own functional complexity, permitting rich classes and treating them generically.
 2.4 From a single-index model to a policy
@@
 Here the likelihood is Eq.
-6, which profiles the link rather than assuming it is known and plugging it in.
+7, which profiles the link rather than assuming it is known and plugging it in.
 3.2 Policy consistency of PSPO
@@
 7 in practice, especially when $\pi_{\theta}$ is a neural network, we can use a first-order (gradient-based) method.
-A simple approach would be to alternate taking a gradient step in $\theta$, taking $\Psi$ as fixed, with optimizing $\Psi$ for the current $\theta$.
+A simple approach would be to alternate between taking a (batched stochastic) gradient step in $\theta$, taking $\Psi$ as fixed, and optimizing $\Psi$ for the current $\theta$.
 The former can be combined with any momentum and/or adaptive learning rate procedure.
@@
 Unfortunately, we do not know $g^{\star}_{\theta}$ in Eq.
-7.
+8.
 Fortunately, we can easily estimate $g_{\theta}$ as a univariate regression of $z$ on $t_{\theta}$.
@@
 The dataset $\mathcal{D}$ can be the same $n$ datapoints used in Eq.
-8.
+9.
 Klein and Spady (1993) use a leave-one-out version where $\mathcal{D}=\{(w_{j},z_{j}):j\neq i\}$ is used to train the kernel regression plugged into the $i^{\text{th}}$ term in Eq.
-8.
+9.
 That is the practical version we recommend.
@@
 This is in contrast to PSPO’s objective Eq.
-6, which restricts the profiling to monotonic nondecreasing link functions so that negating the index would not preserve the profile likelihood objective.
+7, which restricts the profiling to monotonic nondecreasing link functions so that negating the index would not preserve the profile likelihood objective.
 Focusing on identifying policies rather than underlying parameters, we will instead impose a slight strengthening of our up-to-scaling identification condition (Assumption 4) so that near-observational-equivalence translates to nearly pointing in the same (or opposite) directions.
@@
 The same holds for the second Lipschitz equation if $\hat{g}$ is similarly scale invariant (such as is the case for the kernel regression in Eq.
-9).
+10).
 We also need to control the complexity of the index class $\mathcal{T}$.
@@
 4.7 First-order optimization
+Given a differentiable $\hat{g}_{\theta}$ (e.g., kernel regression), we can directly apply standard first-order optimizers to Eq. 9.
 Theoretically, to obtain the aligned policy, we must split the data four ways: one quarter to learn $\hat{g}_{\theta}$ for all $\theta\in\Theta$, one quarter to construct the quasi-likelihood and fit $\hat{\theta}^{\rm OSPO}$ to maximize it, one quarter to learn the sign of the index using $\hat{s}$, and finally one quarter to learn $\hat{\beta}_{\hat{\theta}^{\rm OSPO}}$ to calibrate to the divergence constraint (see Section 6).
@@
 5.1 From AUC maximization to an index and a policy
-Define the AUC as
+Define the AUC of an index as
 $\mathrm{AUC}(t)=$ $~\mathbb{P}_{(w_{0},z_{0}),(w_{1},z_{1})\sim P^{2}}(t(w_{1})>t(w_{0})\mid z_{0}=0,z_{1}=1)$ (11)
@@
 $\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t_{\theta})\geq\lambda_{0}\inf_{a\geq 0}\|t^{\star}-at_{\theta}\|^{2}.$
+Note that unlike Lemma 1, here we can restrict to nonnegative scalings.
+Indeed, AUC is not invariant to flipping the sign – that will yield the complement of the AUC.
+Our regularity conditions above ensure that the optimal AUC (attained by $t^{\star}$) is strictly above $0.5$ and that the AUC is $\frac{2}{3}$-Hölder continuous in the score function.
+Since AUC is invariant to positive scalings, this means that for $\iota$ small enough, the covariance of $t^{\star}$ and $t_{\theta}$ cannot be negative.
 5.2 The RSPO policy learner
@@
 13 and Eq.
-13, respectively) are generally distinct.
+14, respectively) are generally distinct.
 In practice, since the symmetric assumption is quite reasonable, we suggest to use Eq.
@@
 In theory, in order to easily leverage existing results (Clémençon et al., 2008) and since the rates would be unaffected, we analyze Eq.
-12.
+13.
 Since the AUC objective involves a step function, to appropriately control the complexity we need to consider the sets of pairs that ranked in order or anti-order.
@@
 Let us focus on transforming Eq.
-13, as that is the variant we recommend in practice; the transformation is analogous for Eq.
-12.
-Create the $2n^{2}$ data points, where for each $i,j$, we label the features $(x_{j},y_{j,1-z_{j}},y_{j,z_{j}},x_{i},y_{i,z_{i}},y_{i,1-z_{i}})$ with $1$ and the features $(x_{j},y_{j,z_{j}},y_{j,1-z_{j}},x_{i},y_{i,1-z_{i}},y_{i,z_{i}})$ with 0.
+14, as that is the variant we recommend in practice; the transformation is analogous for Eq.
+13.
+Create the $2n^{2}$ data points, where for each $i,j$, we label the features $(x_{j},y_{j,1-z_{j}},y_{j,z_{j}},x_{i},y_{i,z_{i}},y_{i,1-z_{i}})$ with 11 and the features $(x_{j},y_{j,z_{j}},y_{j,1-z_{j}},x_{i},y_{i,1-z_{i}},y_{i,z_{i}})$ with 0.
 Then Eq.
@@
 Now use any convex classification surrogate loss, such as hinge or logistic.
-Now, rather than taking batches over the $n^{2}$ augmented data points, we can take batches of size $B$ from the original $n$ data points and simply consider the gradient for the corresponding $B^{2}$ terms.
+To use batched first-order optimization, rather than taking batches over the $n^{2}$ augmented data points, we can take batches of size $B$ from the original $n$ data points and simply consider the gradient for the corresponding $B^{2}$ terms, as such would still form a stochastic gradient.
+Remark 4 (Pairs-of-Pairs DPO).
+Using the symmetric-link AUC objective in Eq.
+14, replacing $\phi$ with a logistic surrogate loss ($\log(1+e^{-u})$), and instantiating with KL divergence ($f(u)=u\log u$) yields a pairs-of-pairs DPO (PoP-DPO) variant:
+$\hat{\theta}_{\rm PoP-DPO}\in\operatornamewithlimits{argmin}_{\theta\in\Theta}\frac{1}{n^{2}}\sum_{i,j}\log\bigl(1+\frac{\pi_{\theta}(y_{j,1-z_{j}}\mid x_{j})}{\pi_{\rm ref}(y_{j,1-z_{j}}\mid x_{j})}\frac{\pi_{\theta}(y_{i,1-z_{i}}\mid x_{i})}{\pi_{\rm ref}(y_{i,1-z_{i}}\mid x_{i})}\frac{\pi_{\rm ref}(y_{j,z_{j}}\mid x_{j})}{\pi_{\theta}(y_{j,z_{j}}\mid x_{j})}\frac{\pi_{\rm ref}(y_{j,z_{j}}\mid x_{j})}{\pi_{\theta}(y_{j,z_{j}}\mid x_{j})}\bigr).$
+This is similar to DPO’s objective, minimizing $\frac{1}{n}\sum_{i}\log\left(1+\frac{\pi_{\theta}(y_{i,1-z_{i}}\mid x_{i})}{\pi_{\rm ref}(y_{i,1-z_{i}}\mid x_{i})}\frac{\pi_{\rm ref}(y_{i,z_{i}}\mid x_{i})}{\pi_{\theta}(y_{i,z_{i}}\mid x_{i})}\right)$.
+Seen as a surrogate-loss classification algorithm, DPO seeks consistent ranking within each demonstration pair, while PoP-DPO compares pairs of pairs and is able to explore a greater range of contrasts rather than rely on the link to extrapolate from just within-pair comparisons, offering some additional robustness, since we are not interested in simply the classification quality but in the fidelity of the learned reward/index.
 6 Empirical Recalibration to the Divergence Budget
@@
 In practice, we can use the same data since calibration should be rather stable and only ever uses the $x$-data.
+Remark 5 (Sampling from $\pi_{\beta,\theta}(y\mid x)$ with autoregressive LLMs).
+In the case of KL divergence ($f(u)=u\log u$), we have $\pi_{\beta,\theta}(y\mid x)\propto(\pi_{\rm ref}(y\mid x))^{1-1/\beta}(\pi_{\theta}(y\mid x))^{1/\beta}$.
+If $y=(y_{1},y_{2},\dots)$ consists of a sequence of tokens and $\pi_{\rm ref}(y\mid x)$, $\pi_{\theta}(y\mid x)$ are given in the form of token-by-token conditionals $\pi_{\rm ref}(y_{t}\mid x,y_{<t})$, $\pi_{\theta}(y_{t}\mid x,y_{<t})$, which are in turn given by a softmax distribution over some (not necessarily normalized) logits $g_{\rm ref}(y_{t}\mid x,y_{<t})$, $g_{\theta}(y_{t}\mid x,y_{<t})$, then to sample from $\pi_{\beta,\theta}(y\mid x)$ we can simply sample token-by-token from the softmax distribution over the logits $(1-1/\beta)g_{\rm ref}(y_{t}\mid x,y_{<t})+(1/\beta)g_{\theta}(y_{t}\mid x,y_{<t})$.
+7 Empirical investigation
+[…]
+7.1 Synthetic preference optimization experiment
+First, we consider an illustrative small synthetic setting.
+We simulate preference data with 20-dimensional Gaussian contexts $x$ and $\left\lvert\mathcal{Y}\right\rvert=10$ actions.
+We set the reference policy to uniform and consider a policy class $\pi_{\theta}$ given by a feedforward network with two hidden layers.
+We fix some $\theta^{\star}$ and set $r^{\star}(x,y)=10\log(\pi_{\theta^{\star}}(y\mid x)/\pi_{\mathrm{ref}}(y\mid x))$.
+Preferences are generated from an even mixture of two shifted logistic links $0.5\,\sigma(4(u-s))+0.5\,\sigma(4(u+s))$ with shift $s$ controlling misspecification ($s=0$ corresponds to logistic link).
+Each method learns $\pi_{\theta}$ from $n=1000$ preference pairs, and we repeat the experiment 1000 times.
+Additional detail is given in Section G.1.
+(a) Reward at a given KL budget.
+(b) Achievable reward-KL, $s=0$.
+(c) Achievable reward-KL, $s=1.5$.
+Figure 1: Synthetic preference optimization experiment.
+Figure 1 shows the average reward (and 90% confidence intervals) at KL budget $\kappa=0.2$ as $s\in[0,1.5]$ varies and the average reward-divergence curves for $\pi_{\beta,\hat{\theta}}$ as $\beta$ varies, at $s=0$ and $s=1.5$.
+All methods perform well when logistic is well-specified ($s=0$).
+DPO performance deteriorates as $s$ grows, and so does RSPO due to its logistic surrogate loss, albeit providing some robustness at low misspecification.
+PSPO performs less well, consistent with its rough loss surface, but is robust to misspecified link, with performance roughly flat as $s$ varies.
+There is some light deterioration due to the lower SNR given the link is very flat around 0 when $s$ is large (rather than due to misspecification).
+OSPO consistently performs the best across all $s$, consistent with it being efficient no matter what the link is, whether logistic or not.
+7.2 Aligning Qwen3 on UltraFeedback
+We next consider an experiment aligning the Qwen3 model (Yang et al., 2025) using the UltraFeedback data set (Cui et al., 2024) and Skywork-Reward-V2 reward model (Liu et al., 2025).
+We take as our reference model Qwen3-0.6B after some light supervised fine-tuning on UltraFeedback responses to ensure format adherence.
+In order to be able to evaluate and compare learned policies, we take Skywork-Reward-V2-Qwen3-1.7B as the ground truth reward function.
+For prompts from UltraFeedback, we generate two responses from the reference policy, score them with the reward model, and take the difference, normalizing so that differences have variance 1 across the data.
+We generate preference observations using the same mixture of shifted logitic links as in the last section.
+We run each method on the preference observations alone (no reward) to train a rank-16 LoRA.
+We then vary $\beta$, use Remark 5 to generate responses from the aligned and scaled policy, and score using the reward model.
+Additional detail is given in Section G.2.
+Figure 2 depicts the resulting average reward over 4 replications at $\kappa=0.5$ (with 90% confidence intervals for the difference to DPO).
+It overall appears to convey a similar story as the synthetic experiment, albeit with more noise, given only four replications.
+We again see all methods perform similarly with a well-specified logistic link, DPO deteriorating when the logistic link becomes more misspecified, and the SPO methods remaining robust to a non-logistic link.
+OSPO and PSPO seem to be noisier and possibly unstable, consistent with their more intricate training procedures.
+RSPO (as PoP-DPO) seems to strike a good balance of simplicity and robustness, offering a practical approach to still use logistic loss optimization while squeezing more comparisons from preference data rather than rely on a well-specified link to extrapolate, given we are interested in more than just classification.
+Figure 2: Aligning Qwen3 on UltraFeedback.
 8 Extensions
@@
 10 Conclusion
-Divergence-constrained alignment of LLMs naturally induces a semiparametric single-index model with the LLM policy class parametrizing the index.
-This requires only the probability one prefers one demonstration over another be some unknown increasing function of the difference in rewards.
+Divergence-constrained alignment of LLMs induces a semiparametric single-index model with the policy class parametrizing the index.
 This significantly relaxes assuming the link function is known exactly, as in DPO.
-However, it means we no longer have a clean likelihood in the policy parameter alone.
-Inspired by Cosslett (1983), we considered an orthogonalized quasi-likelihood with a plug-in for a high-dimensional nuisance function characterizing the distribution of preference given any candidate index.
-Our analysis gave finite-sample bounds and admitted nonparametric function classes with indices identifiable only up to scale.
-The dependence on plug-in error was squared, meaning it generally does not affect the learning rate, and the learner behaves as though the link function were known.
-The semiparametric preference optimization framework developed thus allows to align LLMs to human and AI feedback in a way robust to how these are linked to rewards.
+We considered policy learning by maximizing the profiled likelihood, orthogonolized quasi-likelihood, or bipartite ranking accuracy of the index implied by the policy.
+Our analysis gave finite-sample bounds and admitted nonparametric indices identifiable only up to scale.
+Empirical demonstrations showed how the SPO framework developed allows efficient alignment of LLMs to human and AI feedback in a way robust to how these are linked to rewards.
 Appendix A Proof of Theorem 2
@@
 Proof of Theorem 6.
-Let $\delta\in(0,0.5)$. Intersect the events of Lemma 16 (at confidence $\delta/3$) and Lemma 18 (at confidence $\delta/3$). On this intersection, Eq. 14 holds and the bounds Eqs. 15 and 16 hold.
+Let $\delta\in(0,0.5)$. Intersect the events of Lemma 16 (at confidence $\delta/3$) and Lemma 18 (at confidence $\delta/3$). On this intersection, Eq. 15 holds and the bounds Eqs. 16 and 17 hold.
 The condition $mh\geq c_{0}((v+2)\log(c_{3}m/h)+\log(1/\delta))$ implies that the right-hand side of Eq.
@@
 Thus, by a union bound, with probability at least $1-\delta$ all conditions of Lemma 21 hold.
-Apply Lemma 21, then substitute Eqs. 15 and 16 and $b_{+}\asymp h$, $b_{-}\asymp h$ from Eq. 14. This yields
+Apply Lemma 21, then substitute Eqs. 16 and 17 and $b_{+}\asymp h$, $b_{-}\asymp h$ from Eq. 15. This yields
 $\sup_{t\in\mathcal{T}}\sup_{u\in[-M,M]}|\hat{g}_{t}(u)-g_{t}^{\star}(u)|\leq c_{1}h+c_{2}\left[\sqrt{\frac{(v+2)\log(\tfrac{c_{3}m}{h})+\log(1/\delta)}{mh}}+\frac{(v+2)\log(\tfrac{c_{3}m}{h})+\log(1/\delta)}{mh}\right].$
@@
 From the second part of Eq.
-18, $\eta(w)$ is a.s.
+19, $\eta(w)$ is a.s.
 constant on level sets of $t(w)$, i.e.
@@
 $\sigma(t)$, and by Eq.
-20, the conditional expectation given $(t_{1},t_{0})$ is maximized by choosing the sign $\mathrm{sgn}_{0}(g(t_{1})-g(t_{0}))$ equal to $\mathrm{sgn}_{0}(\bar{\eta}_{1}-\bar{\eta}_{0})$.
+21, the conditional expectation given $(t_{1},t_{0})$ is maximized by choosing the sign $\mathrm{sgn}_{0}(g(t_{1})-g(t_{0}))$ equal to $\mathrm{sgn}_{0}(\bar{\eta}_{1}-\bar{\eta}_{0})$.
 Hence $\mathrm{AUC}(g(t))\leq\mathrm{AUC}(\bar{\eta})$ for all $g$, in particular $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$.
 Combining Eqs.
-21 and 22 and $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$ yields
+22 and 23 and $\mathrm{AUC}(t)\leq\mathrm{AUC}(\bar{\eta})$ yields
 $\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\ \geq\ \mathrm{AUC}(t^{\star})-\mathrm{AUC}(\bar{\eta})=\frac{1}{4\pi_{0}\pi_{1}}\Big(\mathbb{E}|\eta_{1}-\eta_{0}|-\mathbb{E}|\bar{\eta}_{1}-\bar{\eta}_{0}|\Big).$ (24)
@@
 ∎
-Under (H2)–(H5), $\eta(w)=\Psi^{\star}(t^{\star}(w))$ has support an interval $J=[\Psi^{\star}(\underline{t^{\star}}),\Psi^{\star}(\overline{t^{\star}})]$ and admits a density $q$ given by the change-of-variables formula $q(s)=p(r)/(\Psi^{\star})^{\prime}(r)$ where $s=\Psi^{\star}(r)$.
+Under our assumptions, $\eta(w)=\Psi^{\star}(t^{\star}(w))$ has support an interval $J=[\Psi^{\star}(\underline{t^{\star}}),\Psi^{\star}(\overline{t^{\star}})]$ and admits a density $q$ given by the change-of-variables formula $q(s)=p(r)/(\Psi^{\star})^{\prime}(r)$ where $s=\Psi^{\star}(r)$.
 Thus $q(s)\geq\underline{p}/\overline{\psi}=\underline{q}$ on $J$.
@@
 $\inf_{h:\mathbb{R}\to\mathbb{R}}\ \|\eta-h(t)\|^{2}\ \geq\ \underline{\psi}^{2}\ \inf_{f:\mathbb{R}\to\mathbb{R}}\ \|t^{\star}-f(t)\|^{2}.$
-Plugging into Eq. 27 and using (H6) yields the global inequality
+Plugging into Eq. 28 and using Assumption 5 yields the global inequality
 $\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)\ \geq\ \lambda_{0}\ \inf_{a\in\mathbb{R}}\ \|t^{\star}-at\|^{2},\qquad\lambda_{0}=c_{\rm MSE}\,\underline{q}\,\underline{\psi}^{2}.$ (29)
@@
 Then $s$ is a positive multiple of $t$, so $\mathrm{AUC}(s)=\mathrm{AUC}(t)$, and $d(t)^{2}=\|t^{\star}+s\|^{2}$.
-We need a continuity bound comparing $\mathrm{AUC}(s)$ and $\mathrm{AUC}(-t^{\star})$.
-First, under (H2)–(H4), the conditional density of $r=t^{\star}(w)$ given $z=1$ is $p_{1}(r)=\eta(r)p(r)/\pi_{1}$, hence
+We need a continuity bound comparing $\mathrm{AUC}(s)$ and $\mathrm{AUC}(-t^{\star})$. First, the conditional density of $r=t^{\star}(w)$ given $z=1$ is $p_{1}(r)=\eta(r)p(r)/\pi_{1}$, hence
 $\|p_{1}\|_{\infty}\leq\frac{1-\varepsilon}{\varepsilon}\,\overline{p}=B_{0}.$
@@
 By Eq.
-28, $d(t)^{2}\leq(\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t))/\lambda_{0}$.
+29, $d(t)^{2}\leq(\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t))/\lambda_{0}$.
 Let $\Delta=\mathrm{AUC}(t^{\star})-\mathrm{AUC}(t)$; then $d(t)^{2/3}\leq(\Delta/\lambda_{0})^{1/3}$.
@@
 Since $\mathcal{J}_{n}\leq 1$ (take $\eta=1$ in Eq.
-33), the $M$-term is always bounded by $C_{0}(\log(1/\delta))/n$ and is absorbed by the explicit $(\log(1/\delta))/n$ term in Theorem 10.
+34), the $M$-term is always bounded by $C_{0}(\log(1/\delta))/n$ and is absorbed by the explicit $(\log(1/\delta))/n$ term in Theorem 10.
 Likewise, by $ab\leq(a^{2}+b^{2})/2$,
@@
 By symmetrization and a truncated entropy-integral bound for Rademacher averages (as in the definition Eq.
-33), $\sqrt{n}\,\mathbb{E}\sup_{r}|\nu_{n}(r)|\lesssim\sqrt{A}\log n$ under $p=1$.
+34), $\sqrt{n}\,\mathbb{E}\sup_{r}|\nu_{n}(r)|\lesssim\sqrt{A}\log n$ under $p=1$.
 Thus we may take $\phi(\sigma)\equiv C\sqrt{A}\log n$ (constant in $\sigma$), which is admissible.
@@
 Noting that $\delta\leq 0.5$ means the statement is true for some $c>0$. ∎
+Appendix G Additional detail on experiments
+G.1 Synthetic experiment
+We consider preference observations $(x,y_{0},y_{1},z)$ with $x\in\mathbb{R}^{20}$ drawn i.i.d.
+from a standard Gaussian and $y_{0},y_{1}\in\{1,\dots,10\}$ sampled independently from the uniform reference policy $\pi_{\mathrm{ref}}(y\mid x)=0.1$.
+A teacher policy $\pi_{\theta^{\star}}$ is a randomly chosen feedforward network with two hidden layers (32 units each, ReLU activations) and a 10-way softmax output.
+The reward is defined as $r^{\star}(x,y)=10\log(\pi_{\theta^{\star}}(y\mid x)/\pi_{\mathrm{ref}}(y\mid x))$.
+Given a pair $(y_{0},y_{1})$ we set $u=r^{\star}(x,y_{1})-r^{\star}(x,y_{0})$ and draw $z\sim\mathrm{Bernoulli}(g(u))$ where $g(u)=0.5\,\sigma(4(u-s))+0.5\,\sigma(4(u+s))$, $\sigma$ is the logistic link, and $s\in\{0,0.25,0.5,0.75,1.0,1.25,1.5\}$ controls misspecification.
+Policies $\pi_{\theta}$ are parameterized by the same architecture as the teacher and trained by batched SGD (Adam, learning rate $2\times 10^{-3}$, 100 epochs, batch size 128).
+DPO minimizes the standard logistic loss on the implied reward index.
+PSPO alternates between (i) isotonic regression for the nondecreasing link function via PAVA and (ii) SGD updates of $\theta$ using the fitted link; we use a short DPO warm-start and then 4 outer iterations with 2 inner epochs.
+OSPO estimates the link by kernel regression of $z$ on the index $t_{\theta}(x,y_{0},y_{1})$ using a Gaussian kernel with bandwidth of $n^{-1/5}$ and an in-memory sample of all training points; we use a short DPO warm-start and align the sign of the implicit reward by the empirical covariance between $t_{\theta}$ and $z$ on the training data.
+RSPO uses the PoP-DPO loss computed within batches (so a batch of $B$ preference observations involves $B^{2}$ loss terms).
+For each learned policy $\pi_{\theta}$, we evaluate the family $\pi_{\beta,\theta}(y\mid x)\propto\pi_{\mathrm{ref}}(y\mid x)\exp(\beta^{-1}h_{\theta}(x,y))$ with $h_{\theta}(x,y)=\log(\pi_{\theta}(y\mid x)/\pi_{\mathrm{ref}}(y\mid x))$ on an independent evaluation set of size $m=2000$.
+We scan a fixed grid of $\beta$ values and report the resulting average reward and KL divergence.
+For reward-vs-$s$ plots, we fix a KL budget $\kappa=0.2$ and, for each run, pick the two $\beta$ values that yields KL closest to $\kappa$ and interpolate linearly between them.
+We average the results over 1000 seeds.
+G.2 Qwen3 alignment experiment
+Dataset and prompts.
+We use the UltraFeedback dataset (openbmb/UltraFeedback), split train.
+For each example we extract a prompt using the first available field among prompt, instruction, query, or question.
+We shuffle the dataset with a fixed seed and take a prefix of size $n=5000$ for each experiment.
+Reference model and SFT.
+The base model is Qwen3-0.6B. The reference policy $\pi_{\mathrm{ref}}$ is a light SFT model trained as follows:
+  • Data: 2,000 UltraFeedback examples, shuffled.
+  • Target response: the completion with the highest overall_score (fallback to fine-grained_score).
+  • Training: 1 epoch, learning rate $2\times 10^{-5}$, max length 1024, per-device batch size 1, BF16.
+Reward model and preference generation.
+We use Skywork-Reward-V2-Qwen3-1.7B as the ground-truth reward model.
+For each prompt, we sample two responses from $\pi_{\mathrm{ref}}$ using nucleus sampling (temperature 0.7, top-$p$ 0.95) with a max generation length of 128 tokens.
+The reward model scores each prompt-response pair by concatenating prompt and response and taking the sequence classification logit.
+We define the reward difference $\Delta r=r_{1}-r_{0}$ and rescale it to have standard deviation 1 across the dataset. Preferences are sampled using the shifted-mixture logistic link function
+$g(u)=\tfrac{1}{2}\sigma\left({u-s}\right)+\tfrac{1}{2}\sigma\left({u+s}\right),$
+with shift $s$ and temperature $\tau$. We then draw $z\sim\mathrm{Bern}(g(\Delta r))$.
+Policy parameterization and log-probabilities.
+All optimizers train a LoRA-adapted policy on top of the base model.
+We use LoRA with $r=16$, $\alpha=32$, and dropout 0.05.
+The reference model is kept frozen.
+Log-probabilities are computed only on response tokens by concatenating prompt and response, masking prompt tokens, and summing log-probs over response tokens.
+Optimization methods.
+We train with AdamW (learning rate $5\times 10^{-6}$, weight decay 0), batch size 4, gradient accumulation 4, and a cosine schedule with 10 warmup steps.
+We clip gradient norm at 1.0 and skip updates on non-finite losses.
+Default training length is 100 steps in the main experiments.
+DPO.
+We implement the standard DPO loss with temperature 1:
+$\ell=\log\left(1+\exp\left(\log\frac{\pi_{\theta}(y_{\mathrm{other}}\mid x)}{\pi_{\mathrm{ref}}(y_{\mathrm{other}}\mid x)}-\log\frac{\pi_{\theta}(y_{\mathrm{pref}}\mid x)}{\pi_{\mathrm{ref}}(y_{\mathrm{pref}}\mid x)}\right)\right).$
+PSPO.
+We alternate between gradient updates on $\theta$ and isotonic regression to estimate $\Psi$.
+At each update of $\Psi$, we compute $t_{\theta}(x,y_{0},y_{1})$ over the dataset and fit an isotonic regression with the Pool Adjacent Violators Algorithm (PAVA).
+The isotonic fit is updated every 50 gradient steps.
+To allow gradients through $\Psi$, we use a soft nearest-neighbor interpolation (softmax over distances with temperature 0.1) and mix in a small logistic term (5%) to avoid flat regions.
+Outputs are clipped to $10^{-6},1-10^{-6}]$.
+OSPO.
+We estimate $g^{\star}_{\theta}(u)$ by kernel regression over a memory buffer of past $(t_{\theta},z)$ values.
+The memory is a FIFO buffer with maximum size 4096.
+We normalize by the standard deviation of stored $t_{\theta}$ values, use a Gaussian kernel, and bandwidth 0.5.
+The memory values are detached so gradients do not flow through the kernel regression.
+We clip $\hat{g}$ to $10^{-4},1-10^{-4}]$ and mix with $0.05\cdot\sigma(t_{\theta})$ to avoid flat gradients.
+After training, we align the sign of $t_{\theta}$ using the sign of the empirical covariance between $t_{\theta}$ and $z$.
+RSPO (PoP-DPO).
+We implement the pairs-of-pairs DPO variant by forming all pairwise sums within a minibatch:
+$\ell=\frac{1}{B^{2}}\sum_{i,j}\log\left(1+\exp\left(-\left(s_{i}+s_{j}\right)\right)\right),\quad s_{i}=\log\frac{\pi_{\theta}(y_{i,\mathrm{pref}}\mid x_{i})}{\pi_{\mathrm{ref}}(y_{i,\mathrm{pref}}\mid x_{i})}-\log\frac{\pi_{\theta}(y_{i,\mathrm{other}}\mid x_{i})}{\pi_{\mathrm{ref}}(y_{i,\mathrm{other}}\mid x_{i})}.$
+Evaluation and KL interpolation.
+For each trained policy and each $\beta$ on a grid, we generate responses from the mixed policy
+$\pi_{\beta}(y\mid x)\propto\pi_{\mathrm{ref}}(y\mid x)^{1-1/\beta}\,\pi_{\theta}(y\mid x)^{1/\beta}.$
+We implement this by mixing token logits at every step:
+$\ell_{\beta}=(1-\beta^{-1})\ell_{\mathrm{ref}}+\beta^{-1}\ell_{\theta},$
+with an extra sign flip for OSPO if the sign alignment was negative.
+Sampling uses temperature 0.7 and top-$p$ 0.95 with max 128 new tokens.
+We compute the average reward using the reward model and estimate reverse KL as the mean difference of log-probabilities between $\pi_{\beta}$ and $\pi_{\mathrm{ref}}$ over generated tokens.
+For convergence curves at a target KL $\kappa$, we linearly interpolate the reward between the two neighboring $\beta$ values that bracket $\kappa$ on the KL axis.
+If $\kappa$ is outside the observed range, we use the nearest point.